A Complete Bibliography of Publications in Communications in Statistics: Simulation and

Computation: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.02

Title word cross-reference

2 [Dra97]. 2 × 2 [ATL92]. ^{1} [BHK91]. ^{2} [CS95a, Tsu95]. ^{3} [Hun98]. _{p} [Ekb90, KJ99]. _{pk} [FW92, KJ99, WMF91]. _{pm} [KJ99]. A [GK97a, HK93, SA99]. α [PTP92]. AR(1) [KS98, Mag91, TW98]. ARMA(1, 1) [Ros93]. X̄ [PR94]. β [Pli90]. c_{pk} [PC96]. c_{s} [PC97]. χ^{2} [NM98]. D [MY98]. F [Din97, Gui90, Sut90, Viv90]. Γ [VI94]. i × j [PKD99]. K [Kha97, Hor98]. k(> 2) [TP90]. L^{B} [TJ94]. l_{1} [Dun92, HC97]. L_{2} [Dha92]. L_{B} [TJ90]. M [Hil91, HW96, Kha97, Won92, Abo92, SM98b]. MA(1) [AMJM96]. P [Bri93, AW96]. S_{U} [Cas98]. SETAR(2; 1, 1) [Lim99]. t [Aka95, BD97, Bla91, DM93, Fan94, KTV94, KK91, Lee92, O’G97, Sut90, Vau92]. T^{2} [CMY99, RJ91]. × [LU95]. W [Rah99]. x [SSM98].

-chart [PR94]. -CUSUM [Dra97]. -distance [Dha92]. -distribution [Aka95, KTV94, Vau92, Gui90]. -Estimates [HW96]. -estimation [SM98b, Hil91]. -estimators [Abo92, Won92]. -intercept [SSM98]. -minimax [VI94]. -optimal [HK93, MY98]. -risks [PTP92]. -step [Kha97]. -test [O’G97, Sut90]. -values [AW96, Bri93]. -variate [Hor98].

1 [GK92, Jai92, RM96].

2 [JR96, Kan93, Mar98, SK93].

3 [Don96].

Abbott [Hoe91]. absolute [RRN91, RKN91b, WZ98]. abundant [ALN95]. Accelerated [Hsi94, AGAA98, BLN94, LW99]. Acceptable [GS90b, GS92]. Acceptance [SR97a, HM92, Lin99]. accounts [Mor99]. accuracy [CGR99, OR95, Wen96]. Accurate [CT99]. Act [SK93]. ad [SS98]. ad-hoc [SS98]. Adapting [SW98b]. Adaptive [AG98, SRHW99, Dia99, cLH97, O’G96, O’G97, ON96, Phi98, Ste93, TPSS99]. Additional [BS94b, Pur94, RR92, RR94]. adequacy [Jai91]. Adjusted [ZB94, HRM99]. adjusting [Mag91]. adjustment [GPPL99, Hol91, SC93a]. admissibility [CV94]. affine [Vid93]. against [Cha90, FH98, GD95, LW94, NdL97, TDTC93, Tsu95, Wen96]. age [AWhW93]. agglomerative [CV94]. Algorithm [HW96, LK90, Alb96, AAA97, Cut93, Din94, GS94, HL90, KH92, LU95, LEKR97, Lim95, MMM95, ME94, MABL90, Mor92, Ngu93, Rac93, Was93]. algorithms [Fan94, Hir97]. allocation [TB99]. alphabet [Sch96, SX97]. alternative [Bed95, Cha90, HD94, Nis97, SM98b, TH99, Tsu95, Wol96]. alternatives [FH98, NdL97, Och93, SSW95]. among [KJ99, Sof99]. analogue [MTvdM98]. Analyses [Gil90, HPL95]. Analysis [Ano94a, BLN94, HK90a, JK98, LV99, dFMK92, PC94, PR96, AV92, AKJ97, BM96a, BS99, CC98a, Dor93a, Dra97, DMS99, Fan99, FR95, GB97, GT90, GS94, HK90b, KAKW98, Lun99, MLA^{+}99, McC96, ME91, MN91, Nis97, PK95, Pla98, RR92, RWB97, Ren92, RFL91, SWFK96, SC93a, ST93, Sof99, SW98b, SR97b, TQK97, TYF98, YMS93]. Analytical [BF92, GPPL99]. anchor [Sim95]. Anderson [MTT90]. animal [BM96a, Lee96]. annealing [AAA97, Li96]. ANOM [KR98]. ANOM-type [KR98]. ANOVA [GW98, OR95, RBL91, SKBK98, Wan97]. antithetic [Tan99]. AOQL [GK91]. apparent [Ren92]. application [Cas98]. applications [BCC97, BJ94, Cer92, KDM96, KL90, KT92, ÖA94, ÖR92, Phi90, RF96, SK93, THD98]. applied [HM92, KH92]. approach [Fal99, LH97, LW99, Nis97, OR95, PS99, SSM98, Taa97, TH99]. approaches [KAKW98]. Approximate [Hoe91, KG91, LM97, Viv90, Wan91, BS94b, jCM97, DS91, SR97b, Wan94]. Approximated [NH97]. Approximating [Gra93, SM98a, Sho98]. Approximation [AGV96, Cas98, Aka95, BD97, Bog93, CT99, FWW98, Pen95, SMR95, Zhe95].

Approximations [Dre90, Kim94, KK99, MN99, SSW95, TA92, DM95, HN95, JG95, Ken90, Sho97, Tho95, TW98, Wan94]. AQL [GK92, GK91]. arbitrary [Haw92]. arch [SL97]. ARIMA [Hun90]. arising [CL91]. ARMA [CPS97, KP90, Sal95]. arrays [ALS95]. ARRLA [Sal97]. art [SF93]. artificial [Hal96]. aspects [AWEBW92, CAF96, SM91]. assays [RLR98]. Assessing [RS90, SG93a, Sug93, CPC97, HL93, ÖR92, SS98, Sug94]. Assessment [Gar94, SWFK96, SM98c]. association [Bet93, HG98, JK98]. assumed [KJ98]. assumptions [FR95]. asymmetric [Boy94, Mad97]. asymmetrical [EHAA94]. asymptotic [ALM93, And91, AMH92, Dor93a, Sal97, Won92]. Asymptotically [RF96, RRN94]. attainment [BM96b]. attribute [ALN95, SA90]. Attributes [GS90b, GK91, GS92]. Author [Ano91a]. auto [SC93a]. autocorrelated [CS95b, KHW93, PCH97]. autocorrelation [ACM94, CG93, ME91]. autocovariances [BG93]. Autoregressive [HM92, AG98, BG93, CFBGM^{+}97, Fuk94, Had98, Kul96, SA95, TA92]. autoregressively [LM94]. auxiliary [Smi90, Won92, YSY96]. available [BHK91, RR94]. Average [Gan90, ACM94, AS91, Dun92, GPPL99, Haw92, JC93, KUK96, MP98, SAL92, SAS91, WC99].

B [AGV96, GR92]. B-splines [AGV96, GR92]. bad [Sch94]. Balanced [THWL97, PB94, Wan92, YSM94]. ball [BD93]. bands [ABL90, Ryc92, Yeh96]. bandwidth [Chu95]. Bartlett [Hol91]. Based [LOSB90, Abo92, AY90, Bar96, BN98, BS94b, Cor97, CAF96, ENN92, Gan94, Jai92, KD99, KW96a, LT95, O’G96, PBB95, PHB97, RR99, SM98c, SA94, TYF98, TP90, TI90, Was93, WS93, WZ98, Yeh96]. basic [KJ99]. basis [Par99]. Bayes [KO90, Mad97, MABL90, NGM98, US92]. Bayesian [AK93a, BS93, Fan94, GB97, GS94, Lav92, LW99, Lin99, PC96, PC94, SA95, SR97b, TH99]. Bayesian-like [PC96]. be [KW99]. behavior [LL92, LWC91, Lee96, Mur99]. Behrens [CR97, Meh97, TH99]. Bernoulli [Taa97]. Best [HS90, BC92b, LK97]. beta [JG95]. Better [VC97, ŠŠ92]. between [BD97, Fuk94, HG98, MF93, PM97, SlC97, VC99]. Bias [CNBCF98, RFL91, BHG94, CCN98, PM97, PSCY97, dCQAC97, Ren92]. biased [CT92, TDTC93]. bifurcation [Bet95]. bilinear [LEKR97]. binary [AC98, HD98, Her97, KHW93, MP97, TQK97, YPW94]. binomial [AY90, DM93, Fam97, MST93, Ong92, RLR98]. bioavailability [CPC97]. Bivariate [ORR90, ABL90, jCM97, CD93, GS98, KK99, dFMK92, MW94, Ong92, SG92, Vau90]. Biweight [Wil90a]. black [SS98]. blended [PHB97]. blinded [Gou97]. Block [Gil90, Ano94a, Ngu93, SD96]. blocking [PJ99]. blue [TYF98]. Blum [Bar94]. Board [Ano91b, Ano91c, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95b, Ano95c, Ano95d, Ano96b, Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Smi99b]. Bootstrap [DG96, RK96, STY93, SF94, Yeh96, CFBGM^{+}97, HCNA97, KHW93, PSCY97, SlC97, WMF91]. Bootstrapping [BW97, Kul96, CFMM95, THWL97]. Both [TJ94]. bound [Cut93, De 90]. boundary [MX92]. bounded [GS90a, GPPL99, Vid93]. Bounds [Ken90, KG91]. Bowman [BS94b]. Box [GT90, Hal96, LV99, Rah99, Zhe97]. boxes [SS98]. branch [Cut93]. branching [Dev92, DE99]. brittle [BDP90]. Brocess [Boy94]. Brown [KW99, Meh97]. Buckley [Hil93, Hil94]. Buckley-James [Hil93]. Burr [SG93b].

C [Smi97b, WMF91, FW92, KJ99, Smi97b]. calculate [God94, WMF91]. Calculating [Don96]. Calculation [NR95]. calibrated [Rod99]. calibration [NH97]. can [Sch94]. candidate [SK93]. Canonical [Ct90, AKJ97, SU92]. can’t [KW99]. Capability [LOSB90, Boy94, CP93, KJ99, Luc96, PC97, WMF91, YB98]. CAPM [KO90]. CAPM-Type [KO90]. capture [Lee96]. capture-recapture [Lee96]. carcinogenesis [Zhe98]. Carlo [AH95, BF92, CGR99, DP90, Die92, EK93, Fuk94, GT99, Gil90, HCNA97, JLPR99, Phi98, Sam99, Sch93, Tan99, WAD^{+}92, ZB94]. case [Has95, HO94, JY95, LM94, RB90]. Casella [SE93a]. categorical [ALM93, GB90, JK98, Tho95]. categories [LU95]. Cauchy [Hol95]. cautionary [PA94]. cdf [ME94]. Cell [KG90, SS96]. Censored [HW96, Mag90, BW97, Cha90, ENN92, GS90a, Hil91, JY95, KW96a, LML94, RL96, TP90, Zho92]. censoring [AGAA98, AK93a, CC98b, DS91, Gor90, USS96, Vau90]. censorship [Che94]. Central [KAS90, Aka95, GL97, KK99]. centroid [Bar96]. centroid-based [Bar96]. certain [Dha92, Hol91, Raj92, TI90]. Chacteristic [Per90]. chain [GT99, JR96, PR96, RM96, RLR98]. chains [THWL97]. Change [JW90, BuH97, Nag90, PP98, SO94]. change-points [SO94]. changepoint [SW90, WC90]. changing [And94]. chaos [LH97, Lim95]. chaotic [Lim99]. character [YSY96]. characteristic [Nai96, THD98]. characteristics [Bed95, GGC95]. characterization [EDMY91]. characterizations [EDMAYS97]. Characterizing [Tör90]. chart [CL94, DW94, Gan94, PR94]. Charts [Gan90, Ale99, Haw92, LCW95, Nis97, PTP92, PCH97, SW98b, VR97, Wik97, XGX97]. charts* [MP98]. Chi [KAS90, KG90, ATL92, Din94, FWW98, GB90, KK99, Tho95, Wan94]. Chi-Square [KAS90, GB90, KK99]. Chi-Squared [KG90, ATL92, Din94, FWW98, Tho95, Wan94]. chief [Smi99a]. choice [BG93]. Choosing [Don90, Adj91, KPA91]. Chow [CPC97]. ChSP [JR96, RM96]. ChSP- [JR96, RM96]. chu [PP98]. chu-white [PP98]. circular [Hol95, SJNM91]. class [GS94, OM95, RKN91a, SR97a]. classic [And91]. Classical [AK93a, AR90, Sug93, Sug94]. classification [BHK91, Ren92, SA95, Swi94, YSM94]. clinical [MLA^{+}99]. closed [Luc97]. closed-loop [Luc97]. Cluster [Lun99, dFMK92, dK96]. clustering [CV94, Sof99]. clusters [LG99]. Cochran [CS95b]. Coefficient [Ame90, Bet93, Bro97, Die92, HG98, PB94, SG92, Vau90]. Coefficients [Wan90a, Wan90b, GS98, HC97, HL93, Mil91, SH93]. collected [MT91]. combinations [Wan91]. combined [PBB95]. Combining [Sch93, RGU97]. Comments [Smi99a]. committing [HEB93]. common [BGDH90, GB97, Wee96]. commonly [Sho97]. commonly-used [Sho97]. Comparative [PKD99, BA93, KW96b, YY96]. Comparing [KC95, Wil90a, Wil90b, And91, AMH92, BS94a, Che94, DM93]. Comparison [Bri90, DP90, Dor93a, Hil93, LNZ94, PCH97, Vau90, AC98, BN98, BGDH90, CR97, DLP94, DM95, EHAA94, Hir97, KAKW98, KH96, Mad97, O’G97, SO95, SlC97, Sof99, SK91, SKBK98, WAD^{+}92, WZ98, WW98, WC99]. Comparisons [Bri90, DS91, HS90, SW90, ALM93, BF92, Bri93, Che93, CMY99, DG94, HEB93, HO94, Ken90, RW96, SE93a, Sea97, TI90, WG99]. complete [Ano94a, SD96, TP90]. completion [KDM96]. component [Boi96]. Components [Wan90a, Wan90b, KG91, MMM95, SC93b, Wal90, YSM94]. composite [GL97]. compound [Has95]. Computation [Dha92, KAS90, SJNM91, Bar98, CS95a, CGR99, GT90, Sal97]. Computational [AWEBW92, Fan94, Fan99, CAF96]. Compute [HS90, ME94]. Computer [BK91, ORR90, CFBGM^{+}97, Ong92]. computers [Wol92]. Computing [Gui90, KL90, Par92a, Par92b, DG94, Din94, Din97, Dor93b, Had98, Str97]. concavity [LW94]. concept [Yeh96]. concurrent [RDW96]. condition [Far92, MX92]. Conditional [GH91, AY90, Raj92, Sch96, SX97, VC99]. Confidence [CFMM95, LOSB90, PB94, Ryc92, Wan90a, Wan90b, Adk90, ABL90, ATT97, CC99, CT93, Cor97, DS91, FW92, GS90a, KG91, Luc96, MJ97, NH97, PSCY97, SY93, SF94, TG93, VC97, VC99, Wan91, WMF91, Yeh96]. configurations [HG98]. conforming [XGX97]. confounding [BS94a]. conjugate [HB98, Lav92, Mac94]. conservative [FW92]. considerations [JC96, Lin99]. consistent [CFBGM^{+}97]. constant [Bér92]. constants [JLPR99]. Constrained [KK97, Cut93]. constraints [AAA97]. constructing [CFBGM^{+}97, CK95a, Ngu93, Phi90]. construction [Kan93, MY98]. contamination [Won92]. contemporaneous [RDW96]. contestants [Cro90]. contingency [LU95, PKD99]. continuity [ATL92, SWFK96]. Continuous [Was90, KH92]. contoured [GK97a]. contribution [RS90]. Control [Bri90, Gan90, KC95, Ale99, AS91, BM96b, Bri93, Cas98, CL94, Che93, Che94, CW96, DW94, Gan94, HL92, HEB93, KL90, LCW95, Luc97, MZ91, MP98, PTP92, PCH97, SE93a, SAL92, SAS91, SS95, SW98b, SA94, TG93, VR97, Wik97, XGX97]. controlled [Con92]. controllers [SRHW99]. controls [RL90]. convergence [Rod99]. convex [Mit93]. Cook [KS96]. Copulas [Sun90]. Cornish [BS97]. correct [SK92]. corrected [DM93]. correction [ATL92, FW92, Hoe91, SWFK96, Sug94]. Correlated [Gil90, BG95, Her97, JK98, LM94, ÖA94, PA94, Sch92, SlC97, SS95, TQK97, YPW94]. Correlation [Ame90, KP90, AKJ97, Ano90, Ong92, PS99, SC93a, SH93, SG92, SL97, SU92, VR97, Vau90, WCWB96, dK96]. correlations [Fal99]. cost [AAA97, Lin99]. count [Sil97, XGX97]. counting [Nic93]. Counts [JW90, Zha94]. Covariance [Wol93, dCQAC97, jCM97, De 95, Dor93b, Had98, KAKW98, Mor99, Sal97, Sal99, Sug93, Sug94, TI90]. covariances [BC92a, Par92b, Vau92]. covariates [Her97, LW96, Leu94]. covergence [Luc94]. Cox [Hal96, LW96, Rah99, Sch93, Zhe97]. cramér [RL96, KSS96]. cramér-von [RL96]. Cressie [Par99]. criteria [De 95, HOJ97]. criterion [Bar94, RK99, Swi94]. Critical [FN98, HS90, AMH92, Bla91, HPL95, RWB97, SE93a]. Cross [De 95, LK90, Rac93]. Cross-Validation [LK90, De 95, Rac93]. crossed [KG91]. crossover [AS95]. CSP [Kan93, KG93]. CSP-2 [Kan93]. CSP-T [KG93]. cubic [HK93]. Cumulative [Pan92, ABL90, Bog93, Gan94, Haw92, Sho97, XGX97]. curious [KTV94]. current [RB90]. curvature [Rav94]. curve [SO95, SS98, Yeh96]. curve-fitting [SS98]. curves [MH98]. cusum [JC93, Nis97, VR97, Was93, Wik97, Dra97]. CV [Don96]. cyclic [LWC91].

damage [Pur94]. Darling [MTT90]. Data [HW96, JW90, AC97, BM96a, BP97, BW97, BuH97, Cha90, CT92, CTH^{+}94, CC98b, FR95, FR93, GS90a, GH91, GB90, Haw92, HD98, Her97, Hil91, Hir97, HK90b, JK98, KDM96, KD99, KHW93, KE96, KW96a, LM94, LML94, Lee96, LV99, LM97, LG99, Luc97, Lun99, dFMK92, MMM95, MRS91, Ogd97, OM95, PW92, PS99, PA94, PC94, Pie95, PCH97, RR92, RR94, RL96, RKN91a, SWFK96, SW98a, SD96, SC93a, Sil97, Smi90, SGF^{+}91, SKBK98, Sug93, Sug94, Taa97, TQK97, Tho95, WAD^{+}92, WZ98, YPW94, YSM94, Zho92, dK96]. default [Sch94]. deficiencies [Bér92]. degree [Adj90]. Degrees [HK90a, Mor99, Vau92, Wan94]. Delayed [Ata99]. Delicate [KJ99]. delta [MABL90]. densities [NR95]. Density [JD95, Cer92, GR92, Lye93, Rod99]. departures [HL93]. Dependence [Sun90, MF93, Taa97]. dependent [CFMM95, LW96, OR95, YMS93]. dephased [PSMR95]. depth [Yeh96]. Design [Cor97, GL99, Ano94a, Dra97, Gan94, GD95, Kha97, McC96, PR94, ST93, Was93, XGX97]. Design-based [Cor97]. Designs [Gil90, BS94a, EHAA94, EK93, GL97, GT90, HK93, HAS98, KUK96, LGG97, MN91, MY98, MW94, NSS99, Ngu93, PJ99, Sch92]. Detecting [Pie95, CL94, WZ98]. detection [ALS95, Mur99, WS93]. Determinants [Ct90]. Determination [GD95, GK91, LK97, Jai91]. determined [Lee92]. determining [Lim95]. Development [LT95, SM98b]. deviation [God94, GPPL99]. diagnosis [HRM99]. Diagnostic [HL93, LW96, OR95, Wen96]. Diagnostics [Rod99, CP92, TQK97, Wal90]. dichotomizing [SW98a]. dichotomous [dFMK92]. Difference [Zho92, AW96, CT93, HN95, SY93, SE93b]. differences [WZ98]. different [BF92, Mor99, SO95, SA95]. diffusion [GS99]. digits [MR96]. dilution [RLR98]. dimensional [PR96]. directional [Lun99]. Dirichlet [Mac94, Nar91]. Discrepancy [FZL98]. discrete [ATT97, DN97, FH98, KT92, KTTB92, Luc97, OM95, dOP99, RR99, Shi99]. discretized [dOP99]. discriminant [AK91, Dor93a, Hou93, Leu94, PK95, Ren92, RFL91]. dispersal [KT92]. dispersion [FR93, KT92, Sho99]. distance [CAF96, Dha92, HB94, KS96, PBB95, PHB97]. distance-based [CAF96]. distances [BN98, PBB95]. Distributed [DG90, Fal99, SKG96]. Distribution [ASA90, BSG98, Bri90, BY90, Gui90, KAS90, KC95, Viv90, AGAA98, ABL90, Aka95, AK93b, AY90, AWBW^{+}96, BC92a, BC92b, Bar98, BLN94, Bet93, BDP90, BT92, Bog93, BKS92, Cas98, CS95a, CTH^{+}94, jCM97, Cor97, Din97, DS91, DSW96, EDMY91, EDMAYS97, ET93, Fam97, FL92, FWW98, GS98, Haw92, Hoe91, Hor98, HL90, JG95, KDM96, KD99, KH92, KTV94, KA99, KL90, KK91, Kul96, KW96b, LW93, Mar98, MST93, Nar91, Ong98, Pan92, Phi90, RKN91a, RG96b, SMR95, Sch96, SG92, Sho97, Sho98, SM91, TA92, Vau92, VC97, Wol96, Woo94]. Distribution-Free [Bri90, BT92, RG96b]. distributional [FR95]. Distributions [ORR90, Per90, TJ90, ATT97, App95, AWEBW92, AWBW^{+}96, BD97, BS94b, Cas98, CK95a, jCM97, Don96, DN97, EDMY91, Fan94, Gar94, Gor90, HD98, Hol95, Hsi94, KDM96, KH92, KK99, KTTB92, LS98, Luc94, Lye93, MZ91, MR96, OM95, Pan92, Par99, PHB97, PC97, dOP99, RG96a, RRN91, RKN91b, RR99, SX97, SG93b, Shi99, SE93b, Sho99, SK91, SR97a, Swi94, TJ94, TW98, Wil90b, Zhe95, dK96]. distrlbution [KT92]. disturbances [CPS97, DM90, Has95]. divergence [PVEP97, Par99, Rod99]. divisibility [KH92]. domain [Chr99]. Donoho [Mur99]. dose [APN91, MF93]. dose-ranging [MF93]. dose-response [APN91]. double [AS95, GS92, RRN91, SAS91, SA90]. doubly [EDMAYS97, KK91]. driven [MS98]. drug [WW98]. dynamic [BM96a, LW99, WS93]. dynamics [DG94].

each [Her97]. easy [Smi97b]. echelon [Sal99]. echelon-form [Sal99]. ecological [Fül95]. Economic [PR94, XGX97]. Edgeworth [Mag91]. Edgeworth-adjusting [Mag91]. editor [Aka97, Smi99a]. editor-in-chief [Smi99a]. Editorial [Ano95a, Ano96a, Smi97a, Smi98, Ano91b, Ano91c, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95b, Ano95c, Ano95d, Ano96b, Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Smi99b]. Effect [PK95, ZWZ93, Chu95, O’G95, PJ99, SW98a, SC93a, TO97]. effects [Ano90, Che93, FR93, GB90, HZBW91, MF93, Ros93, TB99, Wan92]. efficacy [MF93]. Efficiency [CG93, Ct91, Jai92, LNZ94, Vid93]. Efficient [CS95a, DL97, LK90, DG94, Din94, Fro91, GT90, Rac93, Sam99, VI94, Was93]. eight [SH93]. elliptic [Sut90]. Elliptically [GK97a]. Empirical [BS93, KO90, LL92, ALM93, AC98, CW96, DN97, HS97, Kul96, MX92, Nai96, SH93]. employed [RFL91]. endpoint [RW96]. Enhancement [Sho97]. Entropy [WG99]. envelope [Rau92]. environments [DE99]. EPC [SRHW99]. equal [jCM97, GS98]. equality [BSB97, ENN92, KW99, KR98, Mil91, RR94, TP90]. equation [Fir91, HOJ97, THD98]. equations [GGC95, Has95, SlC97]. equivalence [CPC97, Raj92]. equivariance [SG93a]. Erlang [Mar98]. Erlang-2 [Mar98]. Errata [Ano94f, Sug94]. Error [Hun90, Pli90, BS93, CH95, CR97, Dor93a, DG96, Fro91, Hai96, HEB93, KJ98, PB94, RFL91, SJ91, SJNM91, TI90]. Errors [Gil90, KO90, Koz90, CD93, LW94, Mag91, Ros93, Sut90]. errors-in-variables [CD93]. estimate [Cer92, Hai96, Rod99, SSM98]. estimated [AK91, AH95]. Estimates [BY90, HW96, Mag90, BC92b, DG96, FR95, God94, MABL90, RL90, RFL91, SO95, Tay90, ZB94]. Estimating [CD93, DN97, Hal96, JW90, KP90, Lee96, Mac94, Mit93, PSMR95, Rah99, RB90, SS95, BS97, Dor93a, GB97, GGC95, HAS98, LY94, Phi90, SlC97, SK91, Vid93, Wan94]. Estimation [AGAA98, AY90, BM96b, CP92, DP90, FL92, GS98, KO90, KA99, KW96a, LM94, Li96, ORR90, Par99, Pla98, SK93, USS96, YSY96, AK93a, Agr95, All99, AMJM96, APN91, AWhW93, AG98, BDP90, BKS92, CA98, CT92, CH95, CPS97, DE99, Dun92, Edl90, ET93, Fam97, FR93, Fro91, Fül95, GR92, GL97, Gor90, Gou97, Hil91, KS98, KTTB92, JLPR99, LM97, Luc97, Mad97, Nar91, Phi98, PSCY97, RG96a, RF96, RLR98, RRN91, RRN94, RDW96, RKN91a, RKN91b, Sal95, SO95, SG93b, Shi99, SM98b, SKG96, Smi90, SG99, SK92, Ste93, TPSS99, VI94, WAD^{+}92, YMS93, Zhe97]. Estimator [Pli90, Adj90, Adk90, Bar96, Bed95, Blo98, CK95b, FZL98, Fir91, HKT93, Has95, Hil94, Hut90, JY97, Lye93, Mur99, PC96, SJ91, SA99, TDTC93, TA92, Wil98, Won92]. Estimators [Hil93, KO90, Abo92, AGV96, Bér92, Ct91, CG93, Dha92, Die92, DLP94, DM90, DL97, Hoe91, JC96, KK97, LNZ94, cLH97, PP98, Sal97, Sal99, SW90, SS98, Str97, TG93, TDTC93, US92, Wal90, WG99, Won92, YSM94]. estimators-improvements [WG99]. evaluate [HN95]. Evaluating [PM97, SS98, JY97, PJ99]. Evaluation [Haw92, Wen96, Her97, Sal99, SM98b, SA94]. Evaluations [FW99]. event [dOP99]. evidence [HCNA97, Phi98]. EVVMA [VR97]. EWMA [DW94]. Exact [ON96, ALM93, Fir91, Hir97]. exchangeable [Alb96, Fro91]. exciting [AG98]. expanded [RL95]. expansion [BS97]. expansions [RJ91]. Expectations [BG95]. Expected [Vau92, Par92a, Wan91]. Experience [Swi94]. experiment [Czi92, Hun98]. experimental [Gar94, MY98]. experiments [EHAA94, JY97, SS96, ST93, ZB94]. Exponent [Per90, Lim99]. exponential [AK93a, Agr95, ATT97, AW96, CC98b, CP92, CCN98, CNBCF98, EDMAYS97, ENN92, Fro91, Hsi94, JY95, KK97, LG99, RG96a, RRN91, SG92, SE93b, TP90, USS96, Wan97, WAD^{+}92, WC99]. Exponentially [Gan90, SAL92, AS91, MP98, SAS91, SKG96]. expressions [GPPL99]. extended [KDM96]. Extending [Hil91]. Extension [Kim99, HKT93]. Extensions [CK95b, Wil98]. extraneous [Mor99]. extravariation [Mor92]. extremal [BCC97]. extreme [BC92a, BC92b, CT99, DS91, Don96, Luc94]. extreme-value [Don96]. extremely [Sho98]. Extremes [Ct90, SG99].

Facilitate [TJ90, TJ94]. Factor [Bet95, HS97, MW94, SJ91, SS96]. factorial [EHAA94, Hun98, LGG97, McC96, SF93]. factorization [GK97a]. factors [GD95]. failure [BLN94, SE93b]. failures [Pur94]. Families [MN91, Wan97, WC99]. family [ATT97, CP92, CCN98, CNBCF98, DL97, LML94, RG96a]. Fast [HS90]. fibers [BDP90]. filter [Tan99]. filtering [RC95]. final [KDM96]. finite [AK93b, KW96b, MR96, RK96, San95, THWL97]. finite-sample [AK93b, KW96b]. first [GS99, SC93a, TA92]. Fisher [BS97, CR97, Meh97, SMR95, TH99, Woo94]. Fit [KG90, Alb96, CTH^{+}94, GB90, Hir97, NM98, Och93, Rau92, RL96, RR99, Swa92]. fits [BF92, KDM96]. fitted [MMM95]. Fitting [KD99, TJ90, AWEBW92, KDM96, Kul96, LU95, OM95, SS98, TJ94, Wan97]. five [HS97]. fixed [KW99, PCH97, TJ94, TO97]. flexible [Lye93]. fold [PB94]. folded [MR96]. following [CH95, RS90]. footrule [AC97, SQ90]. Forecast [Hun90, SA94]. forecast-based [SA94]. forecasting [BS99]. form [CH94, Sal99]. forms [THD98]. formulae [Jai91]. formulation [Ong98]. Forsythe [KW99, Meh97]. forward [LK97]. four [TW98]. Fourier [HS90]. fractional [Wal90]. fragmented [LV99]. Free [Bri90, BT92, Hoe91, RG96b]. Freedom [HK90a, Vau92, Wan94]. frequency [Chr99, Sim95]. full [Jai92]. fully [SR97b]. Function [KAS90, AGAA98, Agr95, Chu95, Con92, Cor97, DN97, Edl90, Hou93, HB98, KA99, KW96a, Nai96, Pan92, RB90, RR99, SJ91, Sho97]. functions [ABL90, AK91, Boi96, Cer92, Don96, Jai92, Kul96, Mad97, Mit93, Raj92, Zhe98]. Further [JC96, Phi98].

G [Jai92]. gamma [BSG98, ET93, LG99, MTvdM98, RKN91a]. Gauss [DG94]. Gaussian [AK91, CH94, CT99, GT99, MN99, Ong98, Sal97]. Gee [PS99, SKBK98]. Geisser [SC93a]. general [Alb96, BJ94, HK90b, PA94, Wol93]. Generalization [GT90, Hil94]. Generalized [PW92, Pli90, Sil97, AW96, AH95, BD97, BLN94, Bla91, Che94, Fam97, FL92, FT93, GGC95, HL90, KDM96, KD99, KA99, Meh97, Sal95, SlC97, SM91, SM98b, SA90, WCWB96]. generated [BK91]. Generating [BD93, DN97, HD94, Mor92, RR99]. Generation [DG90, Ekb90, ORR90, Dev92, Fal99, Ong92]. generator [Wol92]. generators [Bur90]. genetic [WZ98]. get [Sch94]. Gibbs [GB97]. Gini [Bet93]. Given [GS90b, GK91, GK92, GS92, Ong92]. GLM [HK90b]. Goldfeld [Adj90, Adj91]. Gompertz [Gor90]. good [DL97]. Goodness [KG90, Swa92, GB90, Hir97, NM98, Rau92, RL96]. Goodness-of-Fit [KG90, Swa92, GB90, Hir97, NM98, Rau92]. Graphical [BS94a, JY97, KR98, ÖR92, PJ99, SS96]. Greenhouse [SC93a]. GRKPACK [Wan97]. Group [Ste93, HPL95, Kim94]. grouped [HRM99, MMM95, RKN91a]. grouping [RRN94]. growth [BM96a, BS99]. guess [US92]. guide [BHK91].

H [Dia99]. H-splines [Dia99]. handling [Pie95]. Hartley [BCC97, CS95b]. having [BGDH90, LU95, LWC91]. hazard [Zhe98]. Hellinger [HB94, PHB97]. Hermitian [MH98]. heteroscedasticity [CC98a, Pie95]. heteroskedasticity [Adj90, Adj91, CNF95, SA91, SK90]. heuristic [Hil94]. Hierarchical [NGM98, Sof99, BRB95]. higher [ALM93, Aka95]. higher-order [ALM93]. Hilferty [MTvdM98]. histograms [Sim95]. history [KDM96]. hoc [SS98]. Hoeffding [Bar94]. homogeneity [Cha90, CS95b, PKD99, SD96, Sha91]. homogeneous [HB98]. Hotelling [CMY99]. households [YY96]. Huber [Ekb90]. human [WZ98]. hypercubic [HAS98]. hypergeometric [BKS92]. hypersphere [De 90]. Hypotheses [DP90, PVEP97, Sch93, Wee96]. Hypothesis [MP97, OR95, SC93b, RS90, Sut90, TO97, Tsu95, TPSS99].

identical [CL91, Wil90b]. Identification [CT92, LH97]. identity [KTV94]. ii [BC92b, CC98b, LGG97, RJ91, TJ94, TP90, USS96, Vau90]. iid [DSW96]. III [LGG97]. illustration [BRB95]. implementation [RC95, Sal95]. importance [JLPR99, Tan99]. Improved [CC98b, RK99, TG93, CNF95, GS99, KW99]. improvements [WG99]. Improving [BRB95, Meh97, Zha94]. Imputation [GB90]. include [Hil91]. incomplete [AC97, Ngu93, PC94]. incorrectly [KJ98]. indefinite [CH94]. independence [ALM93, Bar94]. independent [And91, AMH92, GB97, Jai91, Ken90, LG99]. index [Ano91a, GS94, Luc96, PC97, WMF91, YB98]. indexed [AS95]. indicators [AR90]. Indices [LOSB90, CP93, KJ99]. individual [Lee96, RS90]. inference [AC98, BA93, CC98b, KHW93, LW99, PBB95, PA94, PC94, SM91, SK90]. Inferences [SE93b, Fan94]. Influence [Wal90, Far92, GL97, Gra93, Hut90, Lav92]. influential [CT92]. Information [ASA90, HOJ97, Sun90, PSMR95, Sho98, Swa92, YSY96]. inspection [GG97]. instrumental [CD93]. integer [ET93]. integral [Dre90, Luc97]. integrals [De 90, DG94]. integrate [SRHW99]. integrated [SJ91]. integration [De 90, Tan99]. intensities [SE93b]. intensity [HB98, RB90]. interaction [Fuk94, SF93, Tho95]. interactions [KG91]. intercept [SSM98]. interim [HPL95]. intermediate [Bog93]. intermediate-precision [Bog93]. internal [Gra93]. interpolation [Kar93]. interpretation [AKJ97, BJ94]. Interquartile [Sho99]. intersection [KH96]. Interval [CH95, SO95, SSM98, All99, CC99, FR95, GPPL99, PSCY97, PCH97, TG93]. Intervals [Wan90a, Wan90b, Adk90, ATT97, CFBGM^{+}97, CT93, CFMM95, DS91, FT93, GS90a, Luc96, MJ97, MR96, Nis97, ON96, PB94, SAL92, SY93, SAS91, SF94, VC97, VC99, Wan91, Wan92, WMF91]. intraclass [All99]. intratwin [WZ98]. Invariant [SY93, FH98, Nag90]. inverse [AK91, Bog93, Don96, Had98, MN99, Ong98, Sho97]. inversion [THD98]. investigation [MX92]. involution [Kro97]. Involving [GS90b, JG95, MW94]. Issues [Gou97]. items [XGX97]. iterated [PSCY97]. Iterative [HW96, AMJM96, APN91].

jackknife [Adk90, BHG94, Sha91]. Jackknifing [HCNA97]. James [Adk90, Hil91, Hil93, Hil94]. Jenkins [LV99]. John [MT91]. Johnson [Cas98, CTH^{+}94, Swi94, TJ90, TJ94].

Kalman [RC95]. Kaplan [CK95b, FZL98]. kappa [All99, BRB95]. kernel [Rac93]. Kernels [LK90]. Kiefer [Bar94]. known [Dor93b, ET93]. Kolmogorov [KSS96, Mar98]. Kriging [BP97]. Kullback [ASA90, Swa92].

L [Lim95, Ekb90]. lack [Och93]. Lagrange [AK93b, CNF95, Dev92, SM91]. Laguerre [DG94]. lambda [App95, KDM96, KD99]. Laplacian [Hut90]. large [BP97, CP92, Hor98]. large-sample [CP92]. largest [BGDH90, CC99, SU92]. latent [YPW94]. LAV [DP90]. law [HB98, RB90]. Lawless [Fir91]. Layout [KC95, Che93]. LD50 [Hoe91]. LD50-estimators [Hoe91]. Least [HW96, LK90, BF92, BN98, Ct91, CG93, Cut93, SJ91, Sal95, Sal99, TDTC93, TA92]. Least-Squares [LK90]. left [BW97]. Leibler [ASA90, Swa92]. length [Gan94, GS90a, JC93]. lengths [ACM94, Haw92]. less [Wen96]. Letter [Aka97]. Level [GS90b, Adj91, GS92, LK97, MW94, Nag90]. levels [CR97]. life [AGAA98, BLN94, Hsi94, JY95, LW99, Lin99]. lifetimes [SKG96]. like [BT92, PC96]. Likelihood [BY90, MW94, Agr95, AMJM96, AY90, AWEBW92, AWBW^{+}96, CA98, jCM97, FW99, Gea91, Gor90, HB94, Hol91, Jai92, KTV94, KK97, MX92, MMM95, Pla98, RLR98, RRN94, SSW95, SA99, Taa97, Tay90, WCWB96, YSM94]. LIME [SJ91]. limit [Blo98]. limitations [HL93]. limited [Sho98]. Limiting [GS90b, GS92, RLR98]. Limits [LOSB90, Cor97, DW94, FW92, KSS96, OM90, RKN91a, Van94]. line [CW96, SSM98]. Linear [CPS97, DP90, KP90, Koz90, AC98, dCQAC97, BF92, BC92b, BLN94, BJ94, Boi96, CL94, Dha92, DS91, EDMAYS97, FT93, Gra93, GK97b, HK90b, Lav92, Leu94, cLH97, NH97, PB94, Pen95, RRN91, RKN91b, Ryc92, SlC97, Str97, SO94, Sut90, Tsu95, Wan91, WS93]. linear-exponential [EDMAYS97]. Liu [CPC97]. LM [Koz90]. Local [Lav92, Far92, Lim99]. Locally [Nag90]. Locating [Tör90]. Location [Don90, Wil90a, Abo92, BS93, Leu94, LY94, Won92]. locations [BT92]. log [Don96, Gea91, HB94, HK90b, Kim99, KJ98, Pla98]. log-logistic [Don96]. log-normal [Don96]. log-rank [HK90b, Kim99]. logistic [App95, BD97, BHG94, Don96, GH91, LK97, SO95, SF94]. logit [AC98]. lognormal [JC96]. long [OM95]. long-tailed [OM95]. longest [Sch96]. Longitudinal [BM96a, FR95, PW92, PS99, PA94]. look [Nic93]. loop [Luc97, SSW95]. loss [Mad97]. lots [GG97]. Lower [BCC97, LOSB90, FW92]. LQL [GK92]. LR [Koz90]. Lyapunov [Lim99].

M [Jai92]. M/G/1 [Jai92]. MANOVA [BN98, SKBK98]. MANOVA-type [BN98]. marginal [PA94, Rod99, TQK97]. Marginals [ORR90, LG99, Ong92]. Markov [GT99, MLA^{+}99, PR96, RLR98, THWL97]. matched [NS96]. Matching [Was90]. math [HN95]. mathematical [BHK91]. matrices [Dor93b, PS99, RR94, Sug93, Sug94]. Matrix [TI90, dCQAC97, BP97, Had98, Hut90, Mor99, Sal97, Sal99]. maxima [BG95]. Maximum [Agr95, BY90, CA98, Gor90, RLR98, Won92, AMJM96, AWEBW92, BG95, KK97, RRN94, SJ91, SH98, SG93b, SA99, Tay90, YSM94]. MCMC [Alb96]. Mean [Fro91, Pli90, BGDH90, CC99, Fan94, GPPL99, GB97, Hai96, KW99, KPA91, LNZ94, PSMR95, SJ91, Smi90, TG93, TYF98, TI90, Vid93, VC97, Wan91, Wee96, YSY96]. mean-squared [Hai96]. mean-squared-error [SJ91]. means [AW96, And91, AMH92, Ano94a, BC92a, BSB97, CGR99, HG98, Ken90, Mac94, PVEP97, Ste93, Wil90b, Zho92]. Measure [ASA90, Gra93]. measured [SC93a]. measurement [NGM98]. measurements [Gea91, HZBW91, KAKW98]. Measures [Wil90a, BuH97, EK93, HS97, JK98, LY94, NSS99, Rav94, Rod99, Sch94]. Measuring [GL97]. mechanisms [PC94]. median [Blo98, Cas98, DSW96, Gan94]. Meier [CK95b, FZL98]. meta [SWFK96]. meta-analysis [SWFK96]. metaanalysis [Zha94]. metamodels [SS95]. method [AH95, CW96, DG94, Dev92, Din97, Dor93b, Had98, Hil91, HD94, HK90b, KD99, KW96a, KR98, MABL90, ÖR92, PJ99, SS98, SR97b]. methodologies [YY96]. methodology [SE93a]. Methods [KP90, Mur98, And91, AMH92, BS94a, FW99, GT99, JY97, Kar93, PSCY97, Sam99, SO95, SlC97, SM98c, SK91, SKBK98, WAD^{+}92]. Metric [CV94]. Mewama [Ale99]. Meyer [GT90]. microcomputers [Bur90]. Midzuno [Bed95]. might [SA95]. minimal [Pur94]. Minimax [HAS98, Vid93, VI94]. minimization [AAA97]. minimizing [Blo98]. Minimum [GS90b, UN95, Dha92, PBB95, TPSS99]. Misclassification [Hou93, Dor93b]. Mises [Hol95, KSS96, RL96, SK91, Woo94]. missing [GH91, HD98]. Misspecified [RJ91, Fir91, GK97b]. mixed [Ong98, Wol93]. mixing [LM94, Phi90]. Mixture [Ame90, AWEBW92, AWBW^{+}96, CA98, jCM97, Czi92, JY97, MMM95, MR96, Par99, Phi98]. mixtures [Gor90, SK91]. ML [RKN91a]. MLES [BHG94]. mode [SH98]. Model [Gil90, Koz90, TQK97, Alb96, And94, APN91, BS99, BSB97, BJ94, BS93, Bro97, CH95, CPS97, Chr99, Ct91, CAF96, Dha92, DS91, DG96, DMS99, Dun92, Fro91, Fuk94, FN98, GK97a, GS94, Hal96, HS97, HOJ97, Kul96, LW96, LY94, MMM95, MF93, PB94, Pen95, Pur94, RK99, RLR98, RBL91, SC93b, ST93, SF94, Taa97, TB99, TW98, TA92, Wan92, WS93, dK96]. Modeling [LWC91, HK90b, Mur98, SW98a]. modelling [BRB95]. Models [Hun90, KO90, KP90, AC98, AMJM96, AK93b, AH95, AG98, dCQAC97, BLN94, BM96a, BG93, CD93, CP92, CCN98, CNBCF98, Fuk94, Gar94, GT99, GK97b, HD98, Her97, HZBW91, KG91, KW96b, Lav92, LU95, LV99, LM97, MLA^{+}99, MP97, PW92, PA94, Rav94, Ros93, Sal97, Sal99, SF93, SC92, SlC97, Tay90, Wan97, Wol93, YPW94, Zhe98]. Modification [Mar98, Bar94]. Modifications [KSS96, PK95, KJ99]. Modified [Gan90, GK92, ALM93, HKT93, JC96, O’G96, Pla98, Sea97]. moment [JC96, KDM96, TW98]. Moments [EDMY91, AH95, BKS92]. Monitoring [Gan90, HRM99, LCW95, Was93]. Monte [AH95, BF92, CGR99, DP90, Die92, EK93, Fuk94, GT99, Gil90, HCNA97, JLPR99, Phi98, Sam99, Sch93, Tan99, WAD^{+}92, ZB94]. Monte-Carlo [WAD^{+}92]. Morgenstern [SG92]. most [HEB93]. MOSUMSQ [Tör90]. Moving [Gan90, AS91, Dun92, DSW96, MP98, SAL92, SAS91]. MPS [SG93b]. multicollinearity [LW96]. multidimensional [NR95]. multigroup [SG93a]. multiletter [Sch96]. multimodal [Cer92, Hol95, Lye93]. multinomial [Hir97, MJ97, MABL90, Mor92, Mor99, WC90]. multinormal [Ken90]. multinormality [Sug93, Sug94]. multiparameter [HKT93]. Multiple [Ame90, Bri90, HS90, WC99, Bed95, DG94, DMS99, GL99, HEB93, HO94, KC96, KH96, NS96, OR95, RW96, SE93a, Sea97]. multiplicative [BS99, SC92, SC93b]. multiplier [AK93b, CNF95]. multiply [USS96]. multisymbol [SX97]. Multivariate [Ame90, HRM99, RR92, BM96a, BS99, BG93, Car98, CK95a, CL94, CR97, Dre90, HS97, HL93, KTV94, LML94, LG99, MRS91, NH97, ÖR92, PW92, RWB97, RS90, SA95, SGF^{+}91, SW98b, TH99, Vid93, dK96]. MVU [ET93].

N [LU95]. natural [HB98]. nature [FW92]. Nearly [DG90, CC99, NdL97]. negative [AY90, BKS92, Fam97, MST93]. neighbour [MN91]. nested [AH95, ME91, PB94, Sch93]. Newton [Din97]. Neyman [Zhe95]. noise [GD95]. noisy [LH97]. Non [KAS90, Aka95, And91, AH95, CL91, CP92, CCN98, DM90, GT99, HB98, Li96, PC94, PSCY97, RGU97, Sch93, SK93, Wil90b]. non-asymptotic [And91]. Non-Central [KAS90, Aka95]. non-exponential [CP92, CCN98]. non-Gaussian [GT99]. non-homogeneous [HB98]. non-identical [CL91, Wil90b]. non-iterated [PSCY97]. non-nested [AH95, Sch93]. non-normal [DM90]. non-parametric [RGU97]. non-random [PC94]. non-stationary [Li96]. non-white [SK93]. nonadditive [SC92, SC93b]. Noncentral [Gui90, CH94, CS95a, Din97, FWW98, KK91, KK99]. Noncentrality [Gui90, Din94, Din97]. nonhomogeneous [LWC91]. nonindependence [ZWZ93]. Nonlinear [Tan99, Chr99, CP92, FN98, GT99, HZBW91, LM97, RL90, SSM98, SS98, Str97]. nonlinearity [Rav94]. nonnormal [Tan99]. nonnormality [Fuk94]. Nonparametric [Che93, Don90, GR92, KC95, AS91, AWhW93, Bar96, Chu95, Dia99, HL92, LH97, MZ91, Rac93, SA99, SS98, Van94, VC97, ZWZ93]. nonstationary [NdL97]. norm [Gra93]. Normal [Ame90, AWEBW92, AWBW^{+}96, BG95, BSB97, BGDH90, Bog93, Cas98, CA98, CL91, CC99, jCM97, De 90, Don96, DM90, Dre90, Fan94, GB97, GS98, Has95, Hor98, LM94, LW94, Mac94, MMM95, ME94, ÖA94, Par99, Par92a, Par92b, Phi98, STY93, Sho97, Ste93, THD98, Vau90, Vid93, Wee96, WC99]. normality [Boi96, HL93, LL92, Men92, Nai96, ÖR92, SG93a]. normalized [GR92]. Normalizing [Viv90, JLPR99]. Note [Czi92, Gui90, AK93b, FW92, Mur99, Ong98, PA94, RGU97, SQ90, Shi99, Sut90, Tho95, Zhe95]. Null [KC95, jCM97]. number [BG93, Bur90, Far92, SX97, SH98, Wol92]. Numerical [Sal97, BS97, Mad97, Sal95, Sal99, THD98].

observation [Her97, SS96]. Observations [Gan90, Mag90, CS95b, DSW96, GL97, OR95, dOP99, Sch92, SK90, SW98b, ZWZ93]. obtained [Sam99]. OC [Raj92]. occurrence [MR96]. occurs [Gor90]. odds [CT93, Pla98]. omnibus [Zha94]. on-line [CW96]. One [Ale99, ORR90, OM90, RW96, Van94, Car98, Cha90, Che93, CH95, CNBCF98, GW98, HEB93, MMM95, MP98, Nis97, PB94, SS96, Wan92, YSM94]. one-fold [PB94]. One-Parameter [ORR90, CNBCF98]. One-sided [Ale99, OM90, RW96, Van94, Cha90, MP98, Nis97]. one-way [Che93, CH95, GW98, Wan92, YSM94]. operating [Luc97]. operational [Fir91]. Optimal [ATL92, EHAA94, Lee92, RBL91, TB99, YC93, AWhW93, Boi96, Bro97, CC99, Gan94, HK93, LT95, MY98, Nag90, Ngu93, RF96, RRN94]. Optimality [Ct90]. optimally [KPA91]. Optimization [Kar93, Con92]. Optimum [Kha97, RRN91, RKN91a, RKN91b, SO95]. Order [BC92a, BC92b, EDMAYS97, ALM93, AC98, Aka95, CT99, EDMY91, HN95, HK93, HAS98, ON96, Par92a, Par92b, RRN91, RKN91b, SC93a, TA92, Vau92, WS93]. order-restricted [AC98]. Ordered [Mag90, ALM93, JK98, LU95, Sug93, Sug94, Tho95]. ordering [SSW95]. orders [SA95]. ordinal [SW98a, SKBK98, Tay90]. ordinary [TDTC93]. Orientation [Pli90]. orthogonal [GT90]. other [BC92b, HOJ97]. outcomes [SlC97]. outlier [Car98, Fro91, FN98, Mur99]. outliers [MRS91, RDW96, SGF^{+}91]. outlying [CT92]. outlyingness [AR90]. overdifferencing [NdL97]. overestimation [BM96b].

pair [dOP99]. paired [Cro90, EHAA94]. pairwise [HO94, Ken90, Sea97]. panel [LM97]. papers [Luc94]. Parabolic [Wik97, DW94, EDMY91]. parallel [Wol92]. parallelism [NSS99]. Parameter [BDP90, Fam97, Gui90, Luc97, ORR90, Per90, Tör90, ATT97, AG98, CC90, CNBCF98, Din97, ET93, Jai92, MP97, Nar91, RRN91, RKN91a, RKN91b, SF94, TPSS99]. Parameterized [Sun90]. Parameters [BY90, Mag90, AGAA98, Agr95, And94, APN91, CGR99, Din94, DN97, JC96, Par99, RG96a, RRN94, STY93, SK91, SR97b, Tay90, TP90, USS96]. parametric [CP92, Lye93, RGU97, ZWZ93]. Pareto [AGAA98]. Part [JR96, RM96]. partial [SH93]. Partially [Phi98]. passage [GS99]. pattern [Che93]. PCA [AGV96]. pdfs [SH98]. peak [FH98]. Pearson [BS94b, Don96, WCWB96]. Pearson-3 [Don96]. penalized [HB94, PHB97]. Percentage [CL91, Aka95, BCC97, BS94b, BS97, Hor98]. percentile [Dun94, Yeh96]. percentiles [CTH^{+}94, KD99, MN99]. perfect [Wen96].

Performance [Hol91, SA99, Was90, Adk90, GW98, GS90a, LCW95, Nis97, PKD99, PC97, SM98c]. performance-based [SM98c]. performances [BGDH90]. permutation [HG98]. perspective [BS97]. Plan [GS90b, GS92, GL99, Lin99, NGM98, SA90]. Plans [Was90, GK91, Hsi94, JY95, JR96, Kan93, KG93, Raj92, RM96, SR97a, YC93]. play [Ata99]. play-the-winner [Ata99]. plot [Rau92]. plots [JD95, KUK96, MH98]. plotting [Luc94]. point [Aka95, AS95, Bro97, DG96, FR95, SJ91, SO95, TG93]. points [BD93, BCC97, Blo98, BS94b, BS97, CL91, Dun94, Hor98, SE93a, SO94, TB99]. Poisson [FL92, Gan90, HB98, LWC91, Ong98, SM91, Sil97, YMS93]. Poisson-inverse [Ong98]. policy [AWhW93]. polygons [Sim95]. Polynomial [LK90, NM98]. polytomous [dK96]. population [BSB97, BGDH90, BT92, Lee96, PSMR95, San95, TYF98, YSY96, YY96]. Populations [Ame90, AK91, BGDH90, ENN92, GB97, LM94, RR92, RK96, ŠŠ92, TP90, VC99, Wee96, WC99]. portable [Wol92]. portmanteau [BG93, KW96b]. position [Sim95]. positive [Sil97, THD98, Wan91]. positively [PCH97]. possible [SK93]. posterior [MABL90]. Power [CP93, CMY99, HO94, KH96, KG90, SU92, Viv90, Agr95, CR97, DM95, DM93, HB98, Hun98, NdL97, O’G95, Ogd97, RG96a, RB90, SM98a, SA91, Sut90, ZB94]. powerful [SF93]. powers [Boi96, SSW95]. Practical [KS98, RC95]. precision [BRB95, Bog93, Dun94]. preconditioning [Ogd97]. predict1on [GL97]. Prediction [Wan92, BS93, CFBGM^{+}97, CAF96, ON96]. Predictive [HD98, Lav92, PHB97]. preliminary [CH95, SW98b]. prescribed [Fal99]. presence [ACM94, CG93, GD95, LM94, PCH97, SL97, SK90, US92, VR97]. Pretest [KTTB92, Adj90, Shi99]. pretesting [Adj91]. prices [DE99]. principal [Wal90]. prior [Fan94, Gar94, HB98, Mac94, PM97]. priors [Lav92]. probabilities [CT93, Dor93b, Hou93, Lee96, MW94, MN99, PM97, SY93, SJNM91]. probability [BM96b, CH94, De 90, DN97, GR92, HEB93, JD95, Luc94, MABL90, Rau92, RR99, RL95, SK92]. Problem [Wil90a, BHK91, CR97, GK97b, LML94, Meh97, TH99, WC90, YMS93]. problems [Ekb90, GS94]. Procedure [Don90, APN91, Che94, CPC97, CW96, Con92, Dra97, HEB93, KH96, Sal95, Sea97, SAS91, SA90, Was93]. Procedures [Bri90, Kan93, AMJM96, BGDH90, Bla91, GB90, WC99]. Process [JW90, LOSB90, AGV96, CP93, CW96, Czi92, Dev92, Had98, HRM99, HB98, KL90, KJ99, LWC91, LCW95, Luc96, Mac94, PC97, RB90, WMF91, Was93, YB98, YMS93]. processes [CFBGM^{+}97, DE99, GS99, HM92, KS98]. product [ME94, SG93b]. profile [NSS99, Pla98]. program [WMF91]. Projection [Kro97, YY96]. Properties [Koz90, MJ97, MRS91, NSS99, SGF^{+}91, Tay90, YSM94, Die92, DM90, Fir91, Has95, MP98, Nar91, San95, SA91]. proportional [Luc97]. proportional-integral [Luc97]. proportionality [Bér92]. proportions [And91, AMH92, DM93, Jai91, LM94, MJ97, SK93]. proposals [MS98]. protected [KH96]. protein [NS96]. Publication [Smi97b]. pursuit [Kro97]. PWM [Don96].

QSS [GK92]. QSS-1 [GK92]. quadratic [CH94, Hou93, SJ91, SO95, THD98]. qualitative [Sim95]. Quality [GS90b, KJ98, Cas98, CGR99, GS92, HL92, NGM98, SA94]. Quandt [Adj90, Adj91]. Quant1les [MTT90]. Quantile [KUK96, DLP94, MH98, RL90]. quantiles [GS90a, MP97]. quantitative [Sim95]. quasi [Taa97]. quasi-likelihood [Taa97]. queueing [Jai92]. Quick [AS95, GK92]. quotient [BCC97, ÖA94].

Random [Bur90, LW93, MF93, Ano90, BD93, BK91, CL91, Che94, CH95, Con92, Dev92, Die92, DE99, Fal99, HZBW91, HD94, ME94, Mor92, PC94, TB99, Wan92, Wol92]. Random-effects [MF93, Wan92]. Randomization [KC96]. Randomized [ATT97, Ano94a, Ata99, SD96]. Range [LOSB90, CL91, God94]. Ranges [Wan90a, Wan90b]. ranging [MF93]. rank [Bet93, Bla91, BT92, EK93, HK90b, Hun98, JK98, Kim99, LML94, O’G97, SF93, Tho95]. rank-like [BT92]. rank-sum [Bla91, O’G97]. ranked [Bar98, KA99, Sam99, TYF98]. ranked-set [Bar98]. Ranking [Cro90, KE96]. rankings [LT95]. ranks [Ano94a, Mil91]. Rao [RL95]. rare [ALN95]. rate [RFL91]. rates [CR97, Dor93a, Ren92]. ratio [AWBW^{+}96, BSG98, jCM97, CT93, FW99, Gea91, Hol91, HD94, HL90, MX92, MMM95, MW94, Pla98, SSW95, Tay90, WCWB96]. ratios [KG91, San95]. re [Gou97]. re-estimation [Gou97]. Read [Par99]. readers [OR95]. recapture [Lee96]. reduced [Jai92]. reduction [BHG94, CCN98, CNBCF98, JC93, PSCY97, RL90]. Reference [KS96, Wen96]. region [HAS98]. regions [HK93, NH97]. Regression [HK90a, KP90, Pli90, dCQAC97, Bar96, BA93, BHG94, Cer92, CPS97, Chr99, Chu95, Ct91, CCN98, CFMM95, Dia99, Die92, DS91, DG96, DMS99, Ekb90, FT93, Fir91, Fül95, FN98, Gar94, GH91, Gra93, Hal96, HRM99, HC97, JY97, KC96, LEKR97, LK97, cLH97, Mur98, NM98, Och93, OM90, PB94, PA94, Pla98, Rac93, RR94, RGU97, Sil97, SM98c, Str97, TQK97, Tay90, Wal90, Wil98]. regression-adjusted [HRM99]. regressions [CA98, Ryc92]. regressors [SA91]. rejection [HD94, RS90]. rejective [Sea97]. relation [BD97]. relations [KJ99]. relationship [Pen95]. Relative [Far92, Jai92, Rav94]. reliability [AGAA98, AK93a, AAA97, CC98b, GS98, SKG96]. reliable [Luc96]. Renewal [JW90]. repair [Pur94]. Repeated [Gea91, AV92, BuH97, EK93, HS97, HZBW91, KAKW98, NSS99, Sch94, SKBK98]. replacement [AK93a, AWhW93]. replicate [RLR98]. replicated [SK90]. requirements [All99, ALN95]. Resampling [Chr99, LY94, THWL97]. research [MLA^{+}99]. residual [BG93]. residuals [Kul96]. resolvable [Ngu93]. respect [FZL98]. Response [ST93, AC98, APN91, Ata99, HAS98, KUK96, McC96, PJ99, PA94, SJ91, SE93a, TG93, Tay90]. responses [JK98, Mor99]. restricted [AC98, GK97b, Tsu95]. restriction [LW94]. resubmitted [GG97]. resultant [SMR95]. results [GK97b, Wil98]. review [And91, AMH92]. revision [HEB93]. revisited [Ata99]. Ridge [Edl90, HK90a, Pli90, CFMM95, DM90, Fan99, Fir91, Fül95, JY97]. Right [Mag90, BW97, Cha90, Che94]. right-censored [Cha90]. right-censorship [Che94]. Rights [SK93]. risk [Mad97]. Risks [GS90b, PTP92]. Robust [AR90, BSB97, BA93, cLH97, Mag90, PBB95, PHB97, WAD^{+}92, Abo92, Bér92, GD95, RC95, SF93, ST93, SM98c, Smi90, Str97, YB98]. Robustness [Ame90, Men92, MST93, SD96, SL97, SA91]. role [Zhe98]. root [SU92, TW98]. Rosenblatt [Bar94]. Rotation [FH98]. rounded [dOP99]. rounding [Ros93, TO97]. rule [Ata99, Men92]. rules [ACM94, Dun94, Vid93, VI94]. run [ACM94, Gan94, GL99, Haw92, JC93, Sch96, WS93]. runs [ACM94, SX97, SH98]. RV [HG98].

s [SH98]. Saddlepoint [Pen95, HN95]. Sample [All99, ALN95, KC95, Koz90, Wil90a, Abo92, Adk90, AK93b, Car98, Cas98, CT93, CP92, Die92, DM90, DM93, FT93, Gou97, GGC95, Hai96, Has95, Jai91, KW96b, LML94, Lee92, MZ91, dFMK92, Nar91, O’G95, O’G96, O’G97, Ogd97, PR94, RWB97, RW96, RG96b, San95, SY93, Sch96, SX97, SG92, SKBK98, TYF98, TO97, YPW94, ZWZ93, dK96]. sampler [GB97]. samples [And91, AMH92, BK91, CT99, ENN92, FR95, FR93, Gea91, PVEP97, RF96, Sam99, SG93a, SG92, TP90, Vau90]. Sampling [Ame90, GS90b, GG97, Was90, ALN95, AS95, Bar98, GK91, GS92, GL99, GS94, HM92, Hsi94, JR96, KA99, LW93, Lin99, Nis97, PCH97, Raj92, RM96, SAL92, SAS91, SA90, SR97a, Tan99, YSY96, YC93]. Satterthwaite [DM95, DM93]. Saving [CFBGM^{+}97]. Scale [Per90, BT92, Dun94, RRN91, RKN91b, RG96b, TP90, Won92, WC99]. schema [Fan99]. scheme [AS91, Bed95]. schemes [GPPL99, JC93, LY94, SAL92, SA94]. scores [HK90b, O’G96]. Screening [NS96, Bet95]. search [Con92, DMS99, Was93]. seasonal [BS99, Dun92]. second [HK93, WS93]. second-order [HK93]. Section [SK93]. seemingly [Ct91]. segments [NS96]. select [Bro97]. selecting [BGDH90, KAKW98]. Selection [GS90b, GS92, HS90, KG93, KG90, JLPR99, AKJ97, DMS99, Fuk94, GK92, GS94, HOJ97, Kan93, Men92, Mur98, Rac93, RK99, RFL91, ŠŠ92, SK92, SA90, Wol93]. selector [Chu95]. self [AG98]. self-exciting [AG98]. Semiparametric [SG99]. Sen [Wil98]. Sensitivity [FR95, TYF98]. sequence [Bro97, DSW96, WC90]. sequences [NS96, Phi90]. Sequential [AWhW93, Dia99, Bet95, Kim94, ŠŠ92, TO97]. sequentially [Sea97]. sequentially-rejective [Sea97]. Serial [Ano90, KP90, SL97]. serie [RDW96]. series [AV92, AK93b, BS99, Dun92, Kar93, KW96b, LNZ94, LV99, Rav94, RJ91]. set [Bar98, KA99, Sam99, TYF98]. sets [BP97, RR94]. seven [CR97]. Several [DG90, Che94, KR98, O’G95, WW98]. severity [Adj90]. Shaffer [Sea97]. shape [CC90, ET93, SH98]. shaped [FH98]. shapes [Wil90b]. Shapiro [Rah99]. Shenton [BS94b]. Shewhart [PTP92, PCH97]. Shift [ALS95, VC99]. Shifted [SC92, ET93, SC93b]. Short [WS93]. Shrinkage [RDW96, SKG96, HKT93]. sided [Ale99, Bri93, Cha90, MP98, Nis97, OM90, RW96, Van94]. sign [Bar98]. signal [GD95]. signals [KK97]. significance [Adj91, LK97]. similarity [LY94]. simple [BD97, BK91, Cha90, DS91, Fal99, Had98, LGG97, MABL90, Mor92, PB94, VI94]. Simulated [ASA90, Bér92, AAA97, Li96, Sam99]. Simulating [Dun94, JC93]. Simulation [Gil90, Her97, HW96, Hol95, Hun90, LG99, Wil98, Woo94, AR90, AWhW93, BDP90, CT92, CPS97, CGR99, CFMM95, Die92, Dun92, GS90a, GS99, LWC91, Lim99, MST93, dOP99, PSCY97, SO95, SS95, SKBK98, TG93, TDTC93]. Simulations [AMJM96, MS98, RL90]. Simultaneous [YY96, MJ97, OM90]. Single [CC98a, GS90b, Was90, Fro91, FN98, GK91, HS97, SJ91, YC93]. single-factor [HS97, SJ91]. Single-stage [CC98a]. singular [De 90]. singularities [CA98]. six [BF92]. Size [GW98, All99, ALN95, Gou97, Jai91, Lee96, Ogd97, PR94, SKBK98, TO97]. sizes [Kim94, Lee92]. skew [Don96]. skewed [EDMY91, LS98, PC97, Sho99]. skewness [Hai96, HL93]. slippage [SSW95]. slope [HAS98, KUK96, SF94]. Small [Adk90, CT93, Die92, DM90, DM93, FR93, FT93, GGC95, KC95, Koz90, MZ91, Nar91, San95, YPW94, FR95, Gea91, Has95, RF96, RWB97, SY93, SG93a, SKBK98]. Small-sample [DM93, San95]. smaller [Hai96]. smallest [DS91]. Smirnov [Mar98, RG96b]. Smith [Hil94]. smooth [cLH97]. smoothing [Hut90, Sim95, Wan97]. smoothness [Chu95]. software [BHK91, Str97]. solution [Meh97]. solutions [CR97]. Some [Abo92, ALM93, CAF96, GK97b, Ros93, SM91, Str97, VI94, BC92b, BN98, CP93, GW98, HCNA97, HK93, KW99, KTTB92, KW96b, LL92, MZ91, PK95, Phi98, RR94, Shi99, ZWZ93]. sources [SA95]. space [GT99]. spacing [RL95]. spagings [SG93b]. sparse [BP97, PKD99, SWFK96]. Spatial [Sch92, CG93, Taa97]. SPC [SRHW99, WS93]. Spearman [AC97, SQ90]. specific [Her97]. Specified [ORR90]. spectra [Li96]. spectral [RF96]. Speed [Luc94]. spline [Wan97]. Splined [Och93]. splines [AGV96, Dia99, GR92, Hut90]. Square [KAS90, Fro91, GB90, KK99, TI90]. Squared [KG90, Pli90, ATL92, Din94, FWW98, GPPL99, Hai96, Mil91, SJ91, Tho95, Wan94]. Squares [HW96, LK90, BF92, Ct91, CG93, Cut93, LGG97, Sal95, Sal99, TDTC93, TA92, Wan91]. Stable [BY90, Per90, NR95, Pan92]. stage [CC98a, HKT93, JR96, dFMK92, MY98, RM96, Zhe98, dK96]. stagewise [KH96]. Stahel [Mur99]. Standard [KO90, DG96, God94, PK95]. standardized [Pan92]. starting [CA98]. state [GT99, THWL97]. state-space [GT99]. stationary [Had98, Li96, TA92]. Statistic [Rah99, Tör90, All99, AWBW^{+}96, Bar98, BCC97, CL91, CMY99, GB90, KL90, KE96, MTT90]. statistical [BJ94, CC98b, HRM99, LH97, MH98, PKD99, SM91, ZWZ93]. statistics [Abo92, BC92a, BC92b, BJ94, CT99, EDMY91, EDMAYS97, HN95, Lye93, Mag91, ON96, PVEP97, Par92a, Par92b, RRN91, RKN91b, Vau90, Vau92]. steel [Che94]. Stein [Adk90, Mad97, Ste93]. step [BHG94, Kha97]. step-stress [Kha97]. stepwise [LK97, Ren92, RFL91]. stochastic [AGV96, Hut90, MP98]. Stock [DE99]. stopping [Dun94]. straight [SSM98]. Strategies [ŠŠ92, Sof99]. strategy [CK95a, MY98]. stratification [Cer92, UN95]. stratified [dFMK92]. Strawderman [SE93a]. strength [BDP90]. stress [Kha97]. strictly [Mit93]. structural [HOJ97, PP98, Pen95]. structure [KJ98, PB94, TI90, Wol93, dK96]. structures [De 95, KAKW98]. Student [BD97, Vau92]. Studentized [CL91, Nai96]. Study [ACM94, Gil90, HW96, Hun90, AR90, AWhW93, AH95, BA93, BDP90, CT92, CPS97, CFMM95, DW94, Dun92, EK93, Fuk94, GS90a, Gar94, HS97, Hun98, KW96b, LGG97, Lim99, MST93, PSCY97, SW98a, SH93, TDTC93, XGX97, YY96]. style [Mac94]. subject [AAA97, Ano90]. Subset [HS90, RR92]. success [CT93, SY93]. successive [YSY96]. suddenly [And94]. suggested [FW92]. suited [OM95]. sum [BG95, Bla91, Gan94, Haw92, KS98, O’G97, Tsu95]. sums [BG95, LGG97]. superimposed [KK97]. superiority [GK97b]. superpopulation [BS93]. supplement [Sug94]. supplementary [ACM94, PSMR95]. supplemented [Hoe91]. surface [HAS98, KUK96, PJ99, SE93a]. surveillance [FW99]. surveys [AV92]. survival [HK90b, KT92, KW96a, MLA^{+}99, Zhe98]. switch [DG96]. switching [AS95, DG96, GK92]. symbolic [HN95]. Symmetric [BY90, LK90, Per90, BD97, Pan92, RKN91b, SR97a, TW98, VC97, Won92]. symmetrical [Cas98]. synergy [WW98]. System [ORR90, AAA97, AS95, BHK91, GK92, Jai92, KDM96].

T [KG93]. table [RL95, SC93b]. Tables [God94, HPL95, Hor98, SE93a, Sea97, TJ90, TJ94, ATL92, BC92a, FW92, GK92, Kan93, KDM96, LU95, PKD99, SC92]. Tabu [DMS99]. Tadikamalla [TJ90, TJ94]. Taguchi [CW96]. Tail [CH94]. tailed [OM95]. TAR [Lim95]. technique [GS99, HL92, TPSS99]. techniques [BP97, Dor93a, JC93, SM98b]. terminals [TJ94]. Test [Don90, Tör90, Abo92, ATL92, AWBW^{+}96, Bar98, BG93, BT92, Car98, Cha90, CPC97, CH95, jCM97, CNF95, DM95, Ekb90, ENN92, HG98, HK90b, Hsi94, Hun98, JY95, KH92, KW99, Kim94, Kim99, KW96a, LT95, LML94, Lee92, Lin99, Lye93, Mag91, Mar98, MX92, MMM95, Mil91, O’G96, O’G97, PP98, RWB97, Rau92, RG96b, RL95, SF93, SM98a, Sha91, Sut90, Tsu95, Vau90, WCWB96, Wen96, Zho92]. testimating [CC90]. Testing [AW96, BuH97, LS98, LW94, ME91, PVEP97, Pur94, RR94, SO94, TW98, Wee96, AGAA98, AC97, Fuk94, FN98, KE96, KSS96, KR98, MZ91, MP97, Nai96, OR95, Och93, RBL91, RW96, SC93b, Sut90, Tsu95, TPSS99]. Tests [ABL90, DP90, KC95, KG90, Koz90, Per90, RR99, ALM93, AK93b, AH95, BLN94, BN98, BSB97, Bro97, CP93, CS95b, CP92, Din94, DM93, EK93, FH98, GW98, Gea91, HS97, Hir97, Hol91, KC96, KSS96, KW96b, LL92, LW99, MZ91, MRS91, MW94, Nag90, NSS99, NM98, NdL97, O’G95, OR95, PKD99, RJ91, RL96, SD96, Sch94, Sch93, SA91, SH93, Sho99, SL97, SSW95, SGF^{+}91, Swa92, Tho95, TO97, WZ98, WW98, Zha94, ZWZ93]. Testtesting [TP90]. Theil [Wil98]. THEIR [KJ99, BCC97, GB90, Luc94, MN91]. theoretical [Sal99]. theory [KDM96]. things [KW99]. third [HAS98]. third-order [HAS98]. Three [Hil93, Her97, PKD99, RBL91, SS96, TW98]. three-factor [SS96]. threshold [AG98, BM96b]. ties [JK98]. Time [AV92, AK93a, AK93b, BS99, BM96b, CFBGM^{+}97, DS91, Dun92, Kar93, KW96a, KW96b, LNZ94, Lee96, LW96, LV99, Luc97, Rav94, RDW96, RB90, TB99, YMS93]. time-dependent [YMS93]. times [GS99, MLA^{+}99]. tolerance [Hoe91, OM90, Van94]. tolerance-distribution [Hoe91]. tolerances [Boy94]. tool [SS96]. total [KW96a]. toxicity [MF93]. trace [Hut90, MS98]. trace-driven [MS98]. tradeoff [PM97]. traditional [Adj90]. training [LM94]. transfer [Edl90]. Transform [HS90, Hun98, Ogd97]. Transformation [Viv90, Hal96, KJ98, MTvdM98, Rah99, Zhe97]. transformations [SG92]. transformed [HD94]. Transforming [Boi96]. Transition [MLA^{+}99]. treatment [Che93]. Treatments [KC95, Che94, MZ91]. tree [Cha90]. trend [AC97, MW94, SO94, WS93]. trended [SA91]. trends [CL94]. trial [Pie95]. trials [MF93]. Trimmed [VC99, God94, KPA91]. trinity [Zhe98]. trinomial [ŠŠ92]. true [SK92]. truncated [BW97, EDMAYS97, MMM95, PVEP97, RB90]. Tukey [HEB93, MT91]. twin [WZ98]. Two [Ame90, Bri90, BHG94, JR96, KC95, MP98, RM96, Wil90a, Abo92, ALS95, AW96, And91, AMH92, AWEBW92, AWBW^{+}96, BS94a, BSB97, Bri93, jCM97, CT93, Dor93a, Gor90, HKT93, Has95, HPL95, Jai91, KG91, KAKW98, LM94, LML94, Lee92, dFMK92, McC96, MMM95, ME94, MRS91, MY98, MW94, MP97, O’G95, O’G96, O’G97, Ogd97, ÖA94, PR96, RW96, RKN91a, RG96b, SY93, SC92, SC93b, SE93b, Sho97, ŠŠ92, SGF^{+}91, Vau92, VC99, Wee96, Zhe98, dK96]. two-dimensional [PR96]. two-factor [MW94]. two-parameter [MP97, RKN91a]. Two-Sample [Wil90a, Abo92, LML94, Lee92, O’G95, O’G96, O’G97, RW96, RG96b]. two-sided [Bri93]. Two-stage [RM96, dFMK92, MY98, Zhe98, dK96]. Two-step [BHG94]. Two-Way [KC95, ALS95, KG91, SC92, SC93b]. Type [KO90, Zhe95, ATT97, BN98, Dha92, Hol95, KR98, RG96b, SH98, BS94b, CR97, CC98b, HEB93, JY95, LGG97, TP90, USS96, Vau90]. Type-I [JY95].

umbrella [Che93]. unbalanced [KG91, LGG97]. unbiased [BC92b, RRN91, RKN91b]. unconditional [SX97, SM98a]. uncorrected [DM93]. underdifferencing [NdL97]. unequal [BT92, HO94, Kim94, O’G95]. unified [Dor93b]. Uniform [DG90, MR96, App95, KH92]. uniformity [FH98, KE96, KSS96]. Uniformly [DG90, Fal99]. uniforms [HD94]. unit [TW98]. univariate [AK93b, HL93, KW96b, LM94, Smi90]. unknown [dCQAC97, BGDH90, DG96, KSS96, RR92, Sho98]. unrelated [Ct91]. unreplicated [FR93]. untrimmed [God94]. unusual [NS96]. Updates [BJ94]. Updating [AK91]. upon [SW98a, WS93]. Upper [De 90]. Use [Mil91, AC97, LGG97, PS99, PM97, SE93a, SWFK96, Smi90]. used [Sho97, TG93, YY96]. uses [BC92b]. usin [AGAA98]. Using [Bro97, Cer92, Gan90, HN95, HS90, Sch94, AC98, AGV96, AW96, Ano94a, AWEBW92, BP97, Cas98, CTH^{+}94, Din97, GR92, GB97, JC93, LV99, cLH97, PSMR95, Pla98, Sam99, Sea97, SS95, Swa92, Tan99, TW98, YSY96, Zha94].

Validation [LK90, De 95, MS98, Rac93]. validity [YPW94]. value [BC92a, BC92b, DS91, Don96, Luc94, US92]. Values [HS90, AW96, Bla91, Bri93, CA98, FN98, HPL95, Hor98, KS96, Pan92, Par92a, RWB97, RRN91, RKN91b, RL95, Vau92]. variability [CPC97]. Variable [AKJ97, CFMM95, Mur98, Nis97, PR94, PCH97, RFL91, SAL92, SAS91, YC93, YPW94]. Variables [Ct90, DG90, ALM93, And94, BG95, BS93, CL91, CD93, Czi92, Fal99, HM92, HRM99, HD94, Kha97, Lin99, ME94, ÖA94, Ong92, RS90, Sof99, SR97a, THD98]. Variance [Hil93, HK90a, Hun90, PP98, RL90, Wan90a, Wan90b, BGDH90, CC98a, CH95, Hil94, HO94, JC93, KG91, KUK96, LS98, ME91, PM97, RR92, RWB97, Sal99, SD96, TPSS99, UN95, WZ98, YSM94]. variance-covariance [Sal99]. variances [BC92a, CS95b, Hoe91, HL90, KR98, MABL90, O’G95, Par92b, Sha91, Vau92, Wil90b, Won92]. variate [Dev92, Hor98]. Variates [DG90, BSG98, SS95, TG93]. Variation [Tör90, GS98, LCW95, Mil91, Mor99]. various [SK91]. Varma [Sal99]. vary [Lee96]. varying [Bro97, Luc97]. Vector [Hun90, Sal97]. vectors [Mor92]. versus [MMM95, PCH97, Pur94]. very [OM95]. VI [BS94b]. via [BLN94, BM96a, Dia99, DN97, Hal96, Hil91, Phi98, Rah99, RL90, Rod99, SO95]. violated [MX92]. violations [SK93]. volume [Ano91a]. vote [Zha94]. voters [SK93]. voting [SK93, SK93]. vs [MZ91].

W [MT91]. Wald [Koz90, RWB97]. Wang [Fir91]. Watson [KSS96, LW93]. wavelet [Ogd97]. Way [KC95, ALS95, Che93, CH95, GW98, KG91, McC96, RBL91, SC92, SC93b, Wan92, YSM94]. Weibull [BLN94, CC90, EDMY91, JC96, McC96, RRN91, RRN94, Wol96]. weighing [Nic93]. weight [PHB97]. Weighted [Gan90, HW96, AS91, DSW96, LT95, LEKR97, MP98, SAL92, SAS91, Tsu95]. weighting [Nai96]. weights [Edl90, JLPR99]. Welch [Lee92]. white [PP98, SK93]. width [Rac93]. Wilcoxon [KL90, O’G96]. Wilk [Rah99]. Wilks [Car98, GK97a]. Wilson [MTvdM98]. window [Rac93]. winner [Ata99]. without [SK90]. WLS [SKBK98]. word [KDM96]. working [PS99]. works [MT91]. wrapped [Hol95].

x [CS95a, Tsu95]. XII [SG93b].

yield [Pie95].

Zellner [Has95].

References

Angus:1997:SAA

[AAA97] John E. Angus, Kevin Ames, and Kevin Ames. A simulated annealing algorithm for system cost minimization subject to reliability constraints. Communications in Statistics: Simulation and Computation, 26(2):783–790, 1997. CODEN CSSCDB. ISSN 0361-0918.

Adler:1990:TCB

[ABL90] Robert J. Adler, Lawrence D. Brown, and Kun-Liang Lu. Tests and confidence bands for bivariate cumulative distribution functions. Communications in Statistics: Simulation and Computation, 19(1):25–36, 1990. CODEN CSSCDB. ISSN 0361-0918.

Aboukalam:1992:SRT

[Abo92] M. A. F. Aboukalam. Some robust two-sample test statistics based on m-estimators of location. Communications in Statistics: Simulation and Computation, 21(1):133–148, 1992. CO-DEN CSSCDB. ISSN 0361-0918.

Alvo:1997:USF

[AC97] Mayer Alvo and Martin Charbonneau. The use of Spearman’s footrule in testing for trend when the data are incomplete. Communications in Statistics: Simulation and Computation, 26(1): 193–213, 1997. CODEN CSSCDB. ISSN 0361-0918.

Agresti:1998:ECI

[AC98] Alan Agresti and Brent A. Coull. An empirical comparison of inference using order-restricted and linear logit models for a binary response. Communications in Statistics: Simulation and Computation, 27(1):147–166, 1998. CODEN CSSCDB. ISSN 0361-0918.

Alwan:1994:SAR

[ACM94] Layth C. Alwan, Charles W. Champ, and Hazem D. Maragah. Study of average run lengths for supplementary runs rules in the presence of autocorrelation. Communications in Statistics: Simulation and Computation, 23(2):373–391, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Adjibolosoo:1990:DSH

[Adj90] Senyo B. S. K. Adjibolosoo. The degree of severity of heteroskedasticity and the traditional Goldfeld and Quandt pretest estimator. Communications in Statistics: Simulation and Computation, 19(3):827–836, 1990. CODEN CSSCDB. ISSN 0361-0918.

Adjibolosoo:1991:CLS

[Adj91] Senyo B. S. K. Adjibolosoo. On choosing the level of significance for the Goldfeld and Quandt heteroskedasticity pretesting. Communications in Statistics: Simulation and Computation, 20(2–3): 437–447, 1991. CODEN CSSCDB. ISSN 0361-0918.

Adkins:1990:SSP

[Adk90] Lee C. Adkins. Small sample performance of jackknife confidence intervals for the James–Stein estimator. Communications in Statistics: Simulation and Computation, 19(2):401–418, 1990. CODEN CSSCDB. ISSN 0361-0918.

Arnold:1998:APE

[AG98] Matthias Arnold and Roland Günther. Adaptive parameter estimation in self-exciting threshold autoregressive models. Communications in Statistics: Simulation and Computation, 27(4): 921–936, 1998. CODEN CSSCDB. ISSN 0361-0918.

Abdel-Ghaly:1998:EPP

[AGAA98] A. A. Abdel-Ghaly, A. F. Attia, and H. M. Aly. Estimation of the parameters of Pareto distribution and the reliability function usin accelerated life testing with censoring. Communications in Statistics: Simulation and Computation, 27(2):469–484, 1998. CODEN CSSCDB. ISSN 0361-0918.

Agro:1995:MLE

[Agr95] G. Agrò. Maximum likelihood estimation for the exponential power function parameters. Communications in Statistics: Simulation and Computation, 24(2):523–536, 1995. CODEN CSS-CDB. ISSN 0361-0918.

Aguilera:1996:AEP

[AGV96] Ana M. Aguilera, Ramón Gutiérrez, and Mariano J. Valderrama. Approximation of estimators in the PCA of a stochastic process using B-splines. Communications in Statistics: Simulation and Computation, 25(3):671–690, 1996. CODEN CSSCDB. ISSN 0361-0918.

Arkonac:1995:MCS

[AH95] Seyhan Z. Arkonac and Matthew L. Higgins. A Monte Carlo study of tests for non-nested models estimated by generalized method of moments. Communications in Statistics: Simulation and Computation, 24(3):745–763, 1995. CODEN CSSCDB. ISSN 0361-0918.

Amoh:1991:UDF

[AK91] R. K. Amoh and K. Kocherlakota. Updating discriminant functions estimated from inverse Gaussian populations. Communications in Statistics: Simulation and Computation, 20(2–3):619– 637, 1991. CODEN CSSCDB. ISSN 0361-0918.

Abedin:1993:CBE

[AK93a] Zairial Abedin and Marvin Karson. Classical and Bayesian estimation of exponential reliability under time censoring with replacement. Communications in Statistics: Simulation and Computation, 22(2):471–496, 1993. CODEN CSSCDB. ISSN 0361-0918.

Andy:1993:NFS

[AK93b] C. Andy and C. Kwan. A note on the finite-sample distribution of Lagrange multiplier tests for univariate time series models. Communications in Statistics: Simulation and Computation, 22 (4):1135–1160, 1993. CODEN CSSCDB. ISSN 0361-0918.

Akahira:1995:HOA

[Aka95] Masafumi Akahira. A higher order approximation to a percentage point of the non-central t-distribution. Communications in Statistics: Simulation and Computation, 24(3):595–605, 1995. CODEN CSSCDB. ISSN 0361-0918.

Akahira:1997:LE

[Aka97] M. Akahira. Letter to the editor. Communications in Statistics: Simulation and Computation, 26(1):397, 1997. CODEN CSSCDB. ISSN 0361-0918.

Al-Kandari:1997:VSI

[AKJ97] Noriah M. Al-Kandari and Ian T. Jolliffe. Variable selection and interpretation in canonical correlation analysis. Communications in Statistics: Simulation and Computation, 26(3):873–900, 1997. CODEN CSSCDB. ISSN 0361-0918.

Albert:1996:MAF

[Alb96] Jim Albert. A MCMC algorithm to fit a general exchangeable model. Communications in Statistics: Simulation and Computation, 25(3):573–592, 1996. CODEN CSSCDB. ISSN 0361-0918.

Alessandro:1999:OSM

[Ale99] Fassó Alessandro. One-sided Mewama control charts. Communications in Statistics: Simulation and Computation, 28(2):381– 401, 1999. CODEN CSSCDB. ISSN 0361-0918.

Allan:1999:SSR

[All99] Donner Allan. Sample size requirements for interval estimation of the intraclass kappa statistic. Communications in Statistics: Simulation and Computation, 28(2):415–429, 1999. CO-DEN CSSCDB. ISSN 0361-0918.

Agresti:1993:SEC

[ALM93] Alan Agresti, Joseph B. Lang, and Cyrus Mehta. Some empirical comparisons of exact, modified exact, and higher-order asymptotic tests of independence for ordered categorical variables. Communications in Statistics: Simulation and Computation, 22(1):1–18, 1993. CODEN CSSCDB. ISSN 0361-0918.

Arnold:1995:SSR

[ALN95] Jesse C. Arnold, G. H. Lemon, and J. M. Norton. Sample size requirements for rare or abundant attribute sampling. Communications in Statistics: Simulation and Computation, 24(1):131– 148, 1995. CODEN CSSCDB. ISSN 0361-0918.

Altekar:1995:SDT

[ALS95] Maneesha Altekar, James M. Lucas, and John H. Schuenemeyer. Shift detection in two-way arrays. Communications in Statistics: Simulation and Computation, 24(3):711–731, 1995. CO-DEN CSSCDB. ISSN 0361-0918.

Amey:1990:RMC

[Ame90] Alphonse K. A. Amey. Robustness of the multiple correlation coefficient when sampling from a mixture of two multivariate normal populations. Communications in Statistics: Simulation and Computation, 19(4):1443–1457, 1990. CODEN CSSCDB. ISSN 0361-0918.

Andres:1992:CRA

[AMH92] A. Martín Andrés, A. Silva Mato, and Tejedor I. Herranz. A critical review of asymptotic methods for comparing two proportions by means of independent samples. Communications in Statistics: Simulation and Computation, 21(2):551–586, 1992. CO-DEN CSSCDB. ISSN 0361-0918.

Anderson:1996:SIP

[AMJM96] T. W. Anderson, R. P. Mentz, N. M. Jarma, and C. I. Martinez. Simulations of iterative procedures for maximum likelihood estimation in MA(1) models. Communications in Statistics: Simulation and Computation, 25(4):851–865, 1996. CODEN CSSCDB. ISSN 0361-0918.

Andres:1991:RCN

[And91] A. Martin Andrés. A review of classic non-asymptotic methods for comparing two proportions by means of independent samples. Communications in Statistics: Simulation and Computation, 20 (2–3):551–583, 1991. CODEN CSSCDB. ISSN 0361-0918.

Andel:1994:MVS

[And94] Jiří Andēl. A model for variables with suddenly changing parameters. Communications in Statistics: Simulation and Computation, 23(1):193–206, 1994. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1990:SCR

[Ano90] Anonymous. Serial correlation or random subject effects. Communications in Statistics: Simulation and Computation, 19(3): 1105–1123, 1990. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1991:AIV

[Ano91a] Anonymous. Author index to volume 20. Communications in Statistics: Simulation and Computation, 20(4):1191–1193, 1991. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1991:EBa

[Ano91b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 20(2–3):ebi–ebii, 1991. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1991:EBb

[Ano91c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 20(4):ebi–ebii, 1991. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1992:EBa

[Ano92a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 21(1):ebi–ebii, 1992. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1992:EBb

[Ano92b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 21(2):ebi–ebii, 1992. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1992:EBc

[Ano92c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 21(3):ebi–ebii, 1992. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1992:EBd

[Ano92d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 21(4):ebi–ebii, 1992. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 22(1):ebi–ebii, 1993. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 22(2):ebi–ebii, 1993. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1993:EBc

[Ano93c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 22(3):ebi–ebii, 1993. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1993:EBd

[Ano93d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 22(4):ebi–ebii, 1993. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1994:AMU

[Ano94a] Anonymous. Analysis of means using ranks for the randomized complete block design. Communications in Statistics: Simulation and Computation, 23(2):547–568, 1994. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1994:EBa

[Ano94b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 23(1):ebi–ebii, 1994. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1994:EBb

[Ano94c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 23(2):ebi–ebii, 1994. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1994:EBc

[Ano94d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 23(3):ebi–ebii, 1994. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1994:EBd

[Ano94e] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 23(4):ebi–ebii, 1994. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1994:E

[Ano94f] Anonymous. Errata. Communications in Statistics: Simulation and Computation, 23(1):285–286, 1994. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1995:E

[Ano95a] Anonymous. Editorial. Communications in Statistics: Simulation and Computation, 24(1):v, 1995. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1995:EBa

[Ano95b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 24(1):ebi–ebii, 1995. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1995:EBb

[Ano95c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 24(2):ebi–ebii, 1995. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1995:EBc

[Ano95d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 24(3):ebi–ebii, 1995. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1996:E

[Ano96a] Anonymous. Editorial. Communications in Statistics: Simulation and Computation, 25(1):v–vi, 1996. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1996:EBa

[Ano96b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 25(1):ebi–ebii, 1996. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1996:EBb

[Ano96c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 25(2):ebi–ebii, 1996. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1996:EBc

[Ano96d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 25(3):ebi–ebii, 1996. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1996:EBd

[Ano96e] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 25(4):ebi–ebii, 1996. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 26(1):ebi–ebii, 1997. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 26(2):ebi–ebii, 1997. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1997:EBc

[Ano97c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 26(3):ebi–ebii, 1997. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1997:EBd

[Ano97d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 26(4):ebi–ebii, 1997. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1998:EBa

[Ano98a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 27(1):ebi–ebii, 1998. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1998:EBb

[Ano98b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 27(2):ebi–ebii, 1998. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1998:EBc

[Ano98c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 27(3):ebi–ebii, 1998. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1998:EBd

[Ano98d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 27(4):ebi–ebii, 1998. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1999:EBa

[Ano99a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 28(1):ebi–ebii, 1999. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1999:EBb

[Ano99b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 28(2):ebi–ebii, 1999. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1999:EBc

[Ano99c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 28(3):ebi–ebii, 1999. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1999:EBd

[Ano99d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 28(4):ebi–ebii, 1999. CODEN CSSCDB. ISSN 0361-0918.

Andre:1991:IPE

[APN91] Carmen D. S. André, Clóvis A. Peres, and Subhash C. Narula. An iterative procedure for the estimation of parameters in a doseresponse model. Communications in Statistics: Simulation and Computation, 20(2–3):763–775, 1991. CODEN CSSCDB. ISSN 0361-0918.

Appleton:1995:ULL

[App95] David R. Appleton. The uniform, logistic, and lambda distributions. Communications in Statistics: Simulation and Computation, 24(4):947–951, 1995. CODEN CSSCDB. ISSN 0361-0918.

Antille:1990:RCO

[AR90] Gérard Antille and Gilbert Ritschard. Robust and classical outlyingness indicators a simulation study. Communications in Statistics: Simulation and Computation, 19(2):505–512, 1990. CO-DEN CSSCDB. ISSN 0361-0918.

Amin:1991:NEW

[AS91] Raid W. Amin and Alice J. Searcy. A nonparametric exponentially weighted moving average control scheme. Communications in Statistics: Simulation and Computation, 20(4):1049– 1072, 1991. CODEN CSSCDB. ISSN 0361-0918.

Arumainayagam:1995:QSD

[AS95] S. Devaraj Arumainayagam and V. Soundararajan. Quick switching double sampling system indexed by the crossover point. Communications in Statistics: Simulation and Computation, 24 (3):765–773, 1995. CODEN CSSCDB. ISSN 0361-0918.

Awad:1990:SDK

[ASA90] Adnan M. Awad, Talib H. Sarie, and Mufid M. Azzam. Simulated distribution of the Kullback–Leibler information measure. Communications in Statistics: Simulation and Computation, 19 (4):1319–1338, 1990. CODEN CSSCDB. ISSN 0361-0918.

Atanu:1999:DRR

[Ata99] Biswas Atanu. Delayed response in randomized play-the-winner rule revisited. Communications in Statistics: Simulation and Computation, 28(3):715–731, 1999. CODEN CSSCDB. ISSN 0361-0918.

Andres:1992:OCC

[ATL92] A. Martín Andrés, I. Herranz Tejedor, and J. D. Luna del Castillo. Optimal correction for continuity in the chi-squared test in 2 × 2 tables. Communications in Statistics: Simulation and Computation, 21(4):1077–1101, 1992. CODEN CSSCDB. ISSN 0361-0918.

Akahira:1997:RCI

[ATT97] Masafumi Akahira, Kunihiko Takahashi, and Kei Takeuchi. Randomized confidence intervals of a parameter for a family of discrete exponential type distributions. Communications in Statistics: Simulation and Computation, 26(3):1103–1128, 1997. CO-DEN CSSCDB. ISSN 0361-0918.

Abraham:1992:TSA

[AV92] Bovas Abraham and K. Vijayan. Time series analysis for repeated surveys. Communications in Statistics: Simulation and Computation, 21(3):893–908, 1992. CODEN CSSCDB. ISSN 0361-0918.

Ananda:1996:TDT

[AW96] Malwane M. A. Ananda and Samaradasa Weerahandi. Testing the difference of two exponential means using generalized p-values. Communications in Statistics: Simulation and Computation, 25(2):521–532, 1996. CODEN CSSCDB. ISSN 0361-0918.

Atwood:1996:DLR

[AWBW^{+}96] Ld Atwood, Af Wilson, Je Bailey-Wilson, Jn Carruth, and Rc Elston. On the distribution of the likelihood ratio test statistic for a mixture of two normal distributions. Communications in Statistics: Simulation and Computation, 25(3):733–740, 1996. CO-DEN CSSCDB. ISSN 0361-0918.

Atwood:1992:CAF

[AWEBW92] Ld Atwood, Af Wilson, Rc Elston, and Je Bailey-Wilson. Computational aspects of fitting a mixture of two normal distributions using maximum likelihood. Communications in Statistics: Simulation and Computation, 21(3):769–781, 1992. CODEN CSS-CDB. ISSN 0361-0918.

Aras:1993:SNE

[AWhW93] Girish Aras, Lyn R. Whitaker, and Yang huang Wu. Sequential nonparametric estimation of an optimal age replacement policy: a simulation study. Communications in Statistics: Simulation and Computation, 22(4):1115–1134, 1993. CODEN CSSCDB. ISSN 0361-0918.

Anraku:1990:ENB

[AY90] K. Anraku and T. Yanagimoto. Estimation for the negative binomial distribution based on the conditional likelihood. Communications in Statistics: Simulation and Computation, 19(3): 771–786, 1990. CODEN CSSCDB. ISSN 0361-0918.

Birch:1993:RIR

[BA93] Jeffrey B. Birch and David B. Agard. Robust inference in regression: a comparative study. Communications in Statistics: Simulation and Computation, 22(1):217–244, 1993. CODEN CSS-CDB. ISSN 0361-0918.

Baringhaus:1994:MHB

[Bar94] L. Baringhaus. On a modification of the Hoeffding–Blum–Kiefer– Rosenblatt independence criterion. Communications in Statistics: Simulation and Computation, 23(3):683–689, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Barry:1996:CBN

[Bar96] Ronald Paul Barry. A centroid-based nonparametric regression estimator. Communications in Statistics: Simulation and Computation, 25(1):81–97, 1996. CODEN CSSCDB. ISSN 0361-0918.

Barabesi:1998:CDS

[Bar98] Lucio Barabesi. The computation of the distribution of the sign test statistic for ranked-set sampling. Communications in Statistics: Simulation and Computation, 27(3):833–842, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Balakrishnan:1992:OSEa

[BC92a] N. Balakrishnan and P. S. Chan. Order statistics from extreme value distribution, i: tables of means, variances and covariances.

Communications in Statistics: Simulation and Computation, 21 (4):1199–1217, 1992. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:1992:OSEb

[BC92b] N. Balakrishnan and P. S. Chan. Order statistics from extreme value distribution, ii: best linear unbiased estimates and some other uses. Communications in Statistics: Simulation and Computation, 21(4):1219–1246, 1992. CODEN CSSCDB. ISSN 0361-0918.

Bau:1997:LPP

[BCC97] J. J. Bau, Hubert J. Chen, and Shun-Yi Chen. Lower percentage points of Hartley’s extremal quotient statistic and their applications. Communications in Statistics: Simulation and Computation, 26(2):443–465, 1997. CODEN CSSCDB. ISSN 0361-0918.

Banerjia:1993:GRP

[BD93] Sanjeev Banerjia and Rex A. Dwyer. Generating random points in a ball. Communications in Statistics: Simulation and Computation, 22(4):1205–1209, 1993. CODEN CSSCDB. ISSN 0361-0918.

Baibing:1997:SAR

[BD97] Li Baibing and Bart De Moor. A simple approximation relation between the symmetric generalized logistic and the Student’s t distributions. Communications in Statistics: Simulation and Computation, 26(3):1029–1039, 1997. CODEN CSSCDB. ISSN 0361-0918.

Black:1990:PEN

[BDP90] C. M. Black, S. D. Durham, and V. J. Padgett. Parameter estimation for a new distribution for the strength of brittle fibers: a simulation study. Communications in Statistics: Simulation and Computation, 19(3):809–825, 1990. CODEN CSSCDB. ISSN 0361-0918.

Bedi:1995:AEM

[Bed95] Pradeep Kumar Bedi. An alternative estimator in Midzuno scheme for multiple characteristics. Communications in Statistics: Simulation and Computation, 24(1):17–30, 1995. CODEN CSSCDB. ISSN 0361-0918.

Berod:1992:SDR

[Bér92] Annick Clerc Bérod. Simulated deficiencies of robust estimators of a proportionality constant. Communications in Statistics: Simulation and Computation, 21(3):753–767, 1992. CO-DEN CSSCDB. ISSN 0361-0918.

Betro:1993:DGR

[Bet93] Bruno Betrò. On the distribution of Gini’s rank association coefficient. Communications in Statistics: Simulation and Computation, 22(2):497–505, 1993. CODEN CSSCDB. ISSN 0361-0918.

Bettonvil:1995:FSS

[Bet95] Bert Bettonvil. Factor screening by sequential bifurcation. Communications in Statistics: Simulation and Computation, 24(1): 165–185, 1995. CODEN CSSCDB. ISSN 0361-0918.

Babu:1992:AMC

[BF92] Gutti Jogesh Babu and Eric D. Feigelson. Analytical and Monte Carlo comparisons of six different linear least squares fits. Communications in Statistics: Simulation and Computation, 21(2): 533–549, 1992. CODEN CSSCDB. ISSN 0361-0918.

Bender:1993:CNR

[BG93] Ralf Bender and Ulrich Grouven. On the choice of the number of residual autocovariances for the portmanteau test of multivariate autoregressive models. Communications in Statistics: Simulation and Computation, 22(1):19–32, 1993. CODEN CSSCDB. ISSN 0361-0918.

Bain:1995:EMS

[BG95] Lee J. Bain and Gaoxiong Gan. Expectations of the maximum of sums and the sum of maxima for correlated normal variables. Communications in Statistics: Simulation and Computation, 24 (3):775–780, 1995. CODEN CSSCDB. ISSN 0361-0918.

Bechhofer:1990:CPP

[BGDH90] Robert E. Bechhofer, David M. Goldsman, Charles W. Dunnett, and Mark Hartmann. A comparison of the performances of procedures for selecting the normal population having the largest mean when the populations have a common unknown variance. Communications in Statistics: Simulation and Computation, 19 (3):971–1006, 1990. CODEN CSSCDB. ISSN 0361-0918.

Bull:1994:TSJ

[BHG94] S. B. Bull, W. W. Hauck, and C. M. T. Greenwood. Two-step jackknife bias reduction for logistic regression MLES. Communications in Statistics: Simulation and Computation, 23(1):59–88, 1994. CODEN CSSCDB. ISSN 0361-0918.

Boisvert:1991:GAM

[BHK91] Ronald F. Boisvert, Sally E. Howe, and David K. Kahaner. The guide to available mathematical software problem classification system^{1}. Communications in Statistics: Simulation and Computation, 20(4):811–842, 1991. CODEN CSSCDB. ISSN 0361-0918.

Bhimasankaram:1994:USG

[BJ94] P. Bhimasankaram and S. Rao Jammalamadaka. Updates of statistics in a general linear model: a statistical interpretation and applications. Communications in Statistics: Simulation and Computation, 23(3):789–801, 1994. CODEN CSSCDB. ISSN 0361-0918.

Bellhouse:1991:CGS

[BK91] D. R. Bellhouse and R. J. Kulperger. Computer generated simple random samples. Communications in Statistics: Simulation and Computation, 20(2–3):539–550, 1991. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1992:NHD

[BKS92] K. O. Bowman, Marvin A. Kastenbaum, and L. R. Shenton. The negative hypergeometric distribution and estimation by moments. Communications in Statistics: Simulation and Computation, 21(2):301–332, 1992. CODEN CSSCDB. ISSN 0361-0918.

Blair:1991:NCV

[Bla91] R. Clifford Blair. New critical values for the generalized t and generalized rank-sum procedures. Communications in Statistics: Simulation and Computation, 20(4):981–994, 1991. CO-DEN CSSCDB. ISSN 0361-0918.

Barbosa:1994:AAL

[BLN94] Emanuel Pimentel Barbosa and Francisco Louzada-Neto. Analysis of accelerated life tests with Weibull failure distribution via generalized linear models. Communications in Statistics: Simulation and Computation, 23(2):455–465, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Blodgett:1998:LMP

[Blo98] Robert J. Blodgett. The limit of minimizing points as a median estimator. Communications in Statistics: Simulation and Computation, 27(3):857–869, 1998. CODEN CSSCDB. ISSN 0361-0918.

Barbosa:1996:LDA

[BM96a] Emanuel P. Barbosa and Helio S. Migon. Longitudinal data analysis of animal growth via multivariate dynamic models. Communications in Statistics: Simulation and Computation, 25(2): 369–380, 1996. CODEN CSSCDB. ISSN 0361-0918.

Bekele:1996:ETA

[BM96b] Isaac Bekele and Charles E. McCulloch. Estimation of the time of attainment of a threshold with control of the probability of overestimation. Communications in Statistics: Simulation and Computation, 25(1):157–168, 1996. CODEN CSSCDB. ISSN 0361-0918.

Bartkiewicz:1998:CSM

[BN98] Jaroslaw Bartkiewicz and Wojciech Niemiro. A comparison of some MANOVA-type tests based on least distances. Communications in Statistics: Simulation and Computation, 27(2):403–430, 1998. CODEN CSSCDB. ISSN 0361-0918.

Bogen:1993:IPA

[Bog93] Kenneth T. Bogen. An intermediate-precision approximation of the inverse cumulative normal distribution. Communications in Statistics: Simulation and Computation, 22(3):797–801, 1993. CODEN CSSCDB. ISSN 0361-0918.

Boik:1996:TLF

[Boi96] Robert J. Boik. Transforming linear functions to normality: optimal component powers. Communications in Statistics: Simulation and Computation, 25(2):351–367, 1996. CODEN CSSCDB. ISSN 0361-0918.

Boyles:1994:BCA

[Boy94] Russell A. Boyles. Brocess capability with asymmetric tolerances. Communications in Statistics: Simulation and Computation, 23 (3):615–635, 1994. CODEN CSSCDB. ISSN 0361-0918.

Barry:1997:KLD

[BP97] Ronald P. Barry and R. Kelley Pace. Kriging with large data sets using sparse matrix techniques. Communications in Statistics: Simulation and Computation, 26(2):619–629, 1997. CO-DEN CSSCDB. ISSN 0361-0918.

Basu:1995:IPT

[BRB95] Srabashi Basu, Aparna Raychaudhuri, and Ayanendranath Basu. Improving precision through modelling: an illustration with hierarchical kappa. Communications in Statistics: Simulation and Computation, 24(2):399–408, 1995. CODEN CSSCDB. ISSN 0361-0918.

Bristol:1990:CTD

[Bri90] David R. Bristol. Comparison of two distribution-free procedures for multiple comparisons with a control. Communications in Statistics: Simulation and Computation, 19(4):1403–1413, 1990. CODEN CSSCDB. ISSN 0361-0918.

Bristol:1993:VTS

[Bri93] David Bristol. P -values for two-sided comparisons with a control. Communications in Statistics: Simulation and Computation, 22 (2):421–435, 1993. CODEN CSSCDB. ISSN 0361-0918.

Brooks:1997:USP

[Bro97] Robert D. Brooks. Using a sequence of point optimal tests to select a varying coefficient model. Communications in Statistics: Simulation and Computation, 26(2):671–685, 1997. CO-DEN CSSCDB. ISSN 0361-0918.

Bolfarine:1993:EBP

[BS93] Heleno Bolfarine and Monica C. Sandoval. Empirical Bayesian prediction in the location error in variables superpopulation model. Communications in Statistics: Simulation and Computation, 22(2):345–370, 1993. CODEN CSSCDB. ISSN 0361-0918.

Barton:1994:GMC

[BS94a] Russell R. Barton and Lee W. Schruben. Graphical methods for comparing confounding in two or more designs. Communications in Statistics: Simulation and Computation, 23(4):953–971, 1994. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1994:ABS

[BS94b] K. O. Bowman and L. R. Shenton. Additional Bowman–Shenton approximate percentage points for Pearson distributions based on Pearson Type VI. Communications in Statistics: Simulation and Computation, 23(3):583–590, 1994. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1997:CFE

[BS97] K. O. Bowman and L. R. Shenton. The Cornish–Fisher expansion for estimating percentage points: a numerical perspective. Communications in Statistics: Simulation and Computation, 26 (1):355–373, 1997. CODEN CSSCDB. ISSN 0361-0918.

Barbosa:1999:MSG

[BS99] Emanuel Pimentel Barbosa and Regina Sadownik. A multiplicative seasonal growth model for multivariate time series analysis and forecasting. Communications in Statistics: Simulation and Computation, 28(2):291–308, 1999. CODEN CSSCDB. ISSN 0361-0918.

Basu:1997:RTE

[BSB97] Srabashi Basu, Sahadeb Sarkar, and Ayanendranath Basu. Robust tests for equality of two population means under the normal model. Communications in Statistics: Simulation and Computation, 26(1):333–353, 1997. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1998:DRG

[BSG98] K. O. Bowman, L. R. Shenton, and Paul C. Gailey. Distribution of the ratio of gamma variates. Communications in Statistics: Simulation and Computation, 27(1):1–19, 1998. CODEN CSS-CDB. ISSN 0361-0918.

Blair:1992:DFR

[BT92] R. Clifford Blair and G. L. Thompson. A distribution-free ranklike test for scale with unequal population locations. Communications in Statistics: Simulation and Computation, 21(2):353–371, 1992. CODEN CSSCDB. ISSN 0361-0918.

BuHamra:1997:TCR

[BuH97] Sana S. BuHamra. Testing for a change in repeated measures data. Communications in Statistics: Simulation and Computation, 26(3):841–872, 1997. CODEN CSSCDB. ISSN 0361-0918.

Burford:1990:RNG

[Bur90] Roger L. Burford. Random number generators for microcomputers. Communications in Statistics: Simulation and Computation, 19(2):649–662, 1990. CODEN CSSCDB. ISSN 0361-0918.

Bilker:1997:BLT

[BW97] Warren B. Bilker and Mei-Cheng Wang. Bootstrapping left truncated and right censored data. Communications in Statistics: Simulation and Computation, 26(1):141–171, 1997. CODEN CSSCDB. ISSN 0361-0918.

Brorsen:1990:MLE

[BY90] B. Wade Brorsen and Seung Ryong Yang. Maximum likelihood estimates of symmetric stable distribution parameters. Communications in Statistics: Simulation and Computation, 19(4): 1459–1464, 1990. CODEN CSSCDB. ISSN 0361-0918.

Caudill:1998:MLE

[CA98] Steven B. Caudill and Ram N. Acharya. Maximum likelihood estimation of a mixture of normal regressions: starting values and singularities. Communications in Statistics: Simulation and Computation, 27(3):667–674, 1998. CODEN CSSCDB. ISSN 0361-0918.

Cuadras:1996:SCA

[CAF96] C. M. Cuadras, C. Areans, and J. Fortiana. Some computational aspects of a distance-based model for prediction. Communications in Statistics: Simulation and Computation, 25(3):593–609, 1996. CODEN CSSCDB. ISSN 0361-0918.

Caroni:1998:WOT

[Car98] C. Caroni. Wilks’ outlier test in more than one multivariate sample. Communications in Statistics: Simulation and Computation, 27(1):79–94, 1998. CODEN CSSCDB. ISSN 0361-0918.

Castagliola:1998:ANS

[Cas98] Philippe Castagliola. Approximation of the normal sample median distribution using symmetrical Johnson S_{U} distributions: application to quality control. Communications in Statistics: Simulation and Computation, 27(2):289–301, 1998. CODEN CSSCDB. ISSN 0361-0918.

Chandra:1990:TWS

[CC90] Nimai Kumar Chandra and Arijit Chaudhuri. On testimating the Weibull shape parameter. Communications in Statistics: Simulation and Computation, 19(2):637–648, 1990. CODEN CSSCDB. ISSN 0361-0918.

Chen:1998:SSA

[CC98a] Shun-Yi. Chen and Hubert J. Chen. Single-stage analysis of variance under heteroscedasticity. Communications in Statistics: Simulation and Computation, 27(3):641–666, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Colosimo:1998:ISI

[CC98b] Enrico A. Colosimo and Gauss M. Cordeiro. Improved statistical inference for exponential reliability data under Type II censoring. Communications in Statistics: Simulation and Computation, 27 (1):127–136, 1998. CODEN CSSCDB. ISSN 0361-0918.

Chen:1999:NOC

[CC99] Hubert J. Chen and Shun-Yi. Chen. A nearly optimal confidence interval for the largest normal mean. Communications in Statistics: Simulation and Computation, 28(1):131–146, 1999. CODEN CSSCDB. ISSN 0361-0918.

Cordeiro:1998:BRE

[CCN98] Gauss M. Cordeiro and Francisco Cribari-Neto. On bias reduction in exponential and non-exponential family regression models. Communications in Statistics: Simulation and Computation, 27(2):485–500, 1998. CODEN CSSCDB. ISSN 0361-0918.

Clapp:1993:EBE

[CD93] John M. Clapp and Dipak K. Dey. Estimating bivariate errors-in-variables models with instrumental variables. Communications in Statistics: Simulation and Computation, 22(3):863–876, 1993. CODEN CSSCDB. ISSN 0361-0918.

Cerrito:1992:USE

[Cer92] Patricia B. Cerrito. Using stratification to estimate multimodal density functions with applications to regression. Communications in Statistics: Simulation and Computation, 21(4):1149– 1164, 1992. CODEN CSSCDB. ISSN 0361-0918.

Cao:1997:SCT

[CFBGM^{+}97] R. Cao, M. Febrero-Bande, W. González-Manteiga, J. M. Prada-Sánchez, and I. García-Jurado. Saving computer time in constructing consistent bootstrap prediction intervals for autoregressive processes. Communications in Statistics: Simulation and Computation, 26(3):961–978, 1997. CODEN CSSCDB. ISSN 0361-0918.

Crivelli:1995:CIR

[CFMM95] Ana Crivelli, Luis Firinguetti, Rosa Montaño, and Margarita Muñóz. Confidence intervals in ridge regression by bootstrapping the dependent variable: a simulation study. Communications in Statistics: Simulation and Computation, 24(3):631–652, 1995. CODEN CSSCDB. ISSN 0361-0918.

Cordy:1993:ELS

[CG93] Clifford B. Cordy and Daniel A. Griffith. Efficiency of least squares estimators in the presence of spatial autocorrelation. Communications in Statistics: Simulation and Computation, 22 (4):1161–1179, 1993. CODEN CSSCDB. ISSN 0361-0918.

Ciarlini:1999:CAQ

[CGR99] P. Ciarlini, A. Gigli, and G. Regoliosi. The computation of accuracy of quality parameters by means of a Monte Carlo simulation. Communications in Statistics: Simulation and Computation, 28 (3):821–848, 1999. CODEN CSSCDB. ISSN 0361-0918.

Campbell:1994:TPN

[CH94] L. L. Campbell and J. Huang. Tail probability of a noncentral indefinite Gaussian quadratic form. Communications in Statistics: Simulation and Computation, 23(2):431–439, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Chiou:1995:IEE

[CH95] Paul Chiou and Chien-Pai Han. Interval estimation of error variance following a preliminary test in one-way random model. Communications in Statistics: Simulation and Computation, 24 (4):817–824, 1995. CODEN CSSCDB. ISSN 0361-0918.

Chakraborti:1990:OST

[Cha90] S. Chakraborti. A one-sided test of homogeneity against simple tree alternative for right-censored data. Communications in Statistics: Simulation and Computation, 19(3):879–889, 1990. CODEN CSSCDB. ISSN 0361-0918.

Chen:1993:NCU

[Che93] Yuh-Ing Chen. Nonparametric comparisons of umbrella pattern treatment effects with a control in a one-way layout. Communications in Statistics: Simulation and Computation, 22(3):749–764, 1993. CODEN CSSCDB. ISSN 0361-0918.

Chen:1994:GSP

[Che94] Yuh-Ing Chen. A generalized steel procedure for comparing several treatments with a control under random right-censorship. Communications in Statistics: Simulation and Computation, 23 (1):1–16, 1994. CODEN CSSCDB. ISSN 0361-0918.

Christoffersson:1999:RNR

[Chr99] Jan Christoffersson. Resampling a nonlinear regression model in the frequency domain. Communications in Statistics: Simulation and Computation, 28(2):329–348, 1999. CODEN CSSCDB. ISSN 0361-0918.

Chu:1995:ESR

[Chu95] C. K. Chu. The effect of the smoothness of the regression function on a bandwidth selector in nonparametric regression. Communications in Statistics: Simulation and Computation, 24(1):61–77, 1995. CODEN CSSCDB. ISSN 0361-0918.

Chakak:1995:SCM

[CK95a] Abderrahmane Chakak and Kenneth J. Koehler. A strategy for constructing multivariate distributions. Communications in Statistics: Simulation and Computation, 24(3):537–550, 1995. CODEN CSSCDB. ISSN 0361-0918.

Chien:1995:EKM

[CK95b] Wei-Ting Kary Chien and Way Kuo. Extensions of the Kaplan– Meier estimator. Communications in Statistics: Simulation and Computation, 24(4):953–964, 1995. CODEN CSSCDB. ISSN 0361-0918.

Chen:1991:PPS

[CL91] Hubert J. Chen and Kin Lam. Percentage points of a Studentized range statistic arising from non-identical normal random variables. Communications in Statistics: Simulation and Computation, 20(4):995–1047, 1991. CODEN CSSCDB. ISSN 0361-

Chan:1994:MCC

[CL94] Lai K. Chan and Guo-Ying Li. A multivariate control chart for detecting linear trends. Communications in Statistics: Simulation and Computation, 23(4):997–1012, 1994. CODEN CSSCDB. ISSN 0361-0918.

Lo:1997:RLR

[cLH97] Shu chuan Lo and Chien-Pai Han. Robust linear regression using smooth adaptive estimators. Communications in Statistics: Simulation and Computation, 26(1):1–19, 1997. CODEN CSSCDB. ISSN 0361-0918.

Chou:1999:PCH

[CMY99] Youn-Min Chou, Robert L. Mason, and John C. Young. Power comparisons for a Hotelling’s T^{2} statistic. Communications in Statistics: Simulation and Computation, 28(4):1031–1050, 1999. CODEN CSSCDB. ISSN 0361-0918.

Cribari-Neto:1998:BRO

[CNBCF98] Francisco Cribari-Neto, Denise A. Botter, Gauss M. Cordeiro, and Silvia L. P. Ferrari. Bias reduction in one-parameter exponential family models. Communications in Statistics: Simulation and Computation, 27(3):761–782, 1998. CODEN CSSCDB. ISSN 0361-0918.

Cribari-Neto:1995:ILM

[CNF95] Francisco Cribari-Neto and Silvia L. P. Ferrari. An improved Lagrange multiplier test for heteroskedasticity. Communications in Statistics: Simulation and Computation, 24(1):31–44, 1995. CODEN CSSCDB. ISSN 0361-0918.

Conlon:1992:CRS

[Con92] Michael Conlon. The controlled random search procedure for function optimization. Communications in Statistics: Simulation and Computation, 21(3):919–923, 1992. CODEN CSSCDB. ISSN 0361-0918.

Cordy:1997:DBC

[Cor97] Clifford B. Cordy. Design-based confidence limits for a distribution function. Communications in Statistics: Simulation and Computation, 26(4):1447–1464, 1997. CODEN CSSCDB. ISSN 0361-0918.

Cordeiro:1992:ELS

[CP92] Gauss M. Cordeiro and Gilberto A. Paula. Estimation, largesample parametric tests and diagnostics for non-exponential family nonlinear models. Communications in Statistics: Simulation and Computation, 21(1):149–172, 1992. CODEN CSSCDB. ISSN 0361-0918.

Chou:1993:PTS

[CP93] Youn-Min Chou and Alan Polansky. Power of tests for some process capability indices. Communications in Statistics: Simulation and Computation, 22(2):523–544, 1993. CODEN CSSCDB. ISSN 0361-0918.

Chen:1997:TLC

[CPC97] Keh-Wei Chen, Alex S. Papadopoulos, and Shein Chung Chow. On the test of Liu and Chow’s procedure for assessing equivalence in variability of bioavailability. Communications in Statistics: Simulation and Computation, 26(3):1129–1138, 1997. CODEN CSSCDB. ISSN 0361-0918.

Choudhury:1997:LER

[CPS97] Askar H. Choudhury, Simon Power, and Robert D. St. Louis. Linear estimation of the regression model with ARMA disturbances: a simulation study. Communications in Statistics: Simulation and Computation, 26(1):315–332, 1997. CODEN CSS-CDB. ISSN 0361-0918.

Christensen:1997:CTE

[CR97] William F. Christensen and Alvin C. Rencher. A comparison of Type I error rates and power levels for seven solutions to the multivariate Behrens–Fisher problem. Communications in Statistics: Simulation and Computation, 26(4):1251–1273, 1997. CODEN CSSCDB. ISSN 0361-0918.

Crow:1990:RPC

[Cro90] Edwin L. Crow. Ranking paired contestants. Communications in Statistics: Simulation and Computation, 19(3):749–769, 1990. CODEN CSSCDB. ISSN 0361-0918.

Chattamvelli:1995:ECN

[CS95a] R. Chattamvelli and R. Shanmugam. Efficient computation of the noncentral x^{2} distribution. Communications in Statistics: Simulation and Computation, 24(3):675–689, 1995. CODEN CSSCDB. ISSN 0361-0918.

Chu:1995:CHT

[CS95b] David P. T. Chu and Brajendra C. Sutradhar. On Cochran’s and Hartley’s tests for homogeneity of variances when observations are autocorrelated. Communications in Statistics: Simulation and Computation, 24(2):327–347, 1995. CODEN CSSCDB. ISSN 0361-0918.

Chun-tu:1990:EDO

[Ct90] Lin Chun-tu. Extremes of determinants and optimality of canonical variables. Communications in Statistics: Simulation and Computation, 19(4):1415–1430, 1990. CODEN CSSCDB. ISSN 0361-0918.

Chun-tu:1991:ELS

[Ct91] Lin Chun-tu. The efficiency of least squares estimators of a seemingly unrelated regression model. Communications in Statistics: Simulation and Computation, 20(4):919–925, 1991. CO-DEN CSSCDB. ISSN 0361-0918.

Chalton:1992:IOI

[CT92] Derek O. Chalton and Cas. G. Troskie. Identification of outlying and influential data with biased estimation: A simulation study. Communications in Statistics: Simulation and Computation, 21 (3):607–626, 1992. CODEN CSSCDB. ISSN 0361-0918.

Coe:1993:SSC

[CT93] Paul R. Coe and Ajit C. Tamhane. Small sample confidence intervals for the difference,ratio and odds ratio of two success probabilities. Communications in Statistics: Simulation and Computation, 22(4):925–938, 1993. CODEN CSSCDB. ISSN 0361-0918.

Chen:1999:AAE

[CT99] Chien-Chung Chen and Christopher W. Tyler. Accurate approximation to the extreme order statistics of Gaussian samples. Communications in Statistics: Simulation and Computation, 28(1): 177–188, 1999. CODEN CSSCDB. ISSN 0361-0918.

Chou:1994:UPF

[CTH^{+}94] Youn-Min Chou, S. Turner, S. Henson, D. Meyer, and K. S. Chen. On using percentiles to fit data by a Johnson distribution. Communications in Statistics: Simulation and Computation, 23 (2):341–354, 1994. CODEN CSSCDB. ISSN 0361-0918.

Cutler:1993:BBA

[Cut93] Adele Cutler. A branch and bound algorithm for constrained least squares. Communications in Statistics: Simulation and Computation, 22(2):305–321, 1993. CODEN CSSCDB. ISSN 0361-0918.

Chen:1994:MAA

[CV94] Zhenmin Chen and John W. Van Ness. Metric admissibility and agglomerative clustering. Communications in Statistics: Simulation and Computation, 23(3):833–845, 1994. CODEN CSSCDB. ISSN 0361-0918.

Chou:1996:TME

[CW96] Shihyu Chou and Min-Chiang Wang. Taguchi’s method and an empirical procedure for on-line process control. Communications in Statistics: Simulation and Computation, 25(2):301–320, 1996. CODEN CSSCDB. ISSN 0361-0918.

Czitrom:1992:NME

[Czi92] Veronica Czitrom. Note on a mixture experiment with process variables. Communications in Statistics: Simulation and Computation, 21(2):493–498, 1992. CODEN CSSCDB. ISSN 0361-0918.

Aubin:1997:BLR

[dCQAC97] Elisete da Conceição Q. Aubin and Gauss M. Cordeiro. BIAS in linear regression models with unknown covariance matrix. Communications in Statistics: Simulation and Computation, 26(3): 813–828, 1997. CODEN CSSCDB. ISSN 0361-0918.

DeLosReyes:1990:UBS

[De 90] J. P. De Los Reyes. Upper bound for singular normal probability integrals by integration over a hypersphere. Communications in Statistics: Simulation and Computation, 19(3):847–861, 1990. CODEN CSSCDB. ISSN 0361-0918.

DeGooijer:1995:CVC

[De 95] Jan G. De Gooijer. Cross-validation criteria for covariance structures. Communications in Statistics: Simulation and Computation, 24(1):1–16, 1995. CODEN CSSCDB. ISSN 0361-0918.

Dion:1999:SPB

[DE99] J. P. Dion and T. W. Epps. Stock prices as branching processes in random environments: estimation. Communications in Statistics: Simulation and Computation, 28(4):957–975, 1999. CODEN CSSCDB. ISSN 0361-0918.

Devroye:1992:BPM

[Dev92] Luc Devroye. A branching process method in Lagrange random variate generation. Communications in Statistics: Simulation and Computation, 21(1):1–14, 1992. CODEN CSSCDB. ISSN 0361-0918.

Marques:1992:ABD

[dFMK92] Eliana H. de F. Marques and Gary G. Koch. Analysis of bivariate dichotomous data from a stratified two-stage cluster sample. Communications in Statistics: Simulation and Computation, 21 (3):731–752, 1992. CODEN CSSCDB. ISSN 0361-0918.

Deng:1990:GUV

[DG90] Lih-Yuan Deng and E. Olusegun George. Generation of uniform variates from several nearly uniformly distributed variables. Communications in Statistics: Simulation and Computation, 19 (1):145–154, 1990. CODEN CSSCDB. ISSN 0361-0918.

Delong:1994:GLD

[DG94] D. M. Delong and G. H. Guirguis. Gauss–Laguerre dynamics: an efficient method of computing integrals in multiple comparisons. Communications in Statistics: Simulation and Computation, 23 (4):1047–1059, 1994. CODEN CSSCDB. ISSN 0361-0918.

Douglas:1996:BSE

[DG96] Stratford Douglas and David K. Guilkey. Bootstrap standard error estimates in a switching regression model with unknown switch point. Communications in Statistics: Simulation and Computation, 25(1):1–23, 1996. CODEN CSSCDB. ISSN 0361-0918.

Dhar:1992:CCM

[Dha92] Sunil K. Dhar. Computation of certain minimum L_{2}-distance type estimators under the linear model. Communications in Statistics: Simulation and Computation, 21(1):203–220, 1992. CODEN CSSCDB. ISSN 0361-0918.

Dias:1999:SAN

[Dia99] Ronaldo Dias. Sequential adaptive nonparametric regression via H-splines. Communications in Statistics: Simulation and Computation, 28(2):501–515, 1999. CODEN CSSCDB. ISSN 0361-0918.

Dielman:1992:SSP

[Die92] Terry E. Dielman. Small sample properties of random coefficient regression estimators: a Monte Carlo simulation. Communications in Statistics: Simulation and Computation, 21(1):103–132, 1992. CODEN CSSCDB. ISSN 0361-0918.

Ding:1994:EAC

[Din94] Cherng Ding. An efficient algorithm for computing the noncentrality parameters of chi-squared tests. Communications in Statistics: Simulation and Computation, 23(3):861–870, 1994. CODEN CSSCDB. ISSN 0361-0918.

Ding:1997:UNM

[Din97] Cherng G. Ding. On using Newton’s method for computing the noncentrality parameter of the noncentral F distribution. Communications in Statistics: Simulation and Computation, 26(1): 259–268, 1997. CODEN CSSCDB. ISSN 0361-0918.

deFMarques:1996:MDC

[dK96] Eliana H. de F. Marques and Gary G. Koch. A model for distributions and correlation structure of multivariate polytomous data from a two-stage cluster sample. Communications in Statistics: Simulation and Computation, 25(1):25–39, 1996. CODEN CSSCDB. ISSN 0361-0918.

Doray:1997:EEG

[DL97] Louis G. Doray and Andrew Luong. Efficient estimators for the good family. Communications in Statistics: Simulation and Computation, 26(3):1075–1088, 1997. CODEN CSSCDB. ISSN 0361-0918.

Dielman:1994:CQE

[DLP94] Terry Dielman, Cynthia Lowry, and Roger Pfaffenberger. A comparison of quantile estimators. Communications in Statistics: Simulation and Computation, 23(2):355–371, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Donatos:1990:SSP

[DM90] George S. Donatos and George C. Michailidis. Small sample properties of ridge estimators with normal and non-normal disturbances. Communications in Statistics: Simulation and Computation, 19(3):935–950, 1990. CODEN CSSCDB. ISSN 0361-0918.

Dozier:1993:SSP

[DM93] W. Gregory Dozier and Keith E. Muller. Small-sample power of uncorrected and Satterthwaite corrected t tests for comparing binomial proportions. Communications in Statistics: Simulation and Computation, 22(1):245–264, 1993. CODEN CSSCDB. ISSN 0361-0918.

Disantostefano:1995:CPA

[DM95] Rachael L. Disantostefano and Keith E. Muller. A comparison of power approximations for Satterthwaite’s test. Communications in Statistics: Simulation and Computation, 24(3):583–593, 1995. CODEN CSSCDB. ISSN 0361-0918.

Drezner:1999:TSM

[DMS99] Zvi Drezner, George A. Marcoulides, and Said Salhi. Tabu search model selection in multiple regression analysis. Communications in Statistics: Simulation and Computation, 28(2):349–367, 1999. CODEN CSSCDB. ISSN 0361-0918.

Dowling:1997:EPD

[DN97] Mary M. Dowling and Miguel Nakamura. Estimating parameters for discrete distributions via the empirical probability generating function. Communications in Statistics: Simulation and Computation, 26(1):301–313, 1997. CODEN CSSCDB. ISSN 0361-0918.

Donegani:1990:NPC

[Don90] M. Donegani. A nonparametric procedure for choosing a location test. Communications in Statistics: Simulation and Computation, 19(4):1393–1401, 1990. CODEN CSSCDB. ISSN 0361-0918.

Donaldson:1996:CIC

[Don96] R. W. Donaldson. Calculating inverse CV, skew and PWM functions for Pearson-3, log-normal, extreme-value and log-logistic distributions. Communications in Statistics: Simulation and Computation, 25(3):741–747, 1996. CODEN CSSCDB. ISSN 0361-0918.

PortaNova:1999:PRD

[dOP99] Acácio M. de O. Porta Nova. A pair of rounded distributions for discretized observations in discrete event simulation. Communications in Statistics: Simulation and Computation, 28(3): 733–742, 1999. CODEN CSSCDB. ISSN 0361-0918.

Dorvlo:1993:CTA

[Dor93a] Atsu S. S. Dorvlo. Comparison of two asymptotic techniques for estimating error rates in discriminant analysis. Communications in Statistics: Simulation and Computation, 22(2):461–470, 1993. CODEN CSSCDB. ISSN 0361-0918.

Dorvlo:1993:UMC

[Dor93b] Atsu S. S. Dorvlo. A unified method for computing probabilities of misclassification when the covariance matrices are known. Communications in Statistics: Simulation and Computation, 22 (3):679–687, 1993. CODEN CSSCDB. ISSN 0361-0918.

Dielman:1990:TLH

[DP90] Terry E. Dielman and Roger C. Pfaffenberger. Tests of linear hypotheses and LAV estimation: A Monte Carlo comparison. Communications in Statistics: Simulation and Computation, 19 (4):1179–1199, 1990. CODEN CSSCDB. ISSN 0361-0918.

Dragalin:1997:DAC

[Dra97] Vladimir Dragalin. The design and analysis of 2-CUSUM procedure. Communications in Statistics: Simulation and Computation, 26(1):67–81, 1997. CODEN CSSCDB. ISSN 0361-0918.

Drezner:1990:AMN

[Dre90] Zvi Drezner. Approximations to the multivariate normal integral. Communications in Statistics: Simulation and Computation, 19 (2):527–534, 1990. CODEN CSSCDB. ISSN 0361-0918.

Doganaksoy:1991:CAC

[DS91] Necip Doganaksoy and Josef Schmee. Comparisons of approximate confidence intervals for the smallest extreme value distribution simple linear regression model under time censoring. Communications in Statistics: Simulation and Computation, 20(4): 1085–1113, 1991. CODEN CSSCDB. ISSN 0361-0918.

Dunsmuir:1996:DWM

[DSW96] William T. M. Dunsmuir, David J. Scott, and Qiu Wang. The distribution of the weighted moving median of a sequence of iid observations. Communications in Statistics: Simulation and Computation, 25(4):1015–1029, 1996. CODEN CSSCDB. ISSN 0361-0918.

Dunsmuir:1992:SSE

[Dun92] William T. M. Dunsmuir. A simulation study of l_{1}: estimation of a seasonal moving average time series model. Communications in Statistics: Simulation and Computation, 21(2):519–531, 1992. CODEN CSSCDB. ISSN 0361-0918.

Dunn:1994:SPP

[Dun94] Charles L. Dunn. Simulating percentile points: stopping rules, precision, and scale. Communications in Statistics: Simulation and Computation, 23(4):1013–1026, 1994. CODEN CSSCDB. ISSN 0361-0918.

Davis:1994:SPC

[DW94] Robert E. Davis and William H. Woodall. A study of parabolic control limits for the EWMA control chart. Communications in Statistics: Simulation and Computation, 23(1):17–26, 1994. CODEN CSSCDB. ISSN 0361-0918.

Edlund:1990:RET

[Edl90] Per-Olov Edlund. Ridge estimation of transfer function weights. Communications in Statistics: Simulation and Computation, 19 (2):451–468, 1990. CODEN CSSCDB. ISSN 0361-0918.

El-Din:1997:OSD

[EDMAYS97] M. M. Mohie El-Din, M. A. W. Mahmoud, S. E. Abu-Youssef, and K. S. Sultan. Order statistics from the doubly truncated linear-exponential distribution and its characterizations. Communications in Statistics: Simulation and Computation, 26(1): 281–290, 1997. CODEN CSSCDB. ISSN 0361-0918.

El-Din:1991:MOS

[EDMY91] M. M. Mohie El-Din, M. A. W. Mahmoud, and S. E. Abo Youssef. Moments of order statistics from parabolic and skewed distributions and a characterization of Weibull distribution. Communications in Statistics: Simulation and Computation, 20(2–3): 639–645, 1991. CODEN CSSCDB. ISSN 0361-0918.

El-Helbawy:1994:ODA

[EHAA94] Abdalla T. El-Helbawy, Essam A. Ahmed, and Abdullah H. Alharbey. Optimal designs for asymmetrical factorial paired comparison experiments. Communications in Statistics: Simulation and Computation, 23(3):663–681, 1994. CODEN CSSCDB. ISSN 0361-0918.

Ernst:1993:MCS

[EK93] Michael D. Ernst and James L. Kepner. A Monte Carlo study of rank tests for repeated measures designs. Communications in Statistics: Simulation and Computation, 22(3):671–678, 1993. CODEN CSSCDB. ISSN 0361-0918.

Ekblom:1990:GTP

[Ekb90] Håkan Ekblom. Generation of test problems for L_{p} — and Huber regression. Communications in Statistics: Simulation and Computation, 19(2):481–489, 1990. CODEN CSSCDB. ISSN 0361-0918.

Elewa:1992:TEE

[ENN92] S. A. Elewa, B. N. Nagarsenker, and P. B. Nagarsenker. On a test of equality of exponential populations based on censored samples. Communications in Statistics: Simulation and Computation, 21 (3):909–917, 1992. CODEN CSSCDB. ISSN 0361-0918.

El-Taha:1993:MES

[ET93] Muhammad El-Taha. MVU estimation in a shifted gamma distribution with shape parameter a known integer. Communications in Statistics: Simulation and Computation, 22(3):831–843, 1993. CODEN CSSCDB. ISSN 0361-0918.

Falk:1999:SAG

[Fal99] Michael Falk. A simple approach to the generation of uniformly distributed random variables with prescribed correlations. Communications in Statistics: Simulation and Computation, 28(3): 785–791, 1999. CODEN CSSCDB. ISSN 0361-0918.

Famoye:1997:PEG

[Fam97] Felix Famoye. Parameter estimation for generalized negative binomial distribution. Communications in Statistics: Simulation and Computation, 26(1):269–279, 1997. CODEN CSSCDB. ISSN 0361-0918.

Fan:1994:CAB

[Fan94] Tsai-Hung Fan. Computational algorithms in Bayesian inferences for a normal mean with t prior distributions. Communications in Statistics: Simulation and Computation, 23(4):1155–1174, 1994. CODEN CSSCDB. ISSN 0361-0918.

Fan:1999:CSR

[Fan99] Shu-Kai S. Fan. Computational schema on ridge analysis. Communications in Statistics: Simulation and Computation, 28(3): 767–783, 1999. CODEN CSSCDB. ISSN 0361-0918.

Farebrother:1992:RLI

[Far92] R. W. Farebrother. Relative local influence and the condition number. Communications in Statistics: Simulation and Computation, 21(3):707–710, 1992. CODEN CSSCDB. ISSN 0361-0918.

Futschik:1998:RIT

[FH98] Andreas Futschik and Marcus Hudec. Rotation invariant tests for discrete uniformity against peak shaped alternatives. Communications in Statistics: Simulation and Computation, 27(2): 431–457, 1998. CODEN CSSCDB. ISSN 0361-0918.

Firinguetti:1991:EPL

[Fir91] Luis Firinguetti. The exact properties of the Lawless and Wang operational ridge regression estimator in a misspecified regression equation. Communications in Statistics: Simulation and Computation, 20(2–3):689–714, 1991. CODEN CSSCDB. ISSN 0361-0918.

Felix:1992:EGP

[FL92] Famoye Felix and Carl M.-S. Lee. Estimation of generalized Poisson distribution. Communications in Statistics: Simulation and Computation, 21(1):173–188, 1992. CODEN CSSCDB. ISSN 0361-0918.

Fung:1998:CVT

[FN98] Wing K. Fung and M. C. Ngai. Critical values for testing for a single outlier in a nonlinear regression model. Communications in Statistics: Simulation and Computation, 27(1):95–105, 1998. CODEN CSSCDB. ISSN 0361-0918.

Ferrer:1993:SSE

[FR93] Alberto J. Ferrer and Rafael Romero. Small samples estimation of dispersion effects from unreplicated data. Communications in Statistics: Simulation and Computation, 22(4):975–995, 1993. CODEN CSSCDB. ISSN 0361-0918.

Fellingham:1995:SPI

[FR95] G. W. Fellingham and T. E. Raghunathan. Sensitivity of point and interval estimates to distributional assumptions in longitudinal data analysis of small samples. Communications in Statistics: Simulation and Computation, 24(3):617–630, 1995. CO-DEN CSSCDB. ISSN 0361-0918.

From:1991:MSE

[Fro91] Steven G. From. Mean square error efficient estimation of an exponential mean under an exchangeable single outlier model. Communications in Statistics: Simulation and Computation, 20 (4):1073–1084, 1991. CODEN CSSCDB. ISSN 0361-0918.

Field:1993:SSI

[FT93] C. A. Field and M. A. Tingley. Small sample intervals for generalized linear regression. Communications in Statistics: Simulation and Computation, 22(3):689–707, 1993. CODEN CSSCDB. ISSN 0361-0918.

Fukushige:1994:MCS

[Fuk94] Mototsugu Fukushige. A Monte Carlo study on the interaction between model selection and testing nonnormality in autoregressive models. Communications in Statistics: Simulation and Computation, 23(4):925–937, 1994. CODEN CSSCDB. ISSN 0361-0918.

Fule:1995:ERR

[Fül95] Erika Füle. On ecological regression and ridge estimation. Communications in Statistics: Simulation and Computation, 24(2): 385–398, 1995. CODEN CSSCDB. ISSN 0361-0918.

Franklin:1992:NCN

[FW92] LeRoy A. Franklin and Gary S. Wasserman. A note on the conservative nature of the tables of lower confidence limits for c_{pk} with a suggested correction. Communications in Statistics: Simulation and Computation, 21(4):1165–1169, 1992. CODEN CSS-CDB. ISSN 0361-0918.

Frisen:1999:ELR

[FW99] M. Frisén and P. Wessman. Evaluations of likelihood ratio methods for surveillance. Communications in Statistics: Simulation and Computation, 28(3):597–622, 1999. CODEN CSSCDB. ISSN 0361-0918.

Fraser:1998:ANC

[FWW98] D. A. S. Fraser, J. Wu, and A. C. M. Wong. An approximation for the noncentral chi-squared distribution. Communications in Statistics: Simulation and Computation, 27(2):275–287, 1998. CODEN CSSCDB. ISSN 0361-0918.

Fang:1998:DRK

[FZL98] Kai-Tai Fang, Zukang Zheng, and Wenliang Lu. Discrepancy with respect to Kaplan–Meier estimator. Communications in Statistics: Simulation and Computation, 27(2):329–344, 1998. CODEN CSSCDB. ISSN 0361-0918.

Gan:1990:MPO

[Gan90] F. F. Gan. Monitoring Poisson observations using modified exponentially weighted moving average control charts. Communications in Statistics: Simulation and Computation, 19(1):103–124, 1990. CODEN CSSCDB. ISSN 0361-0918.

Gan:1994:ODC

[Gan94] F. F. Gan. An optimal design of cumulative sum control chart based on median run length. Communications in Statistics: Simulation and Computation, 23(2):485–503, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Garthwaite:1994:APD

[Gar94] Paul H. Garthwaite. Assessment of prior distributions for regression models: an experimental study. Communications in Statistics: Simulation and Computation, 23(4):871–895, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Gimotty:1990:IPC

[GB90] Phyllis A. Gimotty and Morton B. Brown. Imputation procedures for categorical data: their effects on the goodness-of-fit chi-square statistic. Communications in Statistics: Simulation and Computation, 19(2):681–703, 1990. CODEN CSSCDB. ISSN 0361-0918.

Gregurich:1997:BAE

[GB97] Mary Ann Gregurich, Ph. D. and Lyle D. Broemeling, Ph. D. A Bayesian analysis for estimating the common mean of independent normal populations using the Gibbs sampler. Communications in Statistics: Simulation and Computation, 26(1):35–51, 1997. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:1995:DRD

[GD95] Subir Ghosh and Edna Derderian. Determination of robust design against noise factors and in presence of signal factors. Communications in Statistics: Simulation and Computation, 24(2): 309–326, 1995. CODEN CSSCDB. ISSN 0361-0918.

Geary:1991:RML

[Gea91] D. N. Geary. Repeated measurements, log likelihood ratio tests and small samples. Communications in Statistics: Simulation and Computation, 20(4):887–900, 1991. CODEN CSSCDB. ISSN 0361-0918.

Govindaraju:1997:SIR

[GG97] K. Govindaraju and S. Ganesalingam. Sampling inspection for resubmitted lots. Communications in Statistics: Simulation and Computation, 26(3):1163–1176, 1997. CODEN CSSCDB. ISSN 0361-0918.

Gunsolley:1995:SSC

[GGC95] J. C. Gunsolley, C. Getchell, and V. M. Chinchilli. Small sample characteristics of generalized estimating equations. Communications in Statistics: Simulation and Computation, 24(4):869–878, 1995. CODEN CSSCDB. ISSN 0361-0918.

Gibbons:1991:CLR

[GH91] Laura E. Gibbons and David W. Hosmer, Jr. Conditional logistic regression with missing data. Communications in Statistics: Simulation and Computation, 20(1):109–120, 1991. CO-DEN CSSCDB. ISSN 0361-0918.

Gill:1990:MCS

[Gil90] P. S. Gill. A Monte Carlo simulation study of analyses of block designs under correlated errors model. Communications in Statistics: Simulation and Computation, 19(1):175–188, 1990. CODEN CSSCDB. ISSN 0361-0918.

Govindaraju:1991:DSS

[GK91] K. Govindaraju and V. Kuralmani. Determination of single sampling attributes plans for given (AQL, AOQL). Communications in Statistics: Simulation and Computation, 20(2–3):529– 538, 1991. CODEN CSSCDB. ISSN 0361-0918.

Govindaraju:1992:MTS

[GK92] K. Govindaraju and V. Kuralmani. Modified tables for the selection of QSS-1 quick switching system for a given (AQL, LQL). Communications in Statistics: Simulation and Computation, 21 (4):1103–1123, 1992. CODEN CSSCDB. ISSN 0361-0918.

Gupta:1997:ECM

[GK97a] A. K. Gupta and D. G. Kabe. Elliptically contoured model and factorization of Wilks’ A. Communications in Statistics: Simulation and Computation, 26(3):805–811, 1997. CODEN CSSCDB. ISSN 0361-0918.

Gupta:1997:SRS

[GK97b] A. K. Gupta and D. G. Kabe. Some results for a superiority problem in misspecified restricted linear models. Communications in Statistics: Simulation and Computation, 26(4):1241–1249, 1997. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:1997:MIO

[GL97] Subir Ghosh and Chinglin Lai. Measuring influence of observations in predict1on and estimation for central composite designs. Communications in Statistics: Simulation and Computation, 26 (1):233–257, 1997. CODEN CSSCDB. ISSN 0361-0918.

Govindaraju:1999:DMR

[GL99] K. Govindaraju and C. D. Lai. Design of multiple run sampling plan. Communications in Statistics: Simulation and Computation, 28(1):1–11, 1999. CODEN CSSCDB. ISSN 0361-0918.

Goddard:1994:TCU

[God94] M. J. Goddard. Tables to calculate untrimmed range and standard deviation estimates from trimmed estimates. Communications in Statistics: Simulation and Computation, 23(1):207–221, 1994. CODEN CSSCDB. ISSN 0361-0918.

Gordon:1990:MLE

[Gor90] N. H. Gordon. Maximum likelihood estimation for mixtures of two Gompertz distributions when censoring occurs. Communications in Statistics: Simulation and Computation, 19(2):733–747, 1990. CODEN CSSCDB. ISSN 0361-0918.

Gould:1997:IBS

[Gou97] A. Lawrence Gould. Issues in blinded sample size re-estimation. Communications in Statistics: Simulation and Computation, 26 (3):1229–1239, 1997. CODEN CSSCDB. ISSN 0361-0918.

Gonzalez:1999:AEA

[GPPL99] Francisco J. Gonzalez, Jaime Puig-Pey, and Alberto Luceno. Analytical expressions for the average adjustment interval and mean squared deviation for bounded adjustment schemes. Communications in Statistics: Simulation and Computation, 28(3):623–635, 1999. CODEN CSSCDB. ISSN 0361-0918.

Gehringer:1992:NPD

[GR92] Kevin R. Gehringer and Richard A. Redner. Nonparametric probability density estimation using normalized B-splines. Communications in Statistics: Simulation and Computation, 21(3): 849–878, 1992. CODEN CSSCDB. ISSN 0361-0918.

Gray:1993:AIN

[Gra93] J. Brian Gray. Approximating the internal norm influence measure in linear regression. Communications in Statistics: Simulation and Computation, 22(1):117–135, 1993. CODEN CSSCDB. ISSN 0361-0918.

Gardiner:1990:SSP

[GS90a] Joseph C. Gardiner and James Ii. Stapleton. A simulation study on the performance of bounded length confidence intervals for quantiles with censored data. Communications in Statistics: Simulation and Computation, 19(2):419–432, 1990. CODEN CSSCDB. ISSN 0361-0918.

Govindaraju:1990:SSS

[GS90b] K. Govindaraju and K. Subramani. Selection of single sampling attributes plan for given acceptable quality level and limiting quality level involving minimum risks. Communications in Statistics: Simulation and Computation, 19(4):1293–1302, 1990. CO-DEN CSSCDB. ISSN 0361-0918.

Govindaraju:1992:SDS

[GS92] K. Govindaraju and K. Subramani. Selection of double sampling attributes plan for given acceptable quality level and limiting quality level. Communications in Statistics: Simulation and Computation, 21(1):221–242, 1992. CODEN CSSCDB. ISSN 0361-0918.

Guttman:1994:ISA

[GS94] Irwin Guttman and David P. M. Scollnik. An index sampling algorithm for the Bayesian analysis of a class of model selection problems. Communications in Statistics: Simulation and Computation, 23(2):323–339, 1994. CODEN CSSCDB. ISSN 0361-0918.

Gupta:1998:ERB

[GS98] Ramesh C. Gupta and Sundarraman Subramanian. Estimation of reliability in a bivariate normal distribution with equal coefficients of variation. Communications in Statistics: Simulation and Computation, 27(3):675–698, 1998. CODEN CSSCDB. ISSN 0361-0918.

Giraudo:1999:ITS

[GS99] Maria Teresa Giraudo and Laura Sacerdote. An improved technique for the simulation of first passage times for diffusion processes. Communications in Statistics: Simulation and Computation, 28(4):1135–1163, 1999. CODEN CSSCDB. ISSN 0361-0918.

Guirguis:1990:GEC

[GT90] G. Guirguis and R. D. Tobias. Generalization and efficient computation of the Box–Meyer analysis for orthogonal designs. Communications in Statistics: Simulation and Computation, 19(2): 721–732, 1990. CODEN CSSCDB. ISSN 0361-0918.

Geweke:1999:MCM

[GT99] John Geweke and Hisashi Tanizaki. On Markov chain Monte Carlo methods for nonlinear and non-Gaussian state-space models. Communications in Statistics: Simulation and Computation, 28(4):867–894, 1999. CODEN CSSCDB. ISSN 0361-0918.

Guirguis:1990:NCN

[Gui90] Georges H. Guirguis. A note on computing the noncentrality parameter of the noncentral F -distribution. Communications in Statistics: Simulation and Computation, 19(4):1497–1511, 1990. CODEN CSSCDB. ISSN 0361-0918.

Gamage:1998:SPS

[GW98] Jinadasa Gamage and Sam Weerahandi. Size performance of some tests in one-way ANOVA. Communications in Statistics:

Simulation and Computation, 27(3):625–640, 1998. CODEN CSSCDB. ISSN 0361-0918.

Haddad:1998:SMC

[Had98] John N. Haddad. A simple method for computing the covariance matrix and its inverse of a stationary autoregressive process. Communications in Statistics: Simulation and Computation, 27 (3):617–623, 1998. CODEN CSSCDB. ISSN 0361-0918.

Haim:1996:NES

[Hai96] Shore Haim. A new estimate of skewness with mean-squared error smaller than that of the sample skewness. Communications in Statistics: Simulation and Computation, 25(2):403–414, 1996. CODEN CSSCDB. ISSN 0361-0918.

Halva:1996:EBC

[Hal96] Adel M. Halva. Estimating the Box–Cox transformation via an artificial regression model. Communications in Statistics: Simulation and Computation, 25(2):331–350, 1996. CODEN CSS-CDB. ISSN 0361-0918.

Hasegawa:1995:SSP

[Has95] Hikaru Hasegawa. On small sample properties of Zellner’s estimator for the case of two SUR equations with compound normal disturbances. Communications in Statistics: Simulation and Computation, 24(1):45–59, 1995. CODEN CSSCDB. ISSN 0361-0918.

Huda:1998:MDE

[HAS98] S. Huda and A. A. Al-Shiha. Minimax designs for estimating the slope of a third-order response surface in a hypercubic region. Communications in Statistics: Simulation and Computation, 27 (2):345–356, 1998. CODEN CSSCDB. ISSN 0361-0918.

Hawkins:1992:EAR

[Haw92] Douglas M. Hawkins. Evaluation of average run lengths of cumulative sum charts for an arbitrary data distribution. Communications in Statistics: Simulation and Computation, 21(4): 1001–1020, 1992. CODEN CSSCDB. ISSN 0361-0918.

Harris:1994:HDP

[HB94] Ian R. Harris and Ayanendranath Basu. Hellinger distance as a penalized log likelihood. Communications in Statistics: Simulation and Computation, 23(4):1097–1113, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Huang:1998:NCP

[HB98] Yeu-Shiang Huang and Vicki M. Bier. A natural conjugate prior for the non-homogeneous Poisson process with a power law intensity function. Communications in Statistics: Simulation and Computation, 27(2):525–551, 1998. CODEN CSSCDB. ISSN 0361-0918.

Hong:1997:RC

[HC97] Chong Sun Hong and Hyun Jip Choi. On l_{1} regression coefficients. Communications in Statistics: Simulation and Computation, 26(2):531–537, 1997. CODEN CSSCDB. ISSN 0361-0918.

Hill:1997:JBS

[HCNA97] R. Carter Hill, Phillip A. Cartwright, A. C. Nielsen, and Julia F. Arbaugh. Jackknifing the bootstrap: some Monte Carlo evidence. Communications in Statistics: Simulation and Computation, 26 (1):125–139, 1997. CODEN CSSCDB. ISSN 0361-0918.

Hormann:1994:TRM

[HD94] Wolfgang Hörmann and Gerhard Derflinger. The transformed rejection method for generating random variables, an alternative to the ratio of uniforms method. Communications in Statistics: Simulation and Computation, 23(3):847–860, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Hentges:1998:PDB

[HD98] Adão L. Hentges and Ian R. Dunsmore. Predictive distributions in binary models with missing data. Communications in Statistics: Simulation and Computation, 27(3):735–759, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Hamdy:1993:RTM

[HEB93] M. I. Hamdy and M. Y. El-Bassiouni. A revision of the Tukey multiple comparisons procedure to control the probability of committing at most one Type I error. Communications in Statistics: Simulation and Computation, 22(3):739–748, 1993. CO-DEN CSSCDB. ISSN 0361-0918.

Hertzberg:1997:SET

[Her97] Vicki Stover Hertzberg. Simulation evaluation of three models for correlated binary data with covariates specific to each binary observation. Communications in Statistics: Simulation and Computation, 26(1):375–396, 1997. CODEN CSSCDB. ISSN 0361-0918.

Heo:1998:PTA

[HG98] Moonseong Heo and K. Ruben Gabriel. A permutation test of association between configurations by means of the RV coefficient. Communications in Statistics: Simulation and Computation, 27 (3):843–856, 1998. CODEN CSSCDB. ISSN 0361-0918.

Hillis:1991:EEI

[Hil91] Stephen L. Hillis. Extending M-estimation to include censored data via James’s method. Communications in Statistics: Simulation and Computation, 20(1):121–128, 1991. CODEN CSSCDB. ISSN 0361-0918.

Hillis:1993:CTB

[Hil93] Stephen L. Hillis. A comparison of three Buckley-James variance estimators. Communications in Statistics: Simulation and Computation, 22(4):955–973, 1993. CODEN CSSCDB. ISSN 0361-0918.

Hillis:1994:HGS

[Hil94] Stephen L. Hillis. A heuristic generalization of Smith’s Buckley James variance estimator. Communications in Statistics: Simulation and Computation, 23(3):813–831, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Hirji:1997:CAE

[Hir97] Karim F. Hirji. A comparison of algorithms for exact goodnessof-fit tests for multinomial data. Communications in Statistics: Simulation and Computation, 26(3):1197–1227, 1997. CODEN CSSCDB. ISSN 0361-0918.

Hoerl:1990:RRD

[HK90a] Arthur E. Hoerl and Robert W. Kennard. Ridge regression: Degrees of freedom in the analysis of variance. Communications in Statistics: Simulation and Computation, 19(4):1485–1495, 1990. CODEN CSSCDB. ISSN 0361-0918.

Howard:1990:GLR

[HK90b] P. H. George Howard and Gary G. Koch. The GLM log-rank test: general linear modeling of log-rank scores as a method of analysis for survival data. Communications in Statistics: Simulation and Computation, 19(3):903–917, 1990. CODEN CSSCDB. ISSN 0361-0918.

Huda:1993:SOS

[HK93] S. Huda and I. H. Khan. On some A-optimal second-order designs over cubic regions. Communications in Statistics: Simulation and Computation, 22(1):99–115, 1993. CODEN CSSCDB. ISSN 0361-0918.

Handa:1993:MET

[HKT93] B. R. Handa, Rajeev Kumar, and R. K. Tuteja. A multiparameter extension of two stage modified shrinkage estimator. Communications in Statistics: Simulation and Computation, 22(2): 387–397, 1993. CODEN CSSCDB. ISSN 0361-0918.

Hsu:1990:ADR

[HL90] Chyi Hung Hsu and Martin S. Levy. An algorithm for the distribution of the ratio of generalized variances. Communications in Statistics: Simulation and Computation, 19(3):1127–1142, 1990. CODEN CSSCDB. ISSN 0361-0918.

Hackl:1992:NNQ

[HL92] Peter Hackl and Johannes Ledolter. A new nonparametric quality control technique. Communications in Statistics: Simulation and Computation, 21(2):423–443, 1992. CODEN CSSCDB. ISSN 0361-0918.

Horswell:1993:DLS

[HL93] Ronald L. Horswell and Stephen W. Looney. Diagnostic limitations of skewness coefficients in assessing departures from univariate and multivariate normality. Communications in Statistics: Simulation and Computation, 22(2):437–459, 1993. CO-DEN CSSCDB. ISSN 0361-0918.

Hapuarachchi:1992:APA

[HM92] K. P. Hapuarachchi and B. D. Macpherson. Autoregressive processes applied to acceptance sampling by variables. Communications in Statistics: Simulation and Computation, 21(3):833–848, 1992. CODEN CSSCDB. ISSN 0361-0918.

Harvill:1995:USM

[HN95] Jane L. Harvill and H. Joseph Newton. Using symbolic math to evaluate saddlepoint approximations for the difference of order statistics. Communications in Statistics: Simulation and Computation, 24(3):781–791, 1995. CODEN CSSCDB. ISSN 0361-0918.

Hsiung:1994:PPM

[HO94] Tung-Hsing Hsiung and Stephen Olejnik. Power of pairwise multiple comparisons in the unequal variance case. Communications in Statistics: Simulation and Computation, 23(3):691–710, 1994. CODEN CSSCDB. ISSN 0361-0918.

Hoekstra:1991:AVT

[Hoe91] J. A. Hoekstra. Approximate variances for tolerance-distribution free LD50-estimators supplemented with Abbott’s correction. Communications in Statistics: Simulation and Computation, 20 (2–3):715–723, 1991. CODEN CSSCDB. ISSN 0361-0918.

Haughton:1997:IOC

[HOJ97] Dominique M. A. Haughton, Johan H. L. Oud, and Robert A. R. G. Jansen. Information and other criteria in structural equation model selection. Communications in Statistics: Simulation and Computation, 26(4):1477–1516, 1997. CODEN CSSCDB. ISSN 0361-0918.

Hollas:1991:PBA

[Hol91] E. A. Hollas. Performance of Bartlett adjustment for certain likelihood ratio tests. Communications in Statistics: Simulation and Computation, 20(2–3):449–462, 1991. CODEN CSSCDB. ISSN 0361-0918.

Holmquist:1995:SMC

[Hol95] Björn Holmquist. Simulation of multimodal circular distributions of wrapped Cauchy type and of von Mises type. Communications in Statistics: Simulation and Computation, 24(3):607–615, 1995. CODEN CSSCDB. ISSN 0361-0918.

Horrace:1998:TPP

[Hor98] William C. Horrace. Tables of percentage points of the k-variate normal distribution for large values of k. Communications in Statistics: Simulation and Computation, 27(3):823–831, 1998. CODEN CSSCDB. ISSN 0361-0918.

Houshmand:1993:MPQ

[Hou93] Ali A. Houshmand. Misclassification probabilities for quadratic discriminant function. Communications in Statistics: Simulation and Computation, 22(1):81–98, 1993. CODEN CSSCDB. ISSN 0361-0918.

Hewett:1995:TCV

[HPL95] John E. Hewett, Gregory F. Petroski, and Yiqun Li. Tables of critical values for two group interim analyses. Communications in Statistics: Simulation and Computation, 24(4):825–852, 1995. CODEN CSSCDB. ISSN 0361-0918.

Hauck:1999:MSP

[HRM99] Daryl J. Hauck, George C. Runger, and Douglas C. Montgomery. Multivariate statistical process monitoring and diagnosis with grouped regression-adjusted variables. Communications in Statistics: Simulation and Computation, 28(2):309–328, 1999. CODEN CSSCDB. ISSN 0361-0918.

Hsu:1990:UFF

[HS90] Jason C. Hsu and W. C. Soong. Using the fast Fourier transform to compute multiple comparisons with the best and subset selection critical values. Communications in Statistics: Simulation and Computation, 19(4):1377–1391, 1990. CODEN CSSCDB. ISSN 0361-0918.

Harwell:1997:ESF

[HS97] Michael R. Harwell and Ronald C. Serlin. An empirical study of five multivariate tests for the single-factor repeated measures model. Communications in Statistics: Simulation and Computation, 26(2):605–618, 1997. CODEN CSSCDB. ISSN 0361-0918.

Hsieh:1994:ALT

[Hsi94] Hui K. Hsieh. Accelerated life test sampling plans for exponential distributions. Communications in Statistics: Simulation and Computation, 23(1):27–41, 1994. CODEN CSSCDB. ISSN 0361-0918.

Hung:1990:SSV

[Hun90] Ken Hung. Simulation study on variance of forecast error for vector ARIMA models. Communications in Statistics: Simulation and Computation, 19(1):125–144, 1990. CODEN CSSCDB. ISSN 0361-0918.

Hun:1998:SPR

[Hun98] Choi Young Hun. A study of the power of the rank transform test in a 2^{3} factorial experiment. Communications in Statistics: Simulation and Computation, 27(1):259–274, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Hutchinson:1990:SET

[Hut90] M. F. Hutchinson. A stochastic estimator of the trace of the influence matrix for Laplacian smoothing splines. Communications in Statistics: Simulation and Computation, 19(2):433–450, 1990. CODEN CSSCDB. ISSN 0361-0918.

Hillis:1996:IWL

[HW96] Stephen L. Hillis and Robert F. Woolson. An iterative weighted least squares algorithm and simulation study for censored data M-estimates. Communications in Statistics: Simulation and Computation, 25(1):247–262, 1996. CODEN CSSCDB. ISSN 0361-0918.

Hirst:1991:NRE

[HZBW91] Kathryn Hirst, Gary O. Zerbe, David W. Boyle, and Randall B. Wilkening. On nonlinear random effects models for repeated measurements. Communications in Statistics: Simulation and Computation, 20(2–3):463–478, 1991. CODEN CSSCDB. ISSN 0361-0918.

Jain:1991:ASS

[Jai91] Ram R. Jain. On the adequacy of sample size determination formulae for two independent proportions. Communications in Statistics: Simulation and Computation, 20(2–3):585–617, 1991. CODEN CSSCDB. ISSN 0361-0918.

Jain:1992:REP

[Jai92] Sudha Jain. Relative efficiency of a parameter for a M/G/1 queueing system based on reduced and full likelihood functions. Communications in Statistics: Simulation and Computation, 21 (2):597–606, 1992. CODEN CSSCDB. ISSN 0361-0918.

Jun:1993:SAR

[JC93] Chi-Hyuck Jun and Moon Soo Choi. Simulating the average run length for cusum schemes using variance reduction techniques. Communications in Statistics: Simulation and Computation, 22 (3):877–887, 1993. CODEN CSSCDB. ISSN 0361-0918.

Jones:1996:FCM

[JC96] Whitten Betty Jones and A. Clifford Cohen. Further considerations of modified moment estimators for Weibull and lognormal parameters. Communications in Statistics: Simulation and Computation, 25(3):749–784, 1996. CODEN CSSCDB. ISSN 0361-0918.

Chuang:1997:AND

[jCM97] Rwei ju Chuang and Nancy R. Mendell. The approximate null distribution of the likelihood ratio test for a mixture of two bivariate normal distributions with equal covariance. Communications in Statistics: Simulation and Computation, 26(2):631–648, 1997. CODEN CSSCDB. ISSN 0361-0918.

Jones:1995:DPP

[JD95] M. C. Jones and F. Daly. Density probability plots. Communications in Statistics: Simulation and Computation, 24(4):911–927, 1995. CODEN CSSCDB. ISSN 0361-0918.

Johannesson:1995:AIB

[JG95] Benedikt Jóhannesson and Narayan Giri. On approximations involving the beta distribution. Communications in Statistics: Simulation and Computation, 24(2):489–503, 1995. CODEN CSSCDB. ISSN 0361-0918.

Jung:1998:ACO

[JK98] Jin-Whan Jung and Gary G. Koch. Analysis of correlated ordered categorical data with rank measures of association for responses with ties. Communications in Statistics: Simulation and Computation, 27(1):167–183, 1998. CODEN CSSCDB. ISSN 0361-0918.

Lasinio:1999:SIW

[JLPR99] G. Jona-Lasinio, M. Piccioni, and A. Ramponi. Selection of importance weights for Monte Carlo estimation of normalizing constants. Communications in Statistics: Simulation and Computation, 28(2):441–462, 1999. CODEN CSSCDB. ISSN 0361-0918.

Jothikumar:1996:TSC

[JR96] J. Jothikumar and C. Raju. Two stage chain sampling plans ChSP-(0,2) and ChSP-(1,2) — Part 2. Communications in Statistics: Simulation and Computation, 25(3):817–834, 1996. CODEN CSSCDB. ISSN 0361-0918.

Joseph:1990:ECR

[JW90] Lawrence Joseph and David B. Wolfson. Estimating the change in a renewal process when the data are counts. Communications in Statistics: Simulation and Computation, 19(4):1431– 1441, 1990. CODEN CSSCDB. ISSN 0361-0918.

Jeong:1995:TCL

[JY95] Hyean-Seok Jeong and Bong-Jin Yum. Type-I censored life test plans in the exponential case. Communications in Statistics: Simulation and Computation, 24(1):187–205, 1995. CODEN CSSCDB. ISSN 0361-0918.

Jang:1997:GME

[JY97] Dae Heung Jang and Min Yoon. Graphical methods for evaluating ridge regression estimator in mixture experiments. Communications in Statistics: Simulation and Computation, 26(3): 1049–1061, 1997. CODEN CSSCDB. ISSN 0361-0918.

Kim:1999:EDF

[KA99] YongHee Kim and Barry C. Arnold. Estimation of a distribution function under generalized ranked set sampling. Communications in Statistics: Simulation and Computation, 28(3):657–666, 1999. CODEN CSSCDB. ISSN 0361-0918.

Keselman:1998:CTA

[KAKW98] H. J. Keselman, James Algina, Rhonda K. Kowalchuk, and Russell D. Wolfinger. A comparison of two approaches for selecting covariance structures in the analysis of repeated measurements. Communications in Statistics: Simulation and Computation, 27 (3):591–604, 1998. CODEN CSSCDB. ISSN 0361-0918.

Kandasamy:1993:PTC

[Kan93] C. Kandasamy. Procedures and tables for construction and selection of CSP-2 plans. Communications in Statistics: Simulation and Computation, 22(1):285–303, 1993. CODEN CSSCDB. ISSN 0361-0918.

Karanta:1993:OMT

[Kar93] Ilkka Karanta. Optimization methods in time series interpolation. Communications in Statistics: Simulation and Computation, 22(4):1181–1203, 1993. CODEN CSSCDB. ISSN 0361-0918.

Kamel:1990:CNC

[KAS90] Dr. Ashour Samir Kamel and Ali I. Abdel-Samad. On the computation of non-central chi-square distribution function. Communications in Statistics: Simulation and Computation, 19(4): 1279–1291, 1990. CODEN CSSCDB. ISSN 0361-0918.

Kao:1995:SSN

[KC95] Nancy L. Kao and S. Chakraborti. Small sample null distribution of two nonparametric tests for comparing treatments with a control in a two-way layout. Communications in Statistics: Simulation and Computation, 24(1):149–164, 1995. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1996:RTM

[KC96] Peter E. Kennedy and Brain S. Cade. Randomization tests for multiple regression. Communications in Statistics: Simulation and Computation, 25(4):923–936, 1996. CODEN CSSCDB. ISSN 0361-0918.

Karian:1999:FGL

[KD99] Zaven A. Karian and Edward J. Dudewicz. Fitting the generalized lambda distribution to data: a method based on percentiles. Communications in Statistics: Simulation and Computation, 28 (3):793–819, 1999. CODEN CSSCDB. ISSN 0361-0918.

Karian:1996:EGL

[KDM96] Zaven A. Karian, Edward J. Dudewicz, and Patrick McDonald. The extended generalized lambda distribution system for fitting distributions to data: history, completion of theory, tables, applications, the “final word” on moment fits. Communications in Statistics: Simulation and Computation, 25(3):611–642, 1996. CODEN CSSCDB. ISSN 0361-0918.

Korukoglu:1996:NST

[KE96] Serdar Korukoğlu and Kadir Ertas. A new statistic for testing uniformity of ranking data. Communications in Statistics: Simulation and Computation, 25(1):139–156, 1996. CODEN CSS-CDB. ISSN 0361-0918.

Kenyon:1990:BAP

[Ken90] James R. Kenyon. Bounds and approximations for pairwise comparisons of independent multinormal means. Communications in Statistics: Simulation and Computation, 19(2):555–590, 1990. CODEN CSSCDB. ISSN 0361-0918.

Koehler:1990:CSG

[KG90] Kenneth J. Koehler and F. F. Gan. Chi-squared goodness-of-fit tests: Cell selection and power. Communications in Statistics: Simulation and Computation, 19(4):1265–1278, 1990. CODEN CSSCDB. ISSN 0361-0918.

Kazempour:1991:ACB

[KG91] M. K. Kazempour and Franklin A. Graybill. Approximate confidence bounds for ratios of variance components in two-way unbalanced crossed models with interactions. Communications in Statistics: Simulation and Computation, 20(4):955–967, 1991. CODEN CSSCDB. ISSN 0361-0918.

Kandasamy:1993:SCP

[KG93] C. Kandasamy and K. Govindaraju. Selection of CSP-T plans. Communications in Statistics: Simulation and Computation, 22 (1):265–283, 1993. CODEN CSSCDB. ISSN 0361-0918.

Katti:1992:ATD

[KH92] S. K. Katti and Ferrin Hamson. An algorithm to test divisibility of continuous distributions, applied to the uniform distribution. Communications in Statistics: Simulation and Computation, 21 (2):587–596, 1992. CODEN CSSCDB. ISSN 0361-0918.

Klockars:1996:PSI

[KH96] Alan J. Klockars and Gregory R. Hancock. Power of a stagewise intersection protected multiple comparison procedure. Communications in Statistics: Simulation and Computation, 25(4):953– 960, 1996. CODEN CSSCDB. ISSN 0361-0918.

Khamis:1997:OSS

[Kha97] Imad H. Khamis. Optimum M-step, step-stress design with K stress variables. Communications in Statistics: Simulation and Computation, 26(4):1301–1313, 1997. CODEN CSSCDB. ISSN 0361-0918.

Kim:1993:BBI

[KHW93] Yun Bae Kim, Jorge Haddock, and Thomas R. Willemain. The binary bootstrap: inference with autocorrelated binary data. Communications in Statistics: Simulation and Computation, 22 (1):205–216, 1993. CODEN CSSCDB. ISSN 0361-0918.

Kim:1994:AGS

[Kim94] Hyune-Ju Kim. Approximations in a group sequential test with unequal group sizes. Communications in Statistics: Simulation and Computation, 23(4):915–924, 1994. CODEN CSSCDB. ISSN 0361-0918.

Kim:1999:ELR

[Kim99] Ji-Hyun Kim. Extension of the log-rank test. Communications in Statistics: Simulation and Computation, 28(4):1099–1112, 1999. CODEN CSSCDB. ISSN 0361-0918.

Krehbiel:1998:QLT

[KJ98] Timothy C. Krehbiel and J. Marcus Jobe. Quality of the log transformation for an incorrectly assumed error structure. Communications in Statistics: Simulation and Computation, 27(3): 605–616, 1998. CODEN CSSCDB. ISSN 0361-0918.

Kotz:1999:DRA

[KJ99] Samuel Kotz and Norman L. Johnson. Delicate relations among the basic process capability indices c_{p}, c_{pk} AND c_{pm}, AND THEIR MODIFICATIONS. Communications in Statistics: Simulation and Computation, 28(3):849–866, 1999. CODEN CSS-CDB. ISSN 0361-0918.

Kocherlakota:1991:DND

[KK91] Kathleen Kocherlakota and Subrahmaniam Kocherlakota. On the doubly noncentral t distribution. Communications in Statistics: Simulation and Computation, 20(1):23–31, 1991. CODEN CSSCDB. ISSN 0361-0918.

Kundu:1997:CML

[KK97] Debasis Kundu and Nandini Kannan. Constrained maximum likelihood estimators for superimposed exponential signals. Communications in Statistics: Simulation and Computation, 26(2): 733–764, 1997. CODEN CSSCDB. ISSN 0361-0918.

Kocherlakota:1999:ACN

[KK99] Subrahmaniam Kocherlakota and Kathleen Kocherlakota. Approximations for central and noncentral bivariate chi-square distributions. Communications in Statistics: Simulation and Computation, 28(4):909–930, 1999. CODEN CSSCDB. ISSN 0361-0918.

King:1990:CDW

[KL90] Dennis W. King and Michael T. Longnecker. Computing the distribution of the Wilcoxon statistic with applications to process control. Communications in Statistics: Simulation and Computation, 19(2):381–399, 1990. CODEN CSSCDB. ISSN 0361-0918.

Kadiyala:1990:ESE

[KO90] K. R. Kadiyala and Dennis Oberhelman. Estimation of standard errors of empirical Bayes estimators in CAPM-type models. Communications in Statistics: Simulation and Computation, 19 (1):189–206, 1990. CODEN CSSCDB. ISSN 0361-0918.

Kozumi:1990:SSP

[Koz90] Hideo Kozumi. On small sample properties of the Wald, LR and LM tests in a linear model with AR(1) errors. Communications in Statistics: Simulation and Computation, 19(4):1361–1375, 1990. CODEN CSSCDB. ISSN 0361-0918.

Koreisha:1990:LME

[KP90] Sergio Koreisha and Tarmo Pukkila. Linear methods for estimating ARMA and regression models with serial correlation. Communications in Statistics: Simulation and Computation, 19(1): 71–102, 1990. CODEN CSSCDB. ISSN 0361-0918.

Kishan:1991:COT

[KPA91] Mehrotra Kishan, Jackson Paul, and Schick Anton. On choosing an optimally trimmed mean. Communications in Statistics: Simulation and Computation, 20(1):73–80, 1991. CODEN CSSCDB. ISSN 0361-0918.

Kumar:1998:ATG

[KR98] M. Pran Kumar and C. V. Rao. ANOM-type graphical method for testing the equality of several variances. Communications in Statistics: Simulation and Computation, 27(2):459–468, 1998. CODEN CSSCDB. ISSN 0361-0918.

Krouse:1997:PPI

[Kro97] Donal P. Krouse. Projection pursuit by involution. Communications in Statistics: Simulation and Computation, 26(4):1517– 1546, 1997. CODEN CSSCDB. ISSN 0361-0918.

Kim:1996:RVC

[KS96] Choongrak Kim and Barry E. Storer. Reference values for Cook’s distance. Communications in Statistics: Simulation and Computation, 25(3):691–708, 1996. CODEN CSSCDB. ISSN 0361-0918.

Ku:1998:PES

[KS98] S. Ku and E. Seneta. Practical estimation from the sum of AR(1) processes. Communications in Statistics: Simulation and Computation, 27(4):981–998, 1998. CODEN CSSCDB. ISSN 0361-0918.

Krumbholz:1996:MKC

[KSS96] Wolf Krumbholz, Martin Schader, and Friedrich Schmid. Modifications of the Kolmogorov, Cramér–von Mises, and Watson tests for testing uniformity with unknown limits. Communications in Statistics: Simulation and Computation, 25(4):1093–1104, 1996. CODEN CSSCDB. ISSN 0361-0918.

Kulasekera:1992:NDD

[KT92] K. B. Kulasekera and David W. Tonkyn. A new discrete distrlbution, with applications to survival, dispersal and dispersion. Communications in Statistics: Simulation and Computation, 21 (2):499–518, 1992. CODEN CSSCDB. ISSN 0361-0918.

Kumar:1992:PES

[KTTB92] Santosh Kumar, Ravinder K. Tyagi, Ram C. Tiwari, and Samir K. Bhattacharya. Pretest estimation for some discrete distributions. Communications in Statistics: Simulation and Computation, 21(2):405–421, 1992. CODEN CSSCDB. ISSN 0361-0918.

Kent:1994:CLI

[KTV94] John T. Kent, David E. Tyler, and Yahuda Vard. A curious likelihood identity for the multivariate t-distribution. Communications in Statistics: Simulation and Computation, 23(2):441–453, 1994. CODEN CSSCDB. ISSN 0361-0918.

Kim:1996:QPA

[KUK96] Hyuk Joo Kim, Yonghwan Um, and Andre I. Khuri. Quantile plots of the average slope variance for response surface designs. Communications in Statistics: Simulation and Computation, 25 (4):995–1014, 1996. CODEN CSSCDB. ISSN 0361-0918.

Kulperger:1996:BED

[Kul96] R. J. Kulperger. Bootstrapping empirical distribution functions of residuals from autoregressive model fitting. Communications in Statistics: Simulation and Computation, 25(3):657–670, 1996. CODEN CSSCDB. ISSN 0361-0918.

Kulasekera:1996:ESF

[KW96a] K. B. Kulasekera and William H.. White. Estimation of the survival function from censored data:a method based on total time on test. Communications in Statistics: Simulation and Computation, 25(1):189–200, 1996. CODEN CSSCDB. ISSN 0361-0918.

Kwan:1996:CSF

[KW96b] Andy C. C. Kwan and Yangru Wu. A comparative study of the finite-sample distribution of some portmanteau tests for univariate time series models. Communications in Statistics: Simulation and Computation, 25(4):867–904, 1996. CODEN CSSCDB. ISSN 0361-0918.

Keselman:1999:IBF

[KW99] H. J. Keselman and Rand R. Wilcox. The ‘improved’ Brown and Forsythe test for mean equality: some things can’t be fixed. Communications in Statistics: Simulation and Computation, 28 (3):687–698, 1999. CODEN CSSCDB. ISSN 0361-0918.

Lavine:1992:LPI

[Lav92] Michael Lavine. Local predictive influence in Bayesian linear models with conjugate priors. Communications in Statistics: Simulation and Computation, 21(1):269–283, 1992. CODEN CSSCDB. ISSN 0361-0918.

Lowry:1995:PCC

[LCW95] Cynthia A. Lowry, Charles W. Champ, and William H. Woodall. The performance of control charts for monitoring process variation. Communications in Statistics: Simulation and Computation, 24(2):409–437, 1995. CODEN CSSCDB. ISSN 0361-0918.

Lee:1992:OSS

[Lee92] Austin F. S. Lee. Optimal sample sizes determined by two-sample Welch’s t test. Communications in Statistics: Simulation and Computation, 21(3):689–696, 1992. CODEN CSSCDB. ISSN 0361-0918.

Lee:1996:EPS

[Lee96] Shen-Ming Lee. Estimating population size for capture-recapture data when capture probabilities vary by time, behavior and individual animal. Communications in Statistics: Simulation and Computation, 25(2):431–457, 1996. CODEN CSSCDB. ISSN 0361-0918.

Lee:1997:AWB

[LEKR97] Choon Hwan Lee, Bilal M. Ezzeddine, Choongrak Kim, and Robert T. Ross. An algorithm for weighted bilinear regression.

Communications in Statistics: Simulation and Computation, 26 (2):791–804, 1997. CODEN CSSCDB. ISSN 0361-0918.

Leung:1994:LLD

[Leu94] Chi-Ying Leung. The location linear discriminant with covariates. Communications in Statistics: Simulation and Computation, 23(4):1027–1046, 1994. CODEN CSSCDB. ISSN 0361-0918.

London:1999:SMG

[LG99] Wendy B. London and Chris Gennings. Simulation of multivariate gamma data with exponential marginals for independent clusters. Communications in Statistics: Simulation and Computation, 28(2):487–500, 1999. CODEN CSSCDB. ISSN 0361-0918.

Lewsey:1997:SSU

[LGG97] James D. Lewsey, William P. Gardiner, and George Gettinby. A study of simple unbalanced factorial designs that use Type II and Type III sums of squares. Communications in Statistics: Simulation and Computation, 26(4):1315–1328, 1997. CODEN CSSCDB. ISSN 0361-0918.

Lai:1997:NSA

[LH97] Dejian Lai and Ronald B. Harrist. A nonparametric statistical approach in noisy chaos identification. Communications in Statistics: Simulation and Computation, 26(1):291–300, 1997. CODEN CSSCDB. ISSN 0361-0918.

Li:1996:ENS

[Li96] Chao Li. Estimation of non-stationary spectra by simulated annealing. Communications in Statistics: Simulation and Computation, 25(2):381–402, 1996. CODEN CSSCDB. ISSN 0361-0918.

Lim:1995:ADC

[Lim95] K. S. Lim. An algorithm for determining the chaos of TAR(L). Communications in Statistics: Simulation and Computation, 24 (2):505–522, 1995. CODEN CSSCDB. ISSN 0361-0918.

Lim:1999:SSC

[Lim99] K. S. Lim. A simulation study on the chaotic SETAR(2; 1, 1) and the local Lyapunov exponent. Communications in Statistics: Simulation and Computation, 28(1):147–163, 1999. CO-DEN CSSCDB. ISSN 0361-0918.

Lin:1999:BLT

[Lin99] Darning Lin. A Bayesian life test acceptance sampling plan by variables with cost considerations. Communications in Statistics: Simulation and Computation, 28(4):1113–1133, 1999. CODEN CSSCDB. ISSN 0361-0918.

Lee:1990:EAL

[LK90] Byung Gook Lee and Byung Chun Kim. An efficient algorithm for the least-squares cross-validation with symmetric and polynomial kernels. Communications in Statistics: Simulation and Computation, 19(4):1513–1522, 1990. CODEN CSSCDB. ISSN 0361-0918.

Lee:1997:DBS

[LK97] Kang In Lee and John J. Koval. Determination of the best significance level in forward stepwise logistic regression. Communications in Statistics: Simulation and Computation, 26(2):559–575, 1997. CODEN CSSCDB. ISSN 0361-0918.

Landry:1992:EBS

[LL92] L. Landry and Y. Lepage. Empirical behavior of some tests for normality. Communications in Statistics: Simulation and Computation, 21(4):971–999, 1992. CODEN CSSCDB. ISSN 0361-0918.

Lawoko:1994:EMP

[LM94] C. R. O. Lawoko and G. J. McLachlan. Estimation of mixing proportions in the presence of autoregressively correlated training data: the case of two univariate normal populations. Communications in Statistics: Simulation and Computation, 23(3): 591–613, 1994. CODEN CSSCDB. ISSN 0361-0918.

Lieberman:1997:AEN

[LM97] Offer Lieberman and Laszlo Matyas. Approximate estimation in nonlinear panel data models. Communications in Statistics: Simulation and Computation, 26(3):1177–1195, 1997. CODEN CSSCDB. ISSN 0361-0918.

Leconte:1994:TSP

[LML94] Eve Leconte, Thierry Moreau, and Joseph Lellouch. The twosample problem with multivariate censored data: a new rank test family. Communications in Statistics: Simulation and Computation, 23(4):1061–1076, 1994. CODEN CSSCDB. ISSN 0361-0918.

Lasser:1994:CEM

[LNZ94] Rupert Lasser, Michael Niebner, and Hilmar Zock. Comparison of the efficiency for mean estimators in time series. Communications in Statistics: Simulation and Computation, 23(1):129–141, 1994. CODEN CSSCDB. ISSN 0361-0918.

Li:1990:LCL

[LOSB90] Huaixiang Li, D. B. Owen, A. Salvador, and A. Borrego. Lower confidence limits on process capability indices based on the range. Communications in Statistics: Simulation and Computation, 19 (1):1–24, 1990. CODEN CSSCDB. ISSN 0361-0918.

Lee:1998:TVS

[LS98] S. J. Lee and Ping Sa. Testing the variance of skewed distributions. Communications in Statistics: Simulation and Computation, 27(3):807–822, 1998. CODEN CSSCDB. ISSN 0361-0918.

Larocque:1995:DOT

[LT95] Denis Larocque and Serge Tardif. Development of an optimal test based on weighted rankings. Communications in Statistics: Simulation and Computation, 24(2):367–383, 1995. CODEN CSS-CDB. ISSN 0361-0918.

Lawal:1995:AFM

[LU95] H. Bayo Lawal and G. J. G. Upton. An algorithm for fitting models to n×n contingency tables having ordered categories. Communications in Statistics: Simulation and Computation, 24(3): 793–805, 1995. CODEN CSSCDB. ISSN 0361-0918.

Luceno:1994:SCE

[Luc94] Alberto Luceño. Speed of covergence to the extreme value distributions on their probability plotting papers. Communications in Statistics: Simulation and Computation, 23(2):529–545, 1994. CODEN CSSCDB. ISSN 0361-0918.

Luceno:1996:PCI

[Luc96] Alberto Luceño. A process capability index with reliable confidence intervals. Communications in Statistics: Simulation and Computation, 25(1):235–245, 1996. CODEN CSSCDB. ISSN 0361-0918.

Luceno:1997:PEC

[Luc97] Alberto Luceño. Parameter estimation with closed-loop operating data under time varying discrete proportional-integral control. Communications in Statistics: Simulation and Computation, 26(1):215–232, 1997. CODEN CSSCDB. ISSN 0361-0918.

Lund:1999:CAD

[Lun99] Ulric Lund. Cluster analysis for directional data. Communications in Statistics: Simulation and Computation, 28(4):1001– 1009, 1999. CODEN CSSCDB. ISSN 0361-0918.

Lewellen:1999:AFT

[LV99] Ruth H. Lewellen and Stephen H. Vessey. Analysis of fragmented time series data using Box–Jenkins models. Communications in Statistics: Simulation and Computation, 28(3):667–685, 1999. CODEN CSSCDB. ISSN 0361-0918.

Li:1993:RSW

[LW93] Kim-Hung Li and Carl Ka-Fai Wong. Random sampling from the Watson distribution. Communications in Statistics: Simulation and Computation, 22(4):997–1009, 1993. CODEN CSSCDB. ISSN 0361-0918.

Lei:1994:TAC

[LW94] Xingye Lei and F. T. Wright. Testing for and against a concavity restriction with normal errors. Communications in Statistics: Simulation and Computation, 23(3):711–726, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Lee:1996:MDC

[LW96] Kyung Yul Lee and Lisa A. Weissfeld. A multicollinearity diagnostic for the Cox model with time dependent covariates. Communications in Statistics: Simulation and Computation, 25(1): 41–60, 1996. CODEN CSSCDB. ISSN 0361-0918.

Leu:1999:DBA

[LW99] Lii-Yuh Leu and Ji-Cheng Wu. A dynamic Bayesian approach to inference from accelerated life tests. Communications in Statistics: Simulation and Computation, 28(1):37–49, 1999. CODEN CSSCDB. ISSN 0361-0918.

Lee:1991:MSN

[LWC91] Sanghoon Lee, James R. Wilson, and Melba M. Crawford. Modeling and simulation of a nonhomogeneous Poisson process having cyclic behavior. Communications in Statistics: Simulation and Computation, 20(2–3):777–809, 1991. CODEN CSSCDB. ISSN 0361-0918.

Lu:1994:RSE

[LY94] Rueypyng Lu and Cheng-Hong Yang. Resampling schemes for estimating the similarity measures on location model. Communications in Statistics: Simulation and Computation, 23(4):973–996, 1994. CODEN CSSCDB. ISSN 0361-0918.

Lye:1993:FPD

[Lye93] Jenny N. Lye. A flexible parametric density estimator for multimodal distributions of test statistics. Communications in Statistics: Simulation and Computation, 22(3):813–830, 1993. CO-DEN CSSCDB. ISSN 0361-0918.

Meneses:1990:SAD

[MABL90] Jesse Meneses, C. E. Antle, Mary J. Bartholomew, and R. L. Lengerich. A simple algorithm for delta method variances for multinomial posterior Bayes probability estimates. Communications in Statistics: Simulation and Computation, 19(3):837–845, 1990. CODEN CSSCDB. ISSN 0361-0918.

MacEachern:1994:ENM

[Mac94] Steven N. MacEachern. Estimating normal means with a conjugate style Dirichlet process prior. Communications in Statistics: Simulation and Computation, 23(3):727–741, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Madi:1997:BSE

[Mad97] Mohamed T. Madi. Bayes and Stein estimation under asymmetric loss functions: a numerical risk comparison. Communications in Statistics: Simulation and Computation, 26(1):53–66, 1997. CODEN CSSCDB. ISSN 0361-0918.

Magel:1990:REO

[Mag90] Rhonda C. Magel. Robust estimates of ordered parameters with right censored observations. Communications in Statistics: Simulation and Computation, 19(1):207–223, 1990. CODEN CSS-CDB. ISSN 0361-0918.

Magee:1991:EAT

[Mag91] Lonnie Magee. Edgeworth-adjusting test statistics for AR(1) errors. Communications in Statistics: Simulation and Computation, 20(4):901–917, 1991. CODEN CSSCDB. ISSN 0361-0918.

Marks:1998:MKS

[Mar98] Neil B. Marks. Modification of the Kolmogorov–Smirnov test for the Erlang-2 distribution. Communications in Statistics: Simulation and Computation, 27(1):39–49, 1998. CODEN CSSCDB. ISSN 0361-0918.

McCool:1996:ATW

[McC96] John I. McCool. The analysis of a two way factorial design with Weibull response. Communications in Statistics: Simulation and Computation, 25(1):263–286, 1996. CODEN CSSCDB. ISSN 0361-0918.

McNulty:1991:TAN

[ME91] Mark S. McNulty and William Eisele. Testing for autocorrelation in nested analysis of variance. Communications in Statistics: Simulation and Computation, 20(2–3):399–410, 1991. CODEN CSSCDB. ISSN 0361-0918.

Meeker:1994:ACC

[ME94] William Q. Meeker and Luis A. Escobar. An algorithm to compute the cdf of the product of two normal random variables. Communications in Statistics: Simulation and Computation, 23 (1):271–280, 1994. CODEN CSSCDB. ISSN 0361-0918.

Mehrotra:1997:IBF

[Meh97] Devan V. Mehrotra. Improving the Brown–Forsythe solution to the generalized Behrens–Fisher problem. Communications in Statistics: Simulation and Computation, 26(3):1139–1145, 1997. CODEN CSSCDB. ISSN 0361-0918.

Mengersen:1992:RNS

[Men92] Kerrie Mengersen. Robustness to normality of a selection rule. Communications in Statistics: Simulation and Computation, 21 (1):35–56, 1992. CODEN CSSCDB. ISSN 0361-0918.

Murtaugh:1993:REM

[MF93] Paul A. Murtaugh and Lloyd D. Fisher. Random-effects model of dependence between efficacy and toxicity in dose-ranging trials. Communications in Statistics: Simulation and Computation, 22 (2):507–522, 1993. CODEN CSSCDB. ISSN 0361-0918.

Mudholkar:1998:HQC

[MH98] Govind S. Mudholkar and Alan D. Hutson. Hermitian quantile curves as statistical plots. Communications in Statistics: Simulation and Computation, 27(3):783–805, 1998. CODEN CSSCDB. ISSN 0361-0918.

Miller:1991:USR

[Mil91] G. Edward Miller. Use of the squared ranks test to test for the equality of the coefficients of variation. Communications in Statistics: Simulation and Computation, 20(2–3):743–750, 1991. CODEN CSSCDB. ISSN 0361-0918.

Mitchell:1993:ECS

[Mit93] Heather Mitchell. Estimating convex and strictly convex functions. Communications in Statistics: Simulation and Computation, 22(3):803–811, 1993. CODEN CSSCDB. ISSN 0361-0918.

May:1997:PSC

[MJ97] Warren L. May and William D. Johnson. Properties of simultaneous confidence intervals for multinomial proportions. Communications in Statistics: Simulation and Computation, 26(2): 495–518, 1997. CODEN CSSCDB. ISSN 0361-0918.

Mazumdar:1999:TMM

[MLA^{+}99] Sati Mazumdar, Kenneth Liu, Sang Ahnn, Patricia R. Houck, and Charles F. Reynolds III. Transition (Markov) models for the analysis of survival times in clinical research. Communications in Statistics: Simulation and Computation, 28(1):165–176, 1999. CODEN CSSCDB. ISSN 0361-0918.

McLachlan:1995:ALR

[MMM95] G. J. McLachlan, C. E. McLaren, and D. Matthews. An algorithm for the likelihood ratio test of one versus two components in a normal mixture model fitted to grouped and truncated data. Communications in Statistics: Simulation and Computation, 24 (4):965–985, 1995. CODEN CSSCDB. ISSN 0361-0918.

Misra:1991:FND

[MN91] B. L. Misra and Bhagwandas Nutan. Families of neighbour designs and their analysis. Communications in Statistics: Simulation and Computation, 20(2–3):427–436, 1991. CODEN CSS-CDB. ISSN 0361-0918.

Mudholkar:1999:AIG

[MN99] Govind S. Mudholkar and Rajeshwari Natarajan. Approximations for the inverse Gaussian probabilities and percentiles. Communications in Statistics: Simulation and Computation, 28(4): 1051–1071, 1999. CODEN CSSCDB. ISSN 0361-0918.

Morel:1992:SAG

[Mor92] Jorge G. Morel. A simple algorithm for generating multinomial random vectors with extravariation. Communications in Statistics: Simulation and Computation, 21(4):1255–1268, 1992. CO-DEN CSSCDB. ISSN 0361-0918.

Morel:1999:CMA

[Mor99] Jorge G. Morel. A covariance matrix that accounts for different degrees of extraneous variation in multinomial responses. Communications in Statistics: Simulation and Computation, 28(2): 403–413, 1999. CODEN CSSCDB. ISSN 0361-0918.

Moser:1997:HTQ

[MP97] Barry Kurt Moser and Mark E. Payton. Hypothesis testing for quantiles of two-parameter binary models. Communications in Statistics: Simulation and Computation, 26(2):487–494, 1997. CODEN CSSCDB. ISSN 0361-0918.

Morais:1998:TSP

[MP98] Manuel Cabral Morais and Antònio Pacheco. Two stochastic properties of one-sided exponentially weighted moving average control charts*. Communications in Statistics: Simulation and Computation, 27(4):937–952, 1998. CODEN CSSCDB. ISSN 0361-0918.

Mosimann:1996:UOD

[MR96] James E. Mosimann and Makarand V. Ratnaparkhi. Uniform occurrence of digits for folded and mixture distributions on finite intervals. Communications in Statistics: Simulation and Computation, 25(2):481–506, 1996. CODEN CSSCDB. ISSN 0361-0918.

Menendez:1991:PTT

[MRS91] J. A. Menéndez, C. Rueda, and B. Salvador. Properties of two tests for outliers in multivariate data. Communications in Statistics: Simulation and Computation, 20(2–3):751–763, 1991. CO-DEN CSSCDB. ISSN 0361-0918.

Moors:1998:NPV

[MS98] J. J. A. Moors and L. W. G. Strijbosch. New proposals for the validation of trace-driven simulations. Communications in Statistics: Simulation and Computation, 27(4):1051–1073, 1998. CODEN CSSCDB. ISSN 0361-0918.

Murthi:1993:RNB

[MST93] B. P. S. Murthi, Kannan Srinivasan, and Pandu R. Tadikamalla. Robustness of the negative binomial distribution: a simulation study. Communications in Statistics: Simulation and Computation, 22(2):399–420, 1993. CODEN CSSCDB. ISSN 0361-0918.

Macdonald:1991:CWJ

[MT91] J. Ross Macdonald and William J. Thompson. The collected works of John W. Tukey. Communications in Statistics: Simulation and Computation, 20(4):843–886, 1991. CODEN CSSCDB. ISSN 0361-0918.

Moss:1990:QAD

[MTT90] Linda L. Crawford Moss, Malcolm S. Taylor, and Henry B. Tingey. Quant1les of the Anderson–Darling statistic. Communications in Statistics: Simulation and Computation, 19(3):1007– 1014, 1990. CODEN CSSCDB. ISSN 0361-0918.

Mudholkar:1998:GAW

[MTvdM98] Govind S. Mudholkar, Madhu C. Trivedi, and Edward C. van der Meulen. A gamma analogue of the Wilson–Hilferty transformation. Communications in Statistics: Simulation and Computation, 27(1):21–37, 1998. CODEN CSSCDB. ISSN 0361-0918.

Murtaugh:1998:MVS

[Mur98] Paul A. Murtaugh. Methods of variable selection in regression modeling. Communications in Statistics: Simulation and Computation, 27(3):711–734, 1998. CODEN CSSCDB. ISSN 0361-0918.

Muruzabal:1999:NBS

[Mur99] Jorge Muruzábal. A note on the behavior of the Stahel–Donoho estimator for outlier detection. Communications in Statistics: Simulation and Computation, 28(1):207–222, 1999. CODEN CSSCDB. ISSN 0361-0918.

Moonesinghe:1994:LRT

[MW94] Ramal Moonesinghe and F. T. Wright. Likelihood ratio tests involving a bivariate trend in two-factor designs: the level probabilities. Communications in Statistics: Simulation and Computation, 23(1):143–155, 1994. CODEN CSSCDB. ISSN 0361-0918.

McDonald:1992:EIL

[MX92] James B. McDonald and Yexiao Xu. An empirical investigation of the likelihood ratio test when the boundary condition is violated. Communications in Statistics: Simulation and Computation, 21 (3):879–892, 1992. CODEN CSSCDB. ISSN 0361-0918.

Montepiedra:1998:TSS

[MY98] Grace Montepiedra and Arthur B. Yeh. A two-stage strategy for the construction of D-optimal experimental designs. Communications in Statistics: Simulation and Computation, 27(2): 377–401, 1998. CODEN CSSCDB. ISSN 0361-0918.

Mansouri:1991:SSD

[MZ91] H. Mansouri and X. Zheng. Small sample distributions of some nonparametric tests for testing treatments vs control. Communications in Statistics: Simulation and Computation, 20(4):969– 979, 1991. CODEN CSSCDB. ISSN 0361-0918.

Nagaraj:1990:LOI

[Nag90] Neerchal K. Nagaraj. Locally optimal invariant tests for change in level. Communications in Statistics: Simulation and Computation, 19(3):869–878, 1990. CODEN CSSCDB. ISSN 0361-0918.

Naito:1996:WSE

[Nai96] Kanta Naito. On weighting the Studentized empirical characteristic function for testing normality. Communications in Statistics: Simulation and Computation, 25(1):201–213, 1996. CO-DEN CSSCDB. ISSN 0361-0918.

Narayanan:1991:SSP

[Nar91] A. Narayanan. Small sample properties of parameter estimation in the Dirichlet distribution. Communications in Statistics: Simulation and Computation, 20(2–3):647–666, 1991. CODEN CSSCDB. ISSN 0361-0918.

Nuno:1997:PUO

[NdL97] Crato Nuno and Pedro J. F. de Lima. On the power of underdifferencing and overdifferencing tests against nearly nonstationary alternatives. Communications in Statistics: Simulation and Computation, 26(4):1431–1446, 1997. CODEN CSSCDB. ISSN 0361-0918.

Natarajan:1998:HBQ

[NGM98] Kannan Natarajan, Malay Ghosh, and Tapabrata Maiti. Hierarchical Bayes quality measurement plan. Communications in Statistics: Simulation and Computation, 27(1):199–214, 1998. CODEN CSSCDB. ISSN 0361-0918.

Nguyen:1993:ACO

[Ngu93] Nam-Ky Nguyen. An algorithm for constructing optimal resolvable incomplete block designs. Communications in Statistics: Simulation and Computation, 22(3):911–923, 1993. CO-DEN CSSCDB. ISSN 0361-0918.

Nakao:1997:ACR

[NH97] Hiroyuki Nakao and Hiroto Hyakutake. Approximated confidence regions in multivariate linear calibration. Communications in Statistics: Simulation and Computation, 26(3):829–839, 1997. CODEN CSSCDB. ISSN 0361-0918.

Nickerson:1993:ALC

[Nic93] David M. Nickerson. Another look at counting by weighing. Communications in Statistics: Simulation and Computation, 22(2): 323–343, 1993. CODEN CSSCDB. ISSN 0361-0918.

Nishina:1997:AAA

[Nis97] Ken Nishina. An alternative approach for analysis of performance of one-sided cusum charts with variable sampling intervals. Communications in Statistics: Simulation and Computation, 26(4): 1329–1345, 1997. CODEN CSSCDB. ISSN 0361-0918.

Novo:1998:GFT

[NM98] L. A. Ramil Novo and W. González Manteiga. χ^{2} goodness-of-fit tests for polynomial regression. Communications in Statistics: Simulation and Computation, 27(1):229–258, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Nolan:1995:CMS

[NR95] John P. Nolan and Balram Rajput. Calculation of multidimensional stable densities. Communications in Statistics: Simulation and Computation, 24(3):551–566, 1995. CODEN CSSCDB. ISSN 0361-0918.

Naus:1996:SUM

[NS96] Joseph I. Naus and Ke-Ning Sheng. Screening for unusual matched segments in multiple protein sequences. Communications in Statistics: Simulation and Computation, 25(4):937–952, 1996. CODEN CSSCDB. ISSN 0361-0918.

Newrnan:1999:PPP

[NSS99] R. M. Newrnan, W. B. Smith, and F. M. Speed. Properties of profile parallelism tests in repeated measures designs. Communications in Statistics: Simulation and Computation, 28(4): 1073–1098, 1999. CODEN CSSCDB. ISSN 0361-0918.

Oksoy:1994:QTC

[ÖA94] Dolun Öksoy and Leo A. Aroian. The quotient of two correlated normal variables with applications. Communications in Statistics: Simulation and Computation, 23(1):223–241, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Ochshorn:1993:SAT

[Och93] Peter Ochshorn. Splined alternatives for testing lack of fit in regression. Communications in Statistics: Simulation and Computation, 22(3):709–721, 1993. CODEN CSSCDB. ISSN 0361-0918.

OGorman:1995:EUV

[O’G95] Thomas W. O’Gorman. The effect of unequal variances on the power of several two-sample tests. Communications in Statistics: Simulation and Computation, 24(4):853–867, 1995. CO-DEN CSSCDB. ISSN 0361-0918.

OGorman:1996:ATS

[O’G96] Thomas W. O’Gorman. An adaptive two-sample test based on modified Wilcoxon scores. Communications in Statistics: Simulation and Computation, 25(2):459–479, 1996. CODEN CSS-CDB. ISSN 0361-0918.

OGorman:1997:CAT

[O’G97] Thomas W. O’Gorman. A comparison of an adaptive two-sample test to the t-test and the rank-sum test. Communications in Statistics: Simulation and Computation, 26(4):1393–1411, 1997. CODEN CSSCDB. ISSN 0361-0918.

Ogden:1997:PDW

[Ogd97] R. Todd Ogden. On preconditioning the data for the wavelet transform when the sample size is not a power of two. Communications in Statistics: Simulation and Computation, 26(2): 467–486, 1997. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1990:OSS

[OM90] Robert E. Odeh and Robert W. Mee. One-sided simultaneous tolerance limits for regression. Communications in Statistics: Simulation and Computation, 19(2):663–680, 1990. CODEN CSS-CDB. ISSN 0361-0918.

Ong:1995:CDD

[OM95] S. H. Ong and Subarau Muthaloo. A class of discrete distributions suited to fitting very long-tailed data. Communications in Statistics: Simulation and Computation, 24(4):929–945, 1995. CODEN CSSCDB. ISSN 0361-0918.

Ogunyemi:1996:EAP

[ON96] O. T. Ogunyemi and Paul I. Nelson. Exact and adaptive prediction intervals for order statistics. Communications in Statistics: Simulation and Computation, 25(4):1057–1074, 1996. CODEN CSSCDB. ISSN 0361-0918.

Ong:1992:CGB

[Ong92] S. H. Ong. The computer generation of bivariate binomial variables with given marginals and correlation. Communications in Statistics: Simulation and Computation, 21(2):285–299, 1992. CODEN CSSCDB. ISSN 0361-0918.

Ong:1998:NMP

[Ong98] S. H. Ong. A note on the mixed Poisson formulation of the Poisson-inverse Gaussian distribution. Communications in Statistics: Simulation and Computation, 27(1):67–78, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Ozturk:1992:NMA

[ÖR92] Aydin Öztürk and Jorge L. Romeu. A new method for assessing multivariate normality with graphical applications. Communications in Statistics: Simulation and Computation, 21(1):15–34, 1992. CODEN CSSCDB. ISSN 0361-0918.

Obuchowski:1995:HTD

[OR95] Nancy A. Obuchowski, Jr. and Howard E. Rockette, Jr. Hypothesis testing of diagnostic accuracy for multiple readers and multiple tests an ANOVA approach with dependent observations. Communications in Statistics: Simulation and Computation, 24 (2):285–308, 1995. CODEN CSSCDB. ISSN 0361-0918.

Ocana:1990:CGE

[ORR90] Jordi Ocaña and Carmen Ruiz-Rivas. Computer generation and estimation in a one-parameter system of bivariate distributions with specified marginals. Communications in Statistics: Simulation and Computation, 19(1):37–55, 1990. CODEN CSSCDB. ISSN 0361-0918.

Pepe:1994:CNI

[PA94] Margaret Sullivan Pepe and Garnet L. Anderson. A cautionary note on inference for marginal regression models with longitudinal data and general correlated response data. Communications in Statistics: Simulation and Computation, 23(4):939–951, 1994. CODEN CSSCDB. ISSN 0361-0918.

Panton:1992:CDF

[Pan92] Don B. Panton. Cumulative distribution function values for symmetric standardized stable distributions. Communications in Statistics: Simulation and Computation, 21(2):485–492, 1992. CODEN CSSCDB. ISSN 0361-0918.

Parrish:1992:CEV

[Par92a] Rudolph S. Parrish. Computing expected values of normal order statistics. Communications in Statistics: Simulation and Computation, 21(1):57–70, 1992. CODEN CSSCDB. ISSN 0361-0918.

Parrish:1992:CVC

[Par92b] Rudolph S. Parrish. Computing variances and covariances of normal order statistics. Communications in Statistics: Simulation and Computation, 21(1):71–101, 1992. CODEN CSSCDB. ISSN 0361-0918.

Pardo:1999:EPM

[Par99] M. C. Pardo. Estimation of parameters for a mixture of normal distributions on the basis of the Cressie and Read divergence. Communications in Statistics: Simulation and Computation, 28 (1):115–130, 1999. CODEN CSSCDB. ISSN 0361-0918.

Park:1994:CIR

[PB94] Dong Joon Park and Richard K. Burdick. Confidence intervals on the regression coefficient in a simple linear regression model with a balanced one-fold nested error structure. Communications in Statistics: Simulation and Computation, 23(1):43–58, 1994. CODEN CSSCDB. ISSN 0361-0918.

Park:1995:RMD

[PBB95] Chanseok Park, Ayanendranath Basu, and Srabashi Basu. Robust minimum distance inference based on combined distances. Communications in Statistics: Simulation and Computation, 24 (3):653–673, 1995. CODEN CSSCDB. ISSN 0361-0918.

Pepple:1994:AID

[PC94] Patricia A. Pepple and Sung C. Choi. Analysis of incomplete data under non-random mechanisms: Bayesian inference. Communications in Statistics: Simulation and Computation, 23(3): 743–767, 1994. CODEN CSSCDB. ISSN 0361-0918.

Pearn:1996:BLE

[PC96] W. L. Pearn and K. S. Chen. A Bayesian-like estimator of c_{pk}. Communications in Statistics: Simulation and Computation, 25 (2):321–329, 1996. CODEN CSSCDB. ISSN 0361-0918.

Pearn:1997:PPC

[PC97] W. L. Pearn and C. S. Chang. The performance of process capability index c_{s} on skewed distributions. Communications in Statistics: Simulation and Computation, 26(4):1361–1377, 1997. CODEN CSSCDB. ISSN 0361-0918.

Prybutok:1997:CFV

[PCH97] Victor R. Prybutok, Howard R. Clayton, and Martha M. Harvey. Comparison of fixed versus variable sampling interval Shewhart control charts in the presence of positively autocorrelated data. Communications in Statistics: Simulation and Computation, 26 (1):83–106, 1997. CODEN CSSCDB. ISSN 0361-0918.

Penev:1995:SAL

[Pen95] Spiridon Penev. Saddlepoint approximation in the linear structural relationship model. Communications in Statistics: Simulation and Computation, 24(2):349–366, 1995. CODEN CSSCDB. ISSN 0361-0918.

Pereira:1990:TCE

[Per90] Helena Iglésias Pereira. Tests for the chacteristic exponent and the scale parameter of symmetric stable distributions. Communications in Statistics: Simulation and Computation, 19(4):1465– 1475, 1990. CODEN CSSCDB. ISSN 0361-0918.

Park:1997:RPD

[PHB97] Chanseok Park, Ian R. Harris, and Ayanendranath Basu. Robust predictive distributions based on the penalized blended weight Hellinger distance. Communications in Statistics: Simulation and Computation, 26(1):21–33, 1997. CODEN CSSCDB. ISSN 0361-0918.

Phillips:1990:CSE

[Phi90] Robert F. Phillips. On constructing sequences estimating the mixing distribution with applications. Communications in Statistics: Simulation and Computation, 19(2):705–720, 1990. CO-DEN CSSCDB. ISSN 0361-0918.

Phillips:1998:PAE

[Phi98] Robert F. Phillips. Partially adaptive estimation via a normal mixture: some further Monte Carlo evidence. Communications in Statistics: Simulation and Computation, 27(1):107–114, 1998. CODEN CSSCDB. ISSN 0361-0918.

Piepho:1995:DHH

[Pie95] Hans-Peter Piepho. Detecting and handling heteroscedasticity in yield trial data. Communications in Statistics: Simulation and Computation, 24(1):243–274, 1995. CODEN CSSCDB. ISSN 0361-0918.

Park:1999:GME

[PJ99] Sang-Hyun Park and Dac-Heung Jang. A graphical method for evaluating the effect of blocking in response surface designs. Communications in Statistics: Simulation and Computation, 28 (2):369–380, 1999. CODEN CSSCDB. ISSN 0361-0918.

Park:1995:ESM

[PK95] Pil S. Park and Anant M. Kshirsagar. Effect of some modifications in standard discriminant analysis. Communications in Statistics: Simulation and Computation, 24(4):807–816, 1995. CODEN CSSCDB. ISSN 0361-0918.

Parshall:1999:CPT

[PKD99] Cynthia G. Parshall, Jeffrey D. Kromrey, and Ronald Dailey. Comparative performance of three statistical tests of homogeneity for sparse i×j contingency tables. Communications in Statistics: Simulation and Computation, 28(1):275–289, 1999. CO-DEN CSSCDB. ISSN 0361-0918.

Platt:1998:EUM

[Pla98] Robert W. Platt. Estimation using the modified profile likelihood in log odds ratio regression analysis. Communications in Statistics: Simulation and Computation, 27(4):905–919, 1998. CODEN CSSCDB. ISSN 0361-0918.

Pliskin:1990:MSE

[Pli90] Jeffrey L. Pliskin. The mean squared error of the generalized ridge regression estimator and the orientation of β. Communications in Statistics: Simulation and Computation, 19(4):1477– 1484, 1990. CODEN CSSCDB. ISSN 0361-0918.

Pinsky:1997:ETB

[PM97] Paul F. Pinsky, Ph. D. and Laurence S. Magder, Ph. D. Evaluating the tradeoff between bias and variance through use of prior probabilities. Communications in Statistics: Simulation and Computation, 26(2):399–421, 1997. CODEN CSSCDB. ISSN 0361-0918.

Park:1998:VEC

[PP98] Young J. Park and Sastry G. Pantula. Variance estimators in the chu-white test for structural change. Communications in Statistics: Simulation and Computation, 27(4):1019–1029, 1998. CODEN CSSCDB. ISSN 0361-0918.

Park:1994:EDV

[PR94] Changsoon Park and Marion R. Reynolds. Economic design of a variable sample size X̄-chart. Communications in Statistics: Simulation and Computation, 23(2):467–483, 1994. CO-DEN CSSCDB. ISSN 0361-0918.

Prabhu:1996:ATD

[PR96] Sharad S. Prabhu and George C. Runger. Analysis of a twodimensional Markov chain. Communications in Statistics: Simulation and Computation, 25(1):75–79, 1996. CODEN CSSCDB. ISSN 0361-0918.

Park:1999:UWC

[PS99] Taesung Park and Dong-Yun Shin. On the use of working correlation matrices in the Gee approach for longitudinal data. Communications in Statistics: Simulation and Computation, 28(4): 1011–1029, 1999. CODEN CSSCDB. ISSN 0361-0918.

Prada-Sanchez:1997:SSI

[PSCY97] J. M. Prada-Sánchez and T. Cotos-Yáñez. A simulation study of iterated and non-iterated bootstrap methods for bias reduction and confidence interval estimation. Communications in Statistics: Simulation and Computation, 26(3):927–946, 1997. CO-DEN CSSCDB. ISSN 0361-0918.

Perez-Salvador:1995:EPM

[PSMR95] Blanca Rosa Pérez-Salvador and Ignacio Méndez-Ramírez. Estimating the population mean using supplementary dephased information. Communications in Statistics: Simulation and Computation, 24(3):733–743, 1995. CODEN CSSCDB. ISSN 0361-0918.

Padgett:1992:RSC

[PTP92] C. S. Padgett, L. A. Thombs, and W. J. Padgett. On the αrisks for Shewhart control charts. Communications in Statistics: Simulation and Computation, 21(4):1125–1147, 1992. CODEN CSSCDB. ISSN 0361-0918.

Purohit:1994:TMR

[Pur94] Sudha G. Purohit. Testing for the minimal repair model versus additional damage at failures. Communications in Statistics: Simulation and Computation, 23(1):89–107, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Pardo:1997:THT

[PVEP97] M. C. Pardo, M. L. Vicente, M. D. Esteban, and J. A. Pardo. Testing hypotheses in truncated samples by means of divergence statistics. Communications in Statistics: Simulation and Computation, 26(2):707–732, 1997. CODEN CSSCDB. ISSN 0361-0918.

Park:1992:GMM

[PW92] Taesung Park and Robert F. Woolson. Generalized multivariate models for longitudinal data. Communications in Statistics: Simulation and Computation, 21(4):925–946, 1992. CO-DEN CSSCDB. ISSN 0361-0918.

Racine:1993:ECV

[Rac93] Jeff Racine. An efficient cross-validation algorithm for window width selection for nonparametric kernel regression. Communications in Statistics: Simulation and Computation, 22(4):1107– 1114, 1993. CODEN CSSCDB. ISSN 0361-0918.

Rahman:1999:EBC

[Rah99] Mezbahur Rahman. Estimating the Box–Cox transformation via Shapiro–Wilk W statistic. Communications in Statistics: Simulation and Computation, 28(1):223–241, 1999. CODEN CSS-CDB. ISSN 0361-0918.

Raju:1992:EOF

[Raj92] C. Raju. On equivalence of OC functions of certain conditional sampling plans. Communications in Statistics: Simulation and Computation, 21(4):961–969, 1992. CODEN CSSCDB. ISSN 0361-0918.

Raubertas:1992:EPP

[Rau92] Richard F. Raubertas. The envelope probability plot as a goodness-of-fit test. Communications in Statistics: Simulation and Computation, 21(1):189–202, 1992. CODEN CSSCDB. ISSN 0361-0918.

Ravishanker:1994:RCM

[Rav94] Nalini Ravishanker. Relative curvature measures of nonlinearity for time series models. Communications in Statistics: Simulation and Computation, 23(2):415–430, 1994. CODEN CSSCDB. ISSN 0361-0918.

Rigdon:1990:EIF

[RB90] Steven E. Rigdon and Asit P. Basu. Estimating the intensity function of a power law process at the current time: time truncated case. Communications in Statistics: Simulation and Computation, 19(3):1079–1104, 1990. CODEN CSSCDB. ISSN 0361-0918.

Rayner:1991:OTT

[RBL91] J. C. W. Rayner, D. J. Best, and G. F. Liddell. Optimal testing in a three way ANOVA model. Communications in Statistics: Simulation and Computation, 20(2–3):411–425, 1991. CODEN CSSCDB. ISSN 0361-0918.

Romera:1995:PIR

[RC95] Rosario Romera and Tomas Cipra. On practical implementation of robust Kalman filtering. Communications in Statistics: Simulation and Computation, 24(2):461–488, 1995. CODEN CSS-CDB. ISSN 0361-0918.

Ravishanker:1996:SEC

[RDW96] Nalini Ravishanker, Dipak K. Dey, and Lilian S.-Y. Wu. Shrinkage estimation of contemporaneous outliers in concurrent time serie. Communications in Statistics: Simulation and Computation, 25(3):643–656, 1996. CODEN CSSCDB. ISSN 0361-0918.

Rencher:1992:BAC

[Ren92] Alvin C. Rencher. BIAS in apparent classification rates in stepwise discriminant analysis. Communications in Statistics: Simulation and Computation, 21(2):373–389, 1992. CODEN CSS-CDB. ISSN 0361-0918.

Ramsey:1996:AOS

[RF96] Philip J. Ramsey and Robert V. Foutz. Asymptotically optimal spectral estimation with applications to small samples. Communications in Statistics: Simulation and Computation, 25(1): 61–73, 1996. CODEN CSSCDB. ISSN 0361-0918.

Rutter:1991:BER

[RFL91] Carolyn Rutter, Virgina Flack, and Peter Lachenbruch. Bias in error rate estimates in discriminant analysis when stepwise variable selection is employed. Communications in Statistics: Simulation and Computation, 20(1):1–22, 1991. CODEN CSSCDB. ISSN 0361-0918.

Rahman:1996:EPE

[RG96a] Mezbahur Rahman and D. V. Gokhale. On estimation of parameters of the exponential power family of distributions. Communications in Statistics: Simulation and Computation, 25(2): 291–299, 1996. CODEN CSSCDB. ISSN 0361-0918.

Rothan:1996:TSD

[RG96b] Sister Adele Marie Rothan and Rudy A. Gideon. A two-sample distribution-free scale test of the Smirnov type. Communications in Statistics: Simulation and Computation, 25(3):785–800, 1996. CODEN CSSCDB. ISSN 0361-0918.

Rahman:1997:NCP

[RGU97] Mezbahur Rahman, D. V. Gokhale, and Aman Ullah. A note on combining parametric and non-parametric regression. Communications in Statistics: Simulation and Computation, 26(2): 519–529, 1997. CODEN CSSCDB. ISSN 0361-0918.

Ramirez:1991:MTI

[RJ91] D. E. Ramirez and D. R. Jensen. Misspecified T^{2} tests. II. Series expansions. Communications in Statistics: Simulation and Computation, 20(1):97–108, 1991. CODEN CSSCDB. ISSN 0361-0918.

Rao:1996:BFP

[RK96] Poduri S. R. S. Rao and Myron J. Katzoff. Bootstrap for finite populations. Communications in Statistics: Simulation and Computation, 25(4):979–994, 1996. CODEN CSSCDB. ISSN 0361-0918.

Rahman:1999:IMS

[RK99] M. S. Rahman and Maxwell L. King. Improved model selection criterion. Communications in Statistics: Simulation and Computation, 28(1):51–71, 1999. CODEN CSSCDB. ISSN 0361-0918.

Rosaiah:1991:OCL

[RKN91a] K. Rosaiah, R. R. L. Kantam, and V. L. Narasimham. Optimum class limits for ML estimation in two-parameter gamma distribution from a grouped data. Communications in Statistics: Simulation and Computation, 20(4):1173–1189, 1991. CODEN CSSCDB. ISSN 0361-0918.

Rosaiah:1991:OLU

[RKN91b] K. Rosaiah, R. R. L. Kantam, and V. L. Narasimham. Optimum linear unbiased estimation of scale parameter by absolute values of order statistics in symmetric distributions. Communications in Statistics: Simulation and Computation, 20(4):1159–1172, 1991. CODEN CSSCDB. ISSN 0361-0918.

Ressler:1990:VRQ

[RL90] Richard L. Ressler and Peter A. W. Lewis. Variance reduction for quantile estimates in simulations via nonlinear controls. Communications in Statistics: Simulation and Computation, 19(3): 1045–1077, 1990. CODEN CSSCDB. ISSN 0361-0918.

Russell:1995:ETP

[RL95] Gerald S. Russell and Daniel J. Levitin. An expanded table of probability values for Rao’s spacing test. Communications in Statistics: Simulation and Computation, 24(4):879–888, 1995. CODEN CSSCDB. ISSN 0361-0918.

Ren:1996:CMT

[RL96] Jian-Jian Ren and Glenn Ledder. On cramér-von Mises tests of goodness of fit with censored data. Communications in Statistics: Simulation and Computation, 25(1):169–187, 1996. CO-DEN CSSCDB. ISSN 0361-0918.

Ran:1998:MLE

[RLR98] Shiaohong Ran, David Lansky, and Arthur Roth. Maximum likelihood estimation from replicate limiting dilution assays from a binomial Markov chain model. Communications in Statistics: Simulation and Computation, 27(4):889–903, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Raju:1996:TSC

[RM96] C. Raju and M. N. Narasimha Murthy. Two-stage chain sampling plans ChSP-(0,2) and ChSP-(1,2) — Part 1. Communications in Statistics: Simulation and Computation, 25(2):557–572, 1996. CODEN CSSCDB. ISSN 0361-0918.

Rodrigues:1999:DCE

[Rod99] Josemar Rodrigues. Diagnostics of convergence of an estimate of the marginal density via calibrated divergence measures. Communications in Statistics: Simulation and Computation, 28(4): 895–907, 1999. CODEN CSSCDB. ISSN 0361-0918.

Rose:1993:SER

[Ros93] Elizabeth L. Rose. Some effects of rounding errors on ARMA(1, 1) models. Communications in Statistics: Simulation and Computation, 22(1):155–174, 1993. CODEN CSSCDB. ISSN 0361-0918.

Radhakrishnan:1992:MAV

[RR92] R. Radhakrishnan and Don R. Robinson. Multivariate analysis of variance with additional data from an unknown subset of the populations. Communications in Statistics: Simulation and Computation, 21(3):659–669, 1992. CODEN CSSCDB. ISSN 0361-0918.

Radhakrishnan:1994:TER

[RR94] R. Radhakrishnan and Don R. Robinson. Testing the equality of regression matrices when some additional data sets are available. Communications in Statistics: Simulation and Computation, 23 (1):175–192, 1994. CODEN CSSCDB. ISSN 0361-0918.

Rueda:1999:TFD

[RR99] R. Rueda and F. O’ Reilly. Tests of fit for discrete distributions based on the probability generating function. Communications in Statistics: Simulation and Computation, 28(1):259–274, 1999. CODEN CSSCDB. ISSN 0361-0918.

Rao:1991:OLU

[RRN91] A. Vasudeva Rao, A. V. Dattatreya Rao, and V. L. Narasimham. Optimum linear unbiased estimation of the scale parameter by absolute values of order statistics in the double exponential and the double Weibull distributions. Communications in Statistics: Simulation and Computation, 20(4):1139–1158, 1991. CODEN CSSCDB. ISSN 0361-0918.

Rao:1994:AOG

[RRN94] A. Vasudeva Rao, A. V. Dattatreya Rao, and V. L. Narasimham. Asymptotically optimal grouping for maximum likelihood estimation of Weibull parameters. Communications in Statistics: Simulation and Computation, 23(4):1077–1096, 1994. CODEN CSSCDB. ISSN 0361-0918.

Rencher:1990:ACI

[RS90] Alvin C. Rencher and Del T. Scott. Assessing the contribution of individual variables following rejection of a multivariate hypothesis. Communications in Statistics: Simulation and Computation, 19(2):535–553, 1990. CODEN CSSCDB. ISSN 0361-0918.

Reitmeir:1996:OSM

[RW96] Peter Reitmeir and Gernot Wassmer. One-sided multiple endpoint testing in two-sample comparisons. Communications in Statistics: Simulation and Computation, 25(1):99–117, 1996. CODEN CSSCDB. ISSN 0361-0918.

Randall:1997:WTM

[RWB97] Robert L. Randall, J. Arthur Woodward, and Douglas G. Bonett. A Wald test for the multivariate analysis of variance: small sample critical values. Communications in Statistics: Simulation and Computation, 26(4):1275–1299, 1997. CODEN CSSCDB. ISSN 0361-0918.

Rychlik:1992:CBL

[Ryc92] Igor Rychlik. Confidence bands for linear regressions. Communications in Statistics: Simulation and Computation, 21(2): 333–352, 1992. CODEN CSSCDB. ISSN 0361-0918.

Soundararajan:1990:GPS

[SA90] V. Soundararajan and S. Devaraj Arumainayagam. A generalized procedure for selection of attribute double sampling plan. Communications in Statistics: Simulation and Computation, 19 (3):1015–1034, 1990. CODEN CSSCDB. ISSN 0361-0918.

Senyo:1991:PRP

[SA91] B. Senyo and S. K. Adjibolosoo. The power and robustness properties of tests for heteroskedasticity when the regressors are trended. Communications in Statistics: Simulation and Computation, 20(4):927–953, 1991. CODEN CSSCDB. ISSN 0361-0918.

Superville:1994:EFB

[SA94] Claude R. Superville and Benjamin M. Adams. An evaluation of forecast-based quality control schemes. Communications in Statistics: Simulation and Computation, 23(3):645–661, 1994. CODEN CSSCDB. ISSN 0361-0918.

Shaarawy:1995:BCM

[SA95] Samir M. Shaarawy and Mohammad A. Almahmeed. Bayesian classification with multivariate autoregressive sources that might have different orders. Communications in Statistics: Simulation and Computation, 24(3):567–581, 1995. CODEN CSSCDB. ISSN 0361-0918.

Song:1999:PNM

[SA99] Kyunghee K. Song and Weissfeld Lisa A. Performance of nonparametric maximum likelihood estimator in a. Communications in Statistics: Simulation and Computation, 28(3):637–655, 1999. CODEN CSSCDB. ISSN 0361-0918.

Saccucci:1992:EWM

[SAL92] Michael S. Saccucci, Raid W. Amin, and James M. Lucas. Exponentially weighted moving average control schemes with variable sampling intervals. Communications in Statistics: Simulation and Computation, 21(3):627–657, 1992. CODEN CSSCDB. ISSN 0361-0918.

Salau:1995:NIG

[Sal95] M. O. Salau. On the numerical implementation of the generalized least squares procedure for ARMA estimation. Communications in Statistics: Simulation and Computation, 24(1):111–130, 1995. CODEN CSSCDB. ISSN 0361-0918.

Salau:1997:NCA

[Sal97] M. O. Salau. Numerical computation of asymptotic covariance matrix of the Gaussian estimators for vector ARRLA models. Communications in Statistics: Simulation and Computation, 26 (1):173–192, 1997. CODEN CSSCDB. ISSN 0361-0918.

Salau:1999:NET

[Sal99] M. O. Salau. On the numerical evaluation of the theoretical variance-covariance matrix of least squares estimators for echelon-form Varma models. Communications in Statistics: Simulation and Computation, 28(3):743–766, 1999. CODEN CSS-CDB. ISSN 0361-0918.

Samawi:1999:MEM

[Sam99] Hani M. Samawi. More efficient Monte Carlo methods obtained by using ranked set simulated samples. Communications in Statistics: Simulation and Computation, 28(3):699–713, 1999. CODEN CSSCDB. ISSN 0361-0918.

Sankoh:1995:SSP

[San95] Abdul J. Sankoh. Small-sample properties of finite population ratios. Communications in Statistics: Simulation and Computation, 24(4):897–909, 1995. CODEN CSSCDB. ISSN 0361-0918.

Shamma:1991:DEW

[SAS91] Shawky E. Shamma, Raid W. Amin, and Amal K. Shamma. A double exponentially weighted moving average control procedure with variable sampling intervals. Communications in Statistics: Simulation and Computation, 20(2–3):511–528, 1991. CODEN CSSCDB. ISSN 0361-0918.

Seyedsadr:1992:SMM

[SC92] Mahmoud Seyedsadr and Paul L. Cornelius. Shifted multiplicative models for nonadditive two-way tables. Communications in Statistics: Simulation and Computation, 21(3):807–832, 1992. CODEN CSSCDB. ISSN 0361-0918.

Schwertman:1993:EFO

[SC93a] Neil C. Schwertman and Nancy J. Carter. The effect of first order auto correlation on data analysis as measured by the Geisser– Greenhouse adjustment. Communications in Statistics: Simulation and Computation, 22(3):791–796, 1993. CODEN CSSCDB. ISSN 0361-0918.

Seyedsadr:1993:HTC

[SC93b] Mahmoud S. Seyedsadr and Paul L. Cornelius. Hypothesis testing for components of the shifted multiplicative model for a nonadditive two-way table. Communications in Statistics: Simulation and Computation, 22(4):1065–1078, 1993. CODEN CSS-CDB. ISSN 0361-0918.

Schilling:1992:SDW

[Sch92] Mark F. Schilling. Spatial designs when the observations are correlated. Communications in Statistics: Simulation and Computation, 21(1):243–267, 1992. CODEN CSSCDB. ISSN 0361-0918.

Schork:1993:CMC

[Sch93] Nicholas Schork. Combining Monte Carlo and Cox tests of nonnested hypotheses. Communications in Statistics: Simulation and Computation, 22(4):939–954, 1993. CODEN CSSCDB. ISSN 0361-0918.

Schaefer:1994:UDT

[Sch94] Robert L. Schaefer. Using default tests in repeated measures: how bad can it get? Communications in Statistics: Simulation and Computation, 23(1):109–127, 1994. CODEN CSSCDB. ISSN 0361-0918.

Schuster:1996:CDL

[Sch96] Eugene F. Schuster. The conditional distribution of the longest run in a sample from a multiletter alphabet. Communications in Statistics: Simulation and Computation, 25(1):215–224, 1996. CODEN CSSCDB. ISSN 0361-0918.

Schaalje:1996:RHV

[SD96] G. Bruce Schaalje and Darrin J. Despain. Robustness of homogeneity of variance tests for randomized complete block data.

Communications in Statistics: Simulation and Computation, 25 (4):961–977, 1996. CODEN CSSCDB. ISSN 0361-0918.

Sa:1993:TCS

[SE93a] Ping Sa and Don Edwards. Tables of the Casella and Strawderman (1980) critical points for use in multiple comparisons with a control in response surface methodology. Communications in Statistics: Simulation and Computation, 22(3):765–777, 1993. CODEN CSSCDB. ISSN 0361-0918.

Shiue:1993:IDF

[SE93b] Wei Kei Shiue and Max Engelhardt. Inferences on the difference of failure intensities from two exponential distributions. Communications in Statistics: Simulation and Computation, 22(1): 175–189, 1993. CODEN CSSCDB. ISSN 0361-0918.

Seaman:1997:TPM

[Sea97] Michael A. Seaman. Tables for pairwise multiple comparisons using Shaffer’s modified sequentially-rejective procedure. Communications in Statistics: Simulation and Computation, 26(2): 687–705, 1997. CODEN CSSCDB. ISSN 0361-0918.

Salter:1993:ATI

[SF93] K. C. Salter and R. F. Fawcett. The art test of interaction: a robust and powerful rank test of interaction in factorial models. Communications in Statistics: Simulation and Computation, 22 (1):137–153, 1993. CODEN CSSCDB. ISSN 0361-0918.

Swanepoel:1994:BCI

[SF94] C. J. Swanepoel and C. C. Frangos. Bootstrap confidence intervals for the slope parameter of a logistic model. Communications in Statistics: Simulation and Computation, 23(4):1115– 1126, 1994. CODEN CSSCDB. ISSN 0361-0918.

Shanmugam:1992:DSC

[SG92] M. Shanmugam and A. V. Gajjar. On the distribution of sample correlation coefficient and its transformations in samples from Morgenstern bivariate exponential distribution. Communications in Statistics: Simulation and Computation, 21(4):1021– 1041, 1992. CODEN CSSCDB. ISSN 0361-0918.

Selwyn:1993:ANE

[SG93a] Murray R. Selwyn and Peter P. Gaccione. Assessing normality and equivariance in small multigroup samples. Communications in Statistics: Simulation and Computation, 22(1):61–79, 1993. CODEN CSSCDB. ISSN 0361-0918.

Shah:1993:MPS

[SG93b] Aarti Shah and D. V. Gokhale. On maximum product of spagings (mps) estimation for Burr XII distributions. Communications in Statistics: Simulation and Computation, 22(3):615–641, 1993. CODEN CSSCDB. ISSN 0361-0918.

Sodahl:1999:SEE

[SG99] X. Sodahl and F. Godtliebsen. Semiparametric estimation of extremes. Communications in Statistics: Simulation and Computation, 28(3):569–595, 1999. CODEN CSSCDB. ISSN 0361-0918.

Stapanian:1991:PTT

[SGF^{+}91] Martin A. Stapanian, Forest C. Garner, Kirk E. Fitzgerald, George T. Flatman, and Evan J. Englund. Properties of two tests for outliers in multivariate data. Communications in Statistics: Simulation and Computation, 20(2–3):667–687, 1991. CODEN CSSCDB. ISSN 0361-0918.

Serlin:1993:ESE

[SH93] Ronald C. Serlin and Michael R. Harwell. An empirical study of eight tests of partial correlation coefficients. Communications in Statistics: Simulation and Computation, 22(2):545–567, 1993. CODEN CSSCDB. ISSN 0361-0918.

Schuster:1998:MSP

[SH98] Eugene F. Schuster and Jie He. On mode(s) and shape of the pdfs of the number of runs and of the maximum of runs by type. Communications in Statistics: Simulation and Computation, 27 (1):115–126, 1998. CODEN CSSCDB. ISSN 0361-0918.

Sharma:1991:NJT

[Sha91] Subhash C. Sharma. A new jackknife test for homogeneity of variances. Communications in Statistics: Simulation and Computation, 20(2–3):479–495, 1991. CODEN CSSCDB. ISSN 0361-

Shirke:1999:NPE

[Shi99] D. T. Shirke. A note on pretest estimation for some discrete distributions. Communications in Statistics: Simulation and Computation, 28(2):431–439, 1999. CODEN CSSCDB. ISSN 0361-0918.

Shore:1997:ETC

[Sho97] Haim Shore. Enhancement for two commonly-used approximations for the inverse cumulative function of the normal distribution. Communications in Statistics: Simulation and Computation, 26(3):1041–1047, 1997. CODEN CSSCDB. ISSN 0361-0918.

Shore:1998:AUD

[Sho98] Haim Shore. Approximating an unknown distribution when distribution information is extremely limited. Communications in Statistics: Simulation and Computation, 27(2):501–523, 1998. CODEN CSSCDB. ISSN 0361-0918.

Shoemaker:1999:ITD

[Sho99] Lewis H. Shoemaker. Interquartile tests for dispersion in skewed distributions. Communications in Statistics: Simulation and Computation, 28(1):189–205, 1999. CODEN CSSCDB. ISSN 0361-0918.

Silva:1997:GPR

[Sil97] J. M. C. Santos Silva. Generalized Poisson regression for positive count data. Communications in Statistics: Simulation and Computation, 26(3):1089–1102, 1997. CODEN CSSCDB. ISSN 0361-0918.

Simonoff:1995:APH

[Sim95] Jeffrey S. Simonoff. The anchor position of histograms and frequency polygons: quantitative and qualitative smoothing. Communications in Statistics: Simulation and Computation, 24(3): 691–710, 1995. CODEN CSSCDB. ISSN 0361-0918.

S:1991:LLI

[SJ91] Wang H. S. and Karson M. J. LIME: the least integrated meansquared-error estimator for the maximum point of the singlefactor quadratic response function. Communications in Statistics: Simulation and Computation, 20(1):41–71, 1991. CODEN CSSCDB. ISSN 0361-0918.

Sathe:1991:CCE

[SJNM91] Y. S. Sathe, S. Joshi, S. Nabar, and T. M. Murdeshwar. Computation of circular error probabilities. Communications in Statistics: Simulation and Computation, 20(1):33–40, 1991. CODEN CSSCDB. ISSN 0361-0918.

Sudmant:1990:IPH

[SK90] Walter Sudmant and Peter Kennedy. On inference in the presence of heteroskedasticity without replicated observations. Communications in Statistics: Simulation and Computation, 19(2): 491–504, 1990. CODEN CSSCDB. ISSN 0361-0918.

Spurr:1991:CVM

[SK91] Barrie D. Spurr and Majdi A. Koutbeiy. A comparison of various methods for estimating the parameters in mixtures of von Mises distributions. Communications in Statistics: Simulation and Computation, 20(2–3):725–741, 1991. CODEN CSSCDB. ISSN 0361-0918.

Sohn:1992:ETP

[SK92] Joong Kweon Sohn and Sang Gil Kang. On the estimation of the true probability of a correct selection. Communications in Statistics: Simulation and Computation, 21(2):445–462, 1992. CO-DEN CSSCDB. ISSN 0361-0918.

Spurrier:1993:EPW

[SK93] John D. Spurrier and Sandra J. Kohta. Estimation of proportions of white and non-white voters voting for a candidate with applications to possible violations of Section 2 of the Voting Rights Act. Communications in Statistics: Simulation and Computation, 22(3):845–862, 1993. CODEN CSSCDB. ISSN 0361-0918.

Stiger:1998:ARO

[SKBK98] Thomas R. Stiger, Andrzej S. Kosinski, Huiman X. Barnhart, and David G. Kleinbaum. ANOVA for repeated ordinal data with small sample size? A comparison of ANOVA, MANOVA, WLS and Gee methods by simulation. Communications in Statistics: Simulation and Computation, 27(2):357–375, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Siu-Keung:1996:SER

[SKG96] Tse Siu-Keung and Tso Geoffrey. Shrinkage estimation of reliability for exponentially distributed lifetimes. Communications in Statistics: Simulation and Computation, 25(2):415–430, 1996. CODEN CSSCDB. ISSN 0361-0918.

Silvapulle:1997:RAT

[SL97] Paramsothy Silvapulle and John Lee. Robustness of the arch tests in the presence of serial correlation. Communications in Statistics: Simulation and Computation, 26(2):649–669, 1997. CODEN CSSCDB. ISSN 0361-0918.

Sherman:1997:CBB

[SlC97] Michael Sherman and Saskia le Cessie. A comparison between bootstrap methods and generalized estimating equations for correlated outcomes in generalized linear models. Communications in Statistics: Simulation and Computation, 26(3):901–925, 1997. CODEN CSSCDB. ISSN 0361-0918.

Shoukri:1991:SAS

[SM91] M. M. Shoukri and I. U. H. Mian. Some aspects of statistical inference on the Lagrange (generalized) Poisson distribution. Communications in Statistics: Simulation and Computation, 20 (4):1115–1137, 1991. CODEN CSSCDB. ISSN 0361-0918.

Selicato:1998:APU

[SM98a] Grace R. Selicato and Keith E. Muller. Approximating power of the unconditional test. Communications in Statistics: Simulation and Computation, 27(2):553–564, 1998. CODEN CSSCDB. ISSN 0361-0918.

Simpson:1998:DEA

[SM98b] James R. Simpson and Douglas C. Montgomery. The development and evaluation of alternative generalized m-estimation techniques. Communications in Statistics: Simulation and Computation, 27(4):999–1018, 1998. CODEN CSSCDB. ISSN 0361-0918.

Simpson:1998:PBA

[SM98c] James R. Simpson and Douglas C. Montgomery. A performancebased assessment of robust regression methods. Communications in Statistics: Simulation and Computation, 27(4):1031– 1049, 1998. CODEN CSSCDB. ISSN 0361-0918.

Smith:1990:UAD

[Smi90] Jeffrey W. Smith. The use of auxiliary data in robust univariate mean estimation. Communications in Statistics: Simulation and Computation, 19(3):787–807, 1990. CODEN CSSCDB. ISSN 0361-0918.

Smith:1997:E

[Smi97a] William B. Smith. Editorial. Communications in Statistics: Simulation and Computation, 26(1):ix–x, 1997. CODEN CSSCDB. ISSN 0361-0918.

Smith:1997:PEC

[Smi97b] William B. Smith. Publication is as easy as C–C–C. Communications in Statistics: Simulation and Computation, 26(1):xi–xvi, 1997. CODEN CSSCDB. ISSN 0361-0918.

Smith:1998:E

[Smi98] William B. Smith. Editorial. Communications in Statistics: Simulation and Computation, 27(1):v–vi, 1998. CODEN CSSCDB. ISSN 0361-0918.

Smith:1999:CEC

[Smi99a] William B. Smith. Comments from the editor-in-chief. Communications in Statistics: Simulation and Computation, 28(1):vii, 1999. CODEN CSSCDB. ISSN 0361-0918.

Smith:1999:EB

[Smi99b] William B. Smith. Editorial Board. Communications in Statistics: Simulation and Computation, 28(1):v–vi, 1999. CODEN CSSCDB. ISSN 0361-0918.

Sarkar:1995:ADR

[SMR95] Ila C. Sarkar, Govind S. Mudholkar, and Richard F. Raubertas. An approximation to the distribution of the resultant from a Fisher distribution. Communications in Statistics: Simulation and Computation, 24(1):227–241, 1995. CODEN CSSCDB. ISSN 0361-0918.

Sugiura:1994:TCP

[SO94] N. Sugiura and R. T. Ogden. Testing change-points with linear trend. Communications in Statistics: Simulation and Computation, 23(2):287–322, 1994. CODEN CSSCDB. ISSN 0361-0918.

Scazzero:1995:IEO

[SO95] Joseph A. Scazzero and J. K. Ord. Interval estimates for the optimum point of a quadratic logistic curve — a comparison of different estimation methods via simulation. Communications in Statistics: Simulation and Computation, 24(1):207–226, 1995. CODEN CSSCDB. ISSN 0361-0918.

Soffritti:1999:HCV

[Sof99] Gabriele Soffritti. Hierarchical clustering of variables: a comparison among strategies of analysis. Communications in Statistics: Simulation and Computation, 28(4):977–999, 1999. CO-DEN CSSCDB. ISSN 0361-0918.

Salama:1990:NSF

[SQ90] Ibrahim A. Salama and Dana Quade. A note on Spearman’s footrule. Communications in Statistics: Simulation and Computation, 19(2):591–601, 1990. CODEN CSSCDB. ISSN 0361-0918.

Suresh:1997:ASP

[SR97a] R. P. Suresh and T. V. Ramanathan. Acceptance sampling plans by variables for a class of symmetric distributions. Communications in Statistics: Simulation and Computation, 26(4):1379– 1391, 1997. CODEN CSSCDB. ISSN 0361-0918.

Syversveen:1997:MAF

[SR97b] Anne Randi Syversveen and Håvard Rue. A method for approximate fully Bayesian analysis of parameters. Communications in Statistics: Simulation and Computation, 26(3):1147–1162, 1997. CODEN CSSCDB. ISSN 0361-0918.

Shao:1999:ACI

[SRHW99] Yuehjen E. Shao, George C. Runger, Jorge Haddock, and W. A. Wallace. Adaptive controllers to integrate SPC and EPC. Communications in Statistics: Simulation and Computation, 28(1): 13–36, 1999. CODEN CSSCDB. ISSN 0361-0918.

Sidak:1992:SSS

[ŠŠ92] Zbyněk Šidák and Ondrej Šidák. Strategies for a sequential selection of the better of two trinomial populations. Communications in Statistics: Simulation and Computation, 21(4):1171– 1180, 1992. CODEN CSSCDB. ISSN 0361-0918.

Shih:1995:ESM

[SS95] Neng Hui Shih and Wheyming Tina Song. Estimating simulation metamodels using correlated control variates. Communications in Statistics: Simulation and Computation, 24(1):79–90, 1995. CODEN CSSCDB. ISSN 0361-0918.

See:1996:GTT

[SS96] Kyoungah See and Eric P. Smith. A graphical tool for threefactor experiments with one observation per cell. Communications in Statistics: Simulation and Computation, 25(3):709–732, 1996. CODEN CSSCDB. ISSN 0361-0918.

Speed:1998:EBB

[SS98] F. Michael Speed and Clifford Spiegelman. Evaluating black boxes: an ad-hoc method for assessing nonparametric and nonlinear curve-fitting estimators. Communications in Statistics: Simulation and Computation, 27(3):699–710, 1998. CODEN CSS-CDB. ISSN 0361-0918.

Schechtman:1998:IEI

[SSM98] E. Schechtman, N. Spiegelman, and C. Moran. Interval estimate for the x-intercept of a straight line: a nonlinear approach. Communications in Statistics: Simulation and Computation, 27 (4):1093–1115, 1998. CODEN CSSCDB. ISSN 0361-0918.

Singh:1995:APL

[SSW95] Bahadur Singh, Michael J. Schell, and F. T. Wright. Approximations to the powers of the likelihood ratio tests: the loop ordering and slippage alternatives. Communications in Statistics: Simulation and Computation, 24(1):91–109, 1995. CODEN CSSCDB. ISSN 0361-0918.

Shoemaker:1993:RMA

[ST93] Anne C.. Shoemaker and Kwok-Leung Tsui. Response model analysis for robust design experiments. Communications in Statistics: Simulation and Computation, 22(4):1037–1064, 1993. CODEN CSSCDB. ISSN 0361-0918.

Steel:1993:GAS

[Ste93] S. J. Steel. Group adaptive Stein estimation of normal means. Communications in Statistics: Simulation and Computation, 22 (3):723–737, 1993. CODEN CSSCDB. ISSN 0361-0918.

Stromberg:1997:SSC

[Str97] Arnold J. Stromberg. Some software for computing robust linear or nonlinear regression estimators. Communications in Statistics: Simulation and Computation, 26(3):947–959, 1997. CO-DEN CSSCDB. ISSN 0361-0918.

Seki:1993:BNP

[STY93] Tetsurou Seki, Tadahiro Takashina, and Shinichiro Yokoyama. Bootstrap for the normal parameters. Communications in Statistics: Simulation and Computation, 22(1):191–203, 1993. CO-DEN CSSCDB. ISSN 0361-0918.

Sugiyama:1992:PLR

[SU92] T. Sugiyama and K. Ushizawa. Power of largest root on canonical correlation. Communications in Statistics: Simulation and Computation, 21(4):947–960, 1992. CODEN CSSCDB. ISSN 0361-0918.

Sugiura:1993:AMO

[Sug93] Nariaki Sugiura. Assessing multinormality and ordered covariance matrices in a classical data. Communications in Statistics: Simulation and Computation, 22(3):643–654, 1993. CO-DEN CSSCDB. ISSN 0361-0918.

Sugiura:1994:ECS

[Sug94] Nariaki Sugiura. Errata correction and supplement to assessing multinormality and ordered covariance matrices in a classical data. Communications in Statistics: Simulation and Computation, 23(1):281–284, 1994. CODEN CSSCDB. ISSN 0361-0918.

Sungur:1990:DIP

[Sun90] Engin A. Sungur. Dependence information in parameterized copulas. Communications in Statistics: Simulation and Computation, 19(4):1339–1360, 1990. CODEN CSSCDB. ISSN 0361-0918.

Sutradhar:1990:NPT

[Sut90] Brajendra C. Sutradhar. A note on the power of the F -test for testing linear hypothesis with elliptic t errors. Communications in Statistics: Simulation and Computation, 19(3):863–867, 1990. CODEN CSSCDB. ISSN 0361-0918.

Scariano:1990:CCE

[SW90] Stephen M. Scariano and Terry A. Watkins. Comparisons of changepoint estimators. Communications in Statistics: Simulation and Computation, 19(2):619–636, 1990. CODEN CSSCDB. ISSN 0361-0918.

Sankey:1998:SED

[SW98a] Steadman S. Sankey and Lisa A. Weissfeld. A study of the effect of dichotomizing ordinal data upon modeling. Communications in Statistics: Simulation and Computation, 27(4):871–887, 1998. CODEN CSSCDB. ISSN 0361-0918.

Sullivan:1998:ACC

[SW98b] Joe H. Sullivan and William H. Woodall. Adapting control charts for the preliminary analysis of multivariate observations. Communications in Statistics: Simulation and Computation, 27(4): 953–979, 1998. CODEN CSSCDB. ISSN 0361-0918.

Swartz:1992:GFT

[Swa92] T. Swartz. Goodness-of-fit tests using Kullback–Leibler information. Communications in Statistics: Simulation and Computation, 21(3):711–729, 1992. CODEN CSSCDB. ISSN 0361-0918.

Sankey:1996:AUC

[SWFK96] Steadman S. Sankey, M. Sc., Lisa A. Weissfeld, Ph. D., Michael J. Fine, M. D., M. Sc., and Wishwa Kapoor M. D.. M. P. H. An assessment of the use of the continuity correction for sparse data in meta-analysis. Communications in Statistics: Simulation and Computation, 25(4):1031–1056, 1996. CODEN CSSCDB. ISSN 0361-0918.

Swift:1994:ECC

[Swi94] A. Louise Swift. Experience with a criterion for the classification of Johnson distributions. Communications in Statistics: Simulation and Computation, 23(4):1127–1136, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Schuster:1997:CUD

[SX97] Eugene F. Schuster and Gu Xiangjun. On the conditional and unconditional distributions of the number of runs in a sample from a multisymbol alphabet. Communications in Statistics: Simulation and Computation, 26(2):423–442, 1997. CODEN CSSCDB. ISSN 0361-0918.

Santner:1993:ISS

[SY93] Thomas J. Santner and Shin Yamagami. Invariant small sample confidence intervals for the difference of two success probabilities. Communications in Statistics: Simulation and Computation, 22 (1):33–59, 1993. CODEN CSSCDB. ISSN 0361-0918.

Tsui:1992:ADL

[TA92] Albert K. Tsui and Mukhtar M. Ali. Approximations to the distribution of the least squares estimator in a first order stationary autoregressive model. Communications in Statistics: Simulation and Computation, 21(2):463–484, 1992. CODEN CSSCDB. ISSN 0361-0918.

Taam:1997:QLA

[Taa97] Winson Taam. A quasi-likelihood approach to model Bernoulli data with spatial dependence. Communications in Statistics: Simulation and Computation, 26(2):591–603, 1997. CODEN CSSCDB. ISSN 0361-0918.

Tanizaki:1999:NNF

[Tan99] Hisashi Tanizaki. Nonlinear and nonnormal filter using importance sampling: antithetic Monte Carlo integration. Communications in Statistics: Simulation and Computation, 28(2):463–486, 1999. CODEN CSSCDB. ISSN 0361-0918.

Taylor:1990:PML

[Tay90] Jeremy M. G. Taylor. Properties of maximum likelihood estimates of the ratio of parameters in ordinal response regression models. Communications in Statistics: Simulation and Computation, 19(2):469–480, 1990. CODEN CSSCDB. ISSN 0361-0918.

Tan:1999:OAT

[TB99] Frans E. S. Tan and Martijn P. F. Berger. Optimal allocation of time points for the random effects model. Communications in Statistics: Simulation and Computation, 28(2):517–540, 1999. CODEN CSSCDB. ISSN 0361-0918.

Thiart:1993:SSB

[TDTC93] C. Thiart, T. T. Dunne, C. G. Troskie, and D. O. Chalton. A simulation study of biased estimators against the ordinary least squares estimator. Communications in Statistics: Simulation and Computation, 22(2):569–589, 1993. CODEN CSSCDB. ISSN 0361-0918.

Tan:1993:IPC

[TG93] Ming Tan and Leon J. Gleser. Improved point and confidence interval estimators of mean response in simulation when control variates are used. Communications in Statistics: Simulation and Computation, 22(4):1211–1220, 1993. CODEN CSSCDB. ISSN 0361-0918.

Thabane:1999:ABA

[TH99] L. Thabane and M. Safiul Haq. An alternative Bayesian approach to the multivariate Behrens–Fisher problem. Communications in Statistics: Simulation and Computation, 28(1):243–258, 1999. CODEN CSSCDB. ISSN 0361-0918.

Toffolo:1998:NIC

[THD98] R. W. Toffolo, T. J. Harris, and J. H. Davis. The numerical inversion of the characteristic equation with applications to positive quadratic forms in normal variables. Communications in Statistics: Simulation and Computation, 27(1):215–228, 1998. CO-DEN CSSCDB. ISSN 0361-0918.

Thomas:1995:NCS

[Tho95] G. E. Thomas. A note on chi-squared approximations for tests of rank interaction with ordered categorical data. Communications in Statistics: Simulation and Computation, 24(2):439–459, 1995. CODEN CSSCDB. ISSN 0361-0918.

Tsai-Hung:1997:BRB

[THWL97] Fan Tsai-Hung and Hung Wen-Liang. Balanced resampling for bootstrapping finite state Markov chains. Communications in Statistics: Simulation and Computation, 26(4):1465–1475, 1997. CODEN CSSCDB. ISSN 0361-0918.

Trenkler:1990:MMS

[TI90] G. Trenkler and G. Ihorst. Matrix mean square error comparisons based on a certain covariance structure. Communications in Statistics: Simulation and Computation, 19(3):1035–1043, 1990. CODEN CSSCDB. ISSN 0361-0918.

Tadikamalla:1990:TFF

[TJ90] Pandu R. Tadikamalla and Norman L. Johnson. Tables to facilitate fitting Tadikamalla and Johnson’s L_{B} distributions. Communications in Statistics: Simulation and Computation, 19(4): 1201–1229, 1990. CODEN CSSCDB. ISSN 0361-0918.

Tadikamalla:1994:TFF

[TJ94] Pandu R. Tadikamalla and Norman L. Johnson. Tables to facilitate fitting Tadikamalla and Johnson’s L^{B} distributions II. Both terminals fixed. Communications in Statistics: Simulation and Computation, 23(3):569–581, 1994. CODEN CSSCDB. ISSN 0361-0918.

Tricker:1997:ERS

[TO97] A. R. Tricker and E. Okell. The effect of rounding on sequential and fixed size sample hypothesis tests. Communications in Statistics: Simulation and Computation, 26(4):1413–1429, 1997. CODEN CSSCDB. ISSN 0361-0918.

Tornkvist:1990:LCP

[Tör90] Birgitta Törnkvist. On locating and characterizing parameter variation by the MOSUMSQ test statistic. Communications in Statistics: Simulation and Computation, 19(4):1247–1264, 1990. CODEN CSSCDB. ISSN 0361-0918.

Thiagarajh:1990:TES

[TP90] K. Thiagarajh and S. R. Paul. Testtesting for the equality of scale parameters of k(> 2) exponential populations based on complete and Type II censored samples. Communications in Statistics: Simulation and Computation, 19(3):891–902, 1990. CODEN CSSCDB. ISSN 0361-0918.

Turkheimer:1999:MVA

[TPSS99] F. Turkheimer, K. Pettigrew, L. Sokoloff, and K. Schmidt. A minimum variance adaptive technique for parameter estimation and hypothesis testing. Communications in Statistics: Simulation and Computation, 28(4):931–956, 1999. CODEN CSSCDB. ISSN 0361-0918.

Tan:1997:MDM

[TQK97] Ming Tan, Yingsheng Qu, and Michael H. Kutner. Model diagnostics for marginal regression analysis of correlated binary data. Communications in Statistics: Simulation and Computation, 26 (2):539–558, 1997. CODEN CSSCDB. ISSN 0361-0918.

Tsubaki:1995:NWS

[Tsu95] Michiko Tsubaki. New weighted sum of x^{2} test for testing a linear hypothesis against a restricted alternative hypothesis. Communications in Statistics: Simulation and Computation, 24(4): 987–1004, 1995. CODEN CSSCDB. ISSN 0361-0918.

Tiku:1998:TUR

[TW98] Moti L. Tiku and Wing-Keung Wong. Testing for a unit root in an AR(1) model using three and four moment approximations: symmetric distributions. Communications in Statistics: Simulation and Computation, 27(1):185–198, 1998. CODEN CSSCDB. ISSN 0361-0918.

Tarn:1998:SAB

[TYF98] Cannes Y. C. Tarn, Philip L. H. Yu, and Tony W. K. Fung. Sensitivity analysis of blue for the population mean based on a ranked set sample. Communications in Statistics: Simulation and Computation, 27(4):1075–1091, 1998. CODEN CSSCDB. ISSN 0361-0918.

Unnithan:1995:MVS

[UN95] V. K. G. Unnithan and N. Unnikrishnan Nair. Minimum variance stratification. Communications in Statistics: Simulation and Computation, 24(1):275–284, 1995. CODEN CSSCDB. ISSN 0361-0918.

Upadhyay:1992:BEP

[US92] S. K. Upadhyay and U. Singh. Bayes estimators in the presence of a guess value. Communications in Statistics: Simulation and Computation, 21(4):1181–1198, 1992. CODEN CSSCDB. ISSN 0361-0918.

Upadhyay:1996:EEP

[USS96] S. K. Upadhyay, U. Singh, and V. Shastri. Estimation of exponential parameters under multiply Type II censoring. Communications in Statistics: Simulation and Computation, 25(3): 801–815, 1996. CODEN CSSCDB. ISSN 0361-0918.

Vangel:1994:OSN

[Van94] Mark G. Vangel. One-sided nonparametric tolerance limits. Communications in Statistics: Simulation and Computation, 23 (4):1137–1154, 1994. CODEN CSSCDB. ISSN 0361-0918.

Vaughan:1990:CTS

[Vau90] D. C. Vaughan. Comparison of test statistics for the correlation coefficient in bivariate normal samples with Type II censoring. Communications in Statistics: Simulation and Computation, 19 (2):513–526, 1990. CODEN CSSCDB. ISSN 0361-0918.

Vaughan:1992:EVV

[Vau92] D. C. Vaughan. Expected values, variances and covariances of order statistics for Student’s t-distribution with two degrees of freedom. Communications in Statistics: Simulation and Computation, 21(2):391–404, 1992. CODEN CSSCDB. ISSN 0361-0918.

Vos:1997:BNC

[VC97] Paul W. Vos and Thomas C. Chenier. Better nonparametric confidence intervals for the mean of a symmetric distribution. Communications in Statistics: Simulation and Computation, 26 (2):577–590, 1997. CODEN CSSCDB. ISSN 0361-0918.

Vos:1999:TCC

[VC99] Paul W. Vos and Thomas C. Chenier. Trimmed conditional confidence intervals for a shift between two populations. Communications in Statistics: Simulation and Computation, 28(1):99–114, 1999. CODEN CSSCDB. ISSN 0361-0918.

Vidakovic:1994:SES

[VI94] Brani Vidakovic and David Rios Insua. Some efficient simple rules in Γ-minimax estimation. Communications in Statistics: Simulation and Computation, 23(2):393–413, 1994. CODEN CSSCDB. ISSN 0361-0918.

Vidakovic:1993:EAM

[Vid93] Brani Vidakovic. On the efficiency of affine minimax rules in estimating a bounded multivariate normal mean. Communications in Statistics: Simulation and Computation, 22(3):655–669, 1993. CODEN CSSCDB. ISSN 0361-0918.

Viveros:1990:ANP

[Viv90] R. Viveros. An approximate normalizing power transformation for the F distribution. Communications in Statistics: Simulation and Computation, 19(1):57–69, 1990. CODEN CSSCDB. ISSN 0361-0918.

Vanbrackle:1997:ECC

[VR97] Lewis N. Vanbrackle III and Marion R. Reynolds, Jr. EVVMA and cusum control charts in the presence of correlation. Communications in Statistics: Simulation and Computation, 26(3): 979–1008, 1997. CODEN CSSCDB. ISSN 0361-0918.

Willemain:1992:REM

[WAD^{+}92] Thomas R. Willemain, Ali Allahverdi, Philip Desautels, Janine Eldredge, Ozden Gur, Gregory Panos, Aparna Srinivasan, Johan Surtihadi, Erkan Topal, and Mark Miller. Robust estimation methods for exponential data: a Monte-Carlo comparison. Communications in Statistics: Simulation and Computation, 21(4): 1043–1075, 1992. CODEN CSSCDB. ISSN 0361-0918.

Walker:1990:IDF

[Wal90] Esteban Walker. Influence diagnostics for fractional principal components estimators in regression. Communications in Statistics: Simulation and Computation, 19(3):919–933, 1990. CO-DEN CSSCDB. ISSN 0361-0918.

Wang:1990:RCCa

[Wan90a] Ming C. Wang. On ranges of confidence coefficients for confidence intervals on variance components. Communications in Statistics: Simulation and Computation, 19(4):1165–1178, 1990. CODEN CSSCDB. ISSN 0361-0918.

Wang:1990:RCCb

[Wan90b] Ming C. Wang. On ranges of confidence coefficients for confidence intervals on variance components. Communications in Statistics: Simulation and Computation, 19(4):1165–1178, 1990. CODEN CSSCDB. ISSN 0361-0918.

Wang:1991:ACI

[Wan91] C. Ming Wang. Approximate confidence intervals on positive linear combinations of expected mean squares. Communications in Statistics: Simulation and Computation, 20(1):81–96, 1991. CODEN CSSCDB. ISSN 0361-0918.

Wang:1992:PIB

[Wan92] C. Ming Wang. Prediction intervals for a balanced one-way random-effects model. Communications in Statistics: Simulation and Computation, 21(3):671–687, 1992. CODEN CSSCDB. ISSN 0361-0918.

Wang:1994:EAD

[Wan94] C. Ming Wang. On estimating approximate degrees of freedom of chi-squared approximations. Communications in Statistics: Simulation and Computation, 23(3):769–788, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Wang:1997:GFS

[Wan97] Yuedong Wang. GRKPACK fitting smoothing spline ANOVA models for exponential families. Communications in Statistics: Simulation and Computation, 26(2):765–782, 1997. CODEN CSSCDB. ISSN 0361-0918.

Wasserman:1990:MPC

[Was90] Gary S. Wasserman, Ph. D., P. E. Matching performance of continuous sampling plans with single sampling plans. Communications in Statistics: Simulation and Computation, 19(4): 1303–1317, 1990. CODEN CSSCDB. ISSN 0361-0918.

Wasserman:1993:ESA

[Was93] Gary S. Wasserman. An efficient search algorithm for design of a cusum based procedure for process monitoring. Communications in Statistics: Simulation and Computation, 22(3):889–909, 1993. CODEN CSSCDB. ISSN 0361-0918.

Wolfe:1990:CPM

[WC90] Douglas A. Wolfe and Yun-Shiow Chen. The changepoint problem in a multinomial sequence. Communications in Statistics: Simulation and Computation, 19(2):603–618, 1990. CODEN CSSCDB. ISSN 0361-0918.

Wu:1999:MCP

[WC99] Shu-Fei Wu and Hubert J. Chen. Multiple comparison procedures with the average for scale families: normal and exponential populations. Communications in Statistics: Simulation and Computation, 28(1):73–98, 1999. CODEN CSSCDB. ISSN 0361-0918.

Wan-Ching:1996:GLR

[WCWB96] Liu Wan-Ching, J. Arthur Woodward, and Douglas G. Bonett. The generalized likelihood ratio test for the Pearson correlation. Communications in Statistics: Simulation and Computation, 25 (2):507–520, 1996. CODEN CSSCDB. ISSN 0361-0918.

Weerakkody:1996:THA

[Wee96] Govinda J. Weerakkody. Testing hypotheses about the common mean of two normal populations. Communications in Statistics: Simulation and Computation, 25(4):1075–1091, 1996. CODEN CSSCDB. ISSN 0361-0918.

Weng:1996:END

[Wen96] T. S. Weng. Evaluation of a new diagnostic test against a reference test less than perfect in accuracy. Communications in Statistics: Simulation and Computation, 25(2):533–555, 1996. CODEN CSSCDB. ISSN 0361-0918.

Wieczorkowski:1999:EEI

[WG99] Robert Wieczorkowski and Przemysław Grzegorzewski. Entropy estimators-improvements and comparisons. Communications in Statistics: Simulation and Computation, 28(2):541–567, 1999. CODEN CSSCDB. ISSN 0361-0918.

Wiklund:1997:PCC

[Wik97] Stig Johan Wiklund. Parabolic cusum control charts. Communications in Statistics: Simulation and Computation, 26(1): 107–123, 1997. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:1990:CBM

[Wil90a] Rand R. Wilcox. Comparing biweight measures of location in the two-sample problem. Communications in Statistics: Simulation and Computation, 19(4):1231–1245, 1990. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:1990:CVM

[Wil90b] Rand R. Wilcox. Comparing variances and means when distributions have non-identical shapes. Communications in Statistics: Simulation and Computation, 19(1):155–173, 1990. CO-DEN CSSCDB. ISSN 0361-0918.

Wilcox:1998:SRE

[Wil98] Rand R. Wilcox. Simulation results on extensions of the Theil– Sen regression estimator. Communications in Statistics: Simulation and Computation, 27(4):1117–1126, 1998. CODEN CSS-CDB. ISSN 0361-0918.

Wasserman:1991:PCB

[WMF91] Gary S. Wasserman, Hassan A. Mohsen, and LeRoy A. Franklin. A program to calculate bootstrap confidence intervals for the process capability index c_{pk}. Communications in Statistics: Simulation and Computation, 20(2–3):497–510, 1991. CODEN CSS-CDB. ISSN 0361-0918.

Wollan:1992:PRN

[Wol92] Peter C. Wollan. A portable random number generator for parallel computers. Communications in Statistics: Simulation and Computation, 21(4):1247–1254, 1992. CODEN CSSCDB. ISSN 0361-0918.

Wolfinger:1993:CSS

[Wol93] Russ Wolfinger. Covariance structure selection in general mixed models. Communications in Statistics: Simulation and Computation, 22(4):1079–1106, 1993. CODEN CSSCDB. ISSN 0361-0918.

Wolstenholme:1996:AWD

[Wol96] Linda C. Wolstenholme. An alternative to the Weibull distribution. Communications in Statistics: Simulation and Computation, 25(1):119–137, 1996. CODEN CSSCDB. ISSN 0361-0918.

Wong:1992:MAV

[Won92] H. Y. Su Wong. Maximum asymptotic variances of M-estimators of location with auxiliary scale estimator under symmetric contamination. Communications in Statistics: Simulation and Computation, 21(3):783–805, 1992. CODEN CSSCDB. ISSN 0361-0918.

Wood:1994:SMF

[Woo94] Andrew T. A. Wood. Simulation of the von Mises Fisher distribution. Communications in Statistics: Simulation and Computation, 23(1):157–164, 1994. CODEN CSSCDB. ISSN 0361-0918.

Wasserman:1993:SRS

[WS93] Gary S. Wasserman and Agus Sudjianto. Short run SPC based upon the second order dynamic linear model for trend detection. Communications in Statistics: Simulation and Computation, 22 (4):1011–1036, 1993. CODEN CSSCDB. ISSN 0361-0918.

Wu:1998:CST

[WW98] Minta Wu and Robert F. Woolson. A comparison of several tests for drug synergy. Communications in Statistics: Simulation and Computation, 27(2):303–327, 1998. CODEN CSSCDB. ISSN 0361-0918.

Williams:1998:CTD

[WZ98] Christopher J. Williams and Lifen Zhou. A comparison of tests for detecting genetic variance from human twin data based on absolute intratwin differences. Communications in Statistics: Simulation and Computation, 27(1):51–65, 1998. CODEN CSSCDB. ISSN 0361-0918.

Xie:1997:SED

[XGX97] M. Xie, T. N. Goh, and W. Xie. A study of economic design of control charts for cumulative count of conforming items. Communications in Statistics: Simulation and Computation, 26(3): 1009–1027, 1997. CODEN CSSCDB. ISSN 0361-0918.

Yeh:1998:RPC

[YB98] Arthur B. Yeh and Sumankumar Bhattcharya. A robust process capability index. Communications in Statistics: Simulation and Computation, 27(2):565–589, 1998. CODEN CSSCDB. ISSN 0361-0918.

Yeh:1993:OSV

[YC93] Lam Yeh and Lau Lap Cheung. Optimal single variable sampling plans. Communications in Statistics: Simulation and Computation, 22(2):371–386, 1993. CODEN CSSCDB. ISSN 0361-0918.

Yeh:1996:BPC

[Yeh96] Arthur B. Yeh. Bootstrap percentile confidence bands based on the concept of curve depth. Communications in Statistics: Simulation and Computation, 25(4):905–922, 1996. CODEN CSS-CDB. ISSN 0361-0918.

Yokoyama:1993:EPA

[YMS93] Shinichiro Yokoyama, Norihiko Miyawaki, and Heihachi Sakamoto. On the estimation problem in the analysis of the time-dependent Poisson process. Communications in Statistics: Simulation and Computation, 22(2):591–614, 1993. CODEN CSSCDB. ISSN 0361-0918.

Yinsheng:1994:SSV

[YPW94] Qu Yinsheng, Marion R. Piedmonte, and George V. Williams. Small sample validity of latent variable models for correlated binary data. Communications in Statistics: Simulation and Computation, 23(1):243–269, 1994. CODEN CSSCDB. ISSN 0361-0918.

Yu:1994:PML

[YSM94] Hongjian Yu, Shayle R. Searle, and Charles E. McCulloch. Properties of maximum likelihood estimators of variance components in the one-way classification, balanced data. Communications in Statistics: Simulation and Computation, 23(4):897–914, 1994. CODEN CSSCDB. ISSN 0361-0918.

Yadava:1996:EPM

[YSY96] K. N. S. Yadava, H. Singh, and Surendar S. Yadava. Estimation of population mean in successive sampling using information on auxiliary character. Communications in Statistics: Simulation and Computation, 25(1):225–233, 1996. CODEN CSSCDB. ISSN 0361-0918.

Yadava:1996:SPP

[YY96] K. N. S. Yadava and Surender S. Yadava. Simultaneous projection of population and households: comparative study of methodologies used. Communications in Statistics: Simulation and Computation, 25(3):835–848, 1996. CODEN CSSCDB. ISSN 0361-0918.

Zhang:1994:APE

[ZB94] Ji Zhang and Dennis D. Boos. Adjusted power estimates in Monte Carlo experiments. Communications in Statistics: Simulation and Computation, 23(1):165–173, 1994. CODEN CSS-CDB. ISSN 0361-0918.

Zhang:1994:IOT

[Zha94] Zhiyi Zhang. On improving omnibus tests in metaanalysis using vote counts. Communications in Statistics: Simulation and Computation, 23(3):803–812, 1994. CODEN CSSCDB. ISSN 0361-0918.

Zheng:1995:NAN

[Zhe95] Qi Zheng. A note on an approximation to Neyman Type A distributions. Communications in Statistics: Simulation and Computation, 24(4):889–895, 1995. CODEN CSSCDB. ISSN 0361-0918.

Zhenlin:1997:MEB

[Zhe97] Yang Zhenlin. More on the estimation of Box–Cox transformation. Communications in Statistics: Simulation and Computation, 26(3):1063–1074, 1997. CODEN CSSCDB. ISSN 0361-0918.

Zheng:1998:TRS

[Zhe98] Qi Zheng. On a trinity role of the survival and hazard functions of the two-stage carcinogenesis models. Communications in Statistics: Simulation and Computation, 27(1):137–145, 1998. CODEN CSSCDB. ISSN 0361-0918.

Zhou:1992:DMT

[Zho92] Mai Zhou. Difference of means test with censored data. Communications in Statistics: Simulation and Computation, 21(3): 697–706, 1992. CODEN CSSCDB. ISSN 0361-0918.

Zimmerman:1993:ENS

[ZWZ93] Donald W. Zimmerman, Richard H. Williams, and Bruno D. Zurabo. Effect of nonindependence of sample observations on some parametric and nonparametric statistical tests. Communications in Statistics: Simulation and Computation, 22(3):779– 789, 1993. CODEN CSSCDB. ISSN 0361-0918.